### v1

# MetaTrading: An Immersion-Aware Model Trading Framework for Vehicular Metaverse Services

Abstract
Updates of extensive Internet of Things (IoT) data are critical to the immersion of vehicular metaverse services. However, providing high-quality and sustainable data in unstable and resource-constrained vehicular networks remains a significant challenge. To address this problem, we put forth a novel immersion-aware model trading framework that incentivizes metaverse users (MUs) to contribute learning models trained by their latest local data for augmented reality (AR) services in the vehicular metaverse, while preserving their privacy through federated learning. To comprehensively evaluate the contribution of locally trained learning models provided by MUs to AR services, we design a new immersion metric that captures service immersion by considering the freshness and accuracy of learning models, as well as the amount and potential value of raw data used for training. We model the trading interactions between metaverse service providers (MSPs) and MUs as an equilibrium problem with equilibrium constraints (EPEC) to analyze and balance their costs and gains. Moreover, considering dynamic network conditions and privacy concerns, we formulate the reward decisions of MSPs as a multi-agent Markov decision process. Then, a fully distributed dynamic reward method based on deep reinforcement learning is presented, which operates without any private information about MUs and other MSPs. Experimental results demonstrate that the proposed framework can effectively provide higher-value models for object detection and classification in AR services on real AR-related vehicle datasets compared to benchmark schemes.
Index Terms:
Equilibrium problem with equilibrium constraints, immersion-aware, incentive mechanism, resource allocation, vehicular metaverse.
I Introduction
The metaverse, envisioned as one of the next evolution directions of the Internet, is designed to create a fully immersive, self-sustaining virtual environment in which people can engage in activities such as playing, working, and socializing [1]. This vision is propelled by advances in the fifth-generation (5G) and the forthcoming sixth-generation (6G) communication technologies, renowned for their low latency and high data throughput. These technologies play a critical role in seamlessly integrating the Internet of Things (IoT) data into metaverse services, thus bringing the once-fictional concept of immersive experiences closer to reality.
Metaverse services are beginning to reveal their vast potential across a broad spectrum of industries, from gaming and autonomous driving to education and marketing. Notably, the application of vehicles within the metaverse has captured significant interest due to the enhanced safety features and immersive experiences offered by state-of-the-art augmented reality (AR) technologies.
Figure 1: An example of AR services in the vehicular metaverse [^2] Image source: https://www.jasoren.com/ar-in-automotive/.
Market report [2] forecasted that the global automotive metaverse market will grow from $1.9$ billion in 2022 to $6.5$ billion by 2030. Automakers like BMW have recently doubled down on AR technology in their projects. As shown in Fig. 2, augmented information makes driving safer by showing potential hazards hidden behind the vehicle in front of you. Moreover, Nissan’s [3] upcoming technology utilizes a 3D AR interface that merges the real world with the virtual world to provide the driver with augmented information about the surrounding area. In addition, Vanarama has also implemented a visual parking assistance function using AR [4], which is expected to identify empty spaces in busy car parks and display useful information, such as parking fees, on the windscreen.
The ability to capture information from the “real” world, particularly the ability to collect and process massive data from IoT devices, is the key to determining the success of immersive services (e.g., AR) in the vehicular metaverse. Meanwhile, the data must be processed and presented in a meaningful, responsive, and appropriately protected way. Technically, a high-quality experience with AR services relies on accurate detection and classification of real-world objects (e.g., cars and pedestrians) under complex conditions [5]. To achieve this goal, sufficient valid data needs to be collected and processed in depth to detect and classify objects accurately. Therefore, it is essential to focus on effectively collecting, processing, and protecting the data that supports a safer and more enjoyable driving experience.
Motivations. An existing widely-used data collection method, as adopted by Nissan [6], involves collecting massive amounts of data through vehicle sensors, cameras, and roadside devices, and then processing all the data centrally. However, when it comes to the situation with multiple metaverse users (MUs) and metaverse service providers (MSPs) on behalf of different companies, the centralized data collection approach is not applicable and may lead to the following issues. First, AR services in the vehicular metaverse need to be highly immersive so that MUs feel fully immersed in the rendered environment, such as visualized driving. However, providing a seamless AR experience is challenging due to the latency caused by massive real-time data updates under unstable and resource-limited communication conditions. Note that the value of real-time data diminishes over time [7]. Also, delays can severely impact the MU’s experience and cause dizziness [8]. Second, the data to be transmitted may be sensitive and private, such as location, movement, and biometrics, which can create a better immersive experience but may inevitably increase the privacy risk of MUs [9]. Third, it is difficult to share data between MSPs due to competition. This leads to the same data being sampled multiple times by different MSPs in the physical world, resulting in wasted resources (e.g., sampling costs) and data redundancy. Moreover, tremendous data pressure and computational burden are imposed on MSPs. Therefore, an efficient method for collecting high-quality data while preserving privacy is needed to achieve immersive AR services.
Challenges. To address the above issues, we propose an immersion-aware model trading framework designed to provide data support for AR services and ensure privacy by integrating federated learning (FL). Adopting FL enables MUs to contribute diverse learning models to multiple MSPs using local data, thus effectively protecting sensitive information [10]. Moreover, learning models are uploaded to MSPs for AR services by individual MUs, rather than centrally transmitting large amounts of data, which significantly reduces the communication burden. Nonetheless, developing such a model trading framework still faces the following challenges:
  • MUs are typically self-interested and are reluctant to share learning models with MSPs due to the additional computation, communication, and energy overhead. Thus, the challenge here is how to incentivize them to become contributors in providing high-value learning models that will benefit vehicular metaverse services.
  • It is difficult to design an appropriate metric to comprehensively evaluate the value of learning models contributed by MUs due to the diverse aspects involved in model valuation.
  • MUs have different sampling costs and limited computation and communication resources, while MSPs differ in their model preferences and compete with each other. Hence, the next challenge is to model the dynamic competitive interactions among MSPs and to achieve an equilibrium of interests among MUs and MSPs in the model trading.
Figure 2: The outline of an immersion-aware framework for FL-assisted vehicular metaverse.
Contributions. To address these challenges, we model the dynamic competitive trading interactions as an equilibrium problem with equilibrium constraints (EPEC) involving multiple leaders (MSPs) and multiple followers (MUs), which is a hierarchical optimization problem with equilibria at two levels [11]. Furthermore, to evaluate the value of learning models contributed by MUs to MSPs, we design a new metric called the immersion of the learning model (IoM). The metric jointly considers the freshness and accuracy of the learning model, as well as the amount and potential value of raw data used for training. Moreover, given the dynamic networks and the privacy concerns of MSPs, we formulate the reward decisions of MSPs as a multi-agent Markov decision process (MAMDP) and develop a multi-agent deep reinforcement learning (DRL)-based dynamic reward (MDDR) approach to obtain the reward decisions in a fully distributed manner. The outline of the proposed framework is shown in Fig. 2, and the main contributions of this paper are summarized as follows:
  • We propose an immersion-aware trading framework for learning models with FL assistance, incentivizing MUs to contribute high-value learning models for MSPs in a privacy-preserving and efficient manner. Moreover, a novel metric called “IoM” is designed to quantify the immersion enhancement provided by MUs for AR services.
  • We model the dynamic competitive trading interactions as an EPEC and provide theoretical support for the existence and uniqueness of equilibria at two levels. Moreover, a fully distributed MDDR approach is developed to adapt to dynamic environments and address complex reward decisions of MSPs without accessing any private information of MUs or MSPs.
  • We conduct extensive numerical simulations based on AR-related vehicle datasets to validate the efficacy and efficiency of MDDR and the proposed immersion-aware model trading framework. Compared to benchmark schemes, the framework better motivates MUs to provide higher-value learning models for object detection and classification functions of AR services effectively.
The rest of this paper is organized as follows. Section II discusses the related work. In Section III, we present the system overview and design the immersion metric of the learning model. Section IV gives the game formulation, and Section V analyzes the existence of the equilibria at two levels. In Section VI, we give the detailed design of MDDR. Section VII shows numerical experiments to evaluate the framework performance, and finally Section VIII concludes the paper.
II Related Work
Incentive Mechanism. Incentive Mechanisms have been widely applied to metaverse services [12, 13, 14, 15, 16, 17]. For example, Xu et al. [12] proposed an incentive framework for learning-based virtual reality (VR) services and designed a perceptual quality to measure VR users’ immersion in the virtual world. Sun et al. [13] investigated dynamic digital twins in air-assisted vehicular networking and designed a two-stage resource allocation incentive mechanism based on the Stackelberg game. Jiang et al. [14] developed a hierarchical game-theoretic coded distributed computing (CDC) framework for metaverse services. The framework employs coalition and Stackelberg games to select reliable CDC workers for MSPs and incentivize reliable CDC workers to participate in metaverse services. Xi et al. [16] proposed an incentive-based congestion control scheme for stochastic demand response in digital twin edge networks, using Lyapunov optimization and contract theory. Nguyen et al. [17] introduced MetaChain, an innovative blockchain-based framework designed to address the emerging challenges in developing metaverse applications. The framework integrates game theory to motivate MUs to contribute to the metaverse.
However, the above incentives for metaverse services mainly focus on digital twins, blockchain, or virtual reality, while ignoring the scenarios of multiple MSPs and multiple MUs in the vehicular metaverse. In this scenario, the complexity of interactions among MSPs and MUs requires novel incentive mechanisms to better facilitate the development and innovation of metaverse services.
AR and FL. The AR applications, characterized by smoothness and immersion requirements, provide immersive and interactive experiences for MUs by seamlessly integrating virtual content with the real world. For a better immersion experience, Cozzolino et al. [18] developed a latency-energy efficient task offloading for AR services and designed a best-effort resource provisioning algorithm to solve the problem of heterogeneous resource competition. Further, to reduce latency while protecting data privacy, Chen et al. [19] proposed a framework that integrates mobile edge computing and FL to solve the computational efficiency and low-latency object classification problem for AR services. In FL, MUs can train AI models on their local devices, thus enabling privacy-preserving collaborative machine learning while fully leveraging the computation resources of those MUs [20]. Furthermore, Lu et al. [21] investigated communication-efficient FL for digital twin edge networks in industrial IoT, which provides virtual and real synchronization while protecting privacy. Incorporating FL into AR services brings the benefits of reduced communication latency and privacy protection. However, the existing studies fail to consider the integration of FL into AR for the multi-MSP and multi-MU metaverse scenario, while overlooking the selfishness of MUs, i.e., whether they are willing to participate in FL learning.
Inspired by the above existing works, we design an immersion-aware framework integrated FL to provide a learning model trading platform for MSPs and MUs in the vehicular metaverse service markets, enabling MUs to contribute high-value learning models for AR services.
Figure 3: The workflow of the immersion-aware model trading framework [^4] Image source: https://mixed-news.com/en/google-transforms-google-maps-into-the-backbone-of-our-ar-future/, https://about.facebook.com/meta, https://www.vanarama.com/blog/cars/4-ways-augmented-reality-will-revolutionise-the-automotive-industry..
III System Model
We first present the system overview of the immersion-aware model trading framework in Section III-A. We next introduce the FL mechanism adopted by our framework in Section III-B and design the immersion metric of the learning model in Section III-C.
III-A System Overview
The workflow of the immersion-aware model trading framework is illustrated in Fig. 4, aiming to provide MUs with immersive and interactive AR service experiences. MUs can contribute learning models to the AR services of MSPs for better immersive experiences. For example, services such as object detection can be improved by training on images of streets, pedestrians, and so on, captured by vehicles’ cameras. In this way, AR services can be better applied to visible driving, with more accurate and timely hazard warnings from windscreens. The framework mainly consists of MSPs owned by different companies (e.g., Meta, Nissan, and Google), an infrastructure layer, and an interaction layer with the MUs that enjoy vehicular metaverse services. MSPs are fueled by many technologies that work together to support vehicular metaverse services, including AI engines, AR/VR engines, metaverse trading engines, etc. The infrastructure layer with base stations (equipped with caches and edge servers) provides the basis of 5G/6G communication services for the interactions among MUs and MSPs. MUs play essential roles as both consumers and contributors to vehicular metaverse services.
TABLE I: Parameter Definitions and Notations
Notation Definition
$M$ Number of MUs
$N$ Number of MSPs
$\overline{\Delta}_{mn}$ The average AoI of MU $m$’s model provided for MSP $n$
$V_{mn}$ The IoM contributed by MU $m$ to MSP $n$
$f_{mn}$ MU $m$’s computation resource used for MSP $n$
$B_{mn}$ MU $m$’s communication resource used for MSP $n$
$X_{mn}$ MU $m$’s training data set for MSP $n$
$\omega_{mn}$ The potential value of MU $m$’s local data for MSP $n$
$\theta_{m}$ MU $m$’s local accuracy threshold
The framework is a multi-engine configuration, and this paper focuses on two components: the trading engine and the AI engine of the MSP. The trading engine is responsible for incentivizing MUs to participate in mutually beneficial collaborations with MSPs, providing them with a trading platform to make their decisions (i.e., computation and communication resource allocation of MUs and reward decisions of MSPs). Then, each MU deciding to contribute utilizes the FL mechanisms integrated into AI engines to provide learning models for MSPs under the strategic guidance of the trading results. The specific workflow of the framework involves the following two phases.
Phase I (Incentive process): Initially, MSPs’ AI engines broadcast their global models to all MUs in the vehicular metaverse [^5] The FL global models owned by different MSPs we consider could be object detection or classification models for intelligent terrain mapping, intuitive road safety, visible driving, etc. Note that it is feasible to train on multiple tasks with the same local data. For example, for the same image, intuitive road safety focuses on pedestrians and vehicles, while intelligent terrain mapping focuses on landmarks and routes. (① in Fig. 4). Then, the trading engines of MSPs determine their digital currency prices per IoM, i.e., rewards to MUs, and broadcast them to MUs (②). Next, MUs determine their allocation of computation and communication resources for contributing learning models to various MSPs based on the IoM, costs, and rewards given by MSPs. (③). Based on the responses from MUs, the trading engines of MSPs adjust the digital currency prices to maximize their benefits. Steps ② and ③ are repeated until an agreement is reached among MUs and MSPs.
Phase II (FL process): Each MU, guided by decisions obtained through tradings, allocates computation resources to local training based on different global models from MSPs and generates local learning models for different MSPs (④). Then, MUs upload their updated learning models to the corresponding AI engines of MSPs by consuming communication resources (⑤). Following this, AI engines aggregate models from all MUs to update their global models and transmit them to their AR/VR engines for utilization (⑥). Finally, the AR/VR engines enable better AR services with the assistance of the AI engines (⑦). Steps ④-⑦ are repeated until the guidance time $T$ ends. To ease the presentation, we summarize some important notations in Table I.
III-B FL Framework Mechanism Adopted
Here, we consider a set of =$\mathcal{M}=\left\{1,\ldots,M\right\}$ MUs with the capacities of local computation and a set of $\mathcal{N}=\left\{1,\ldots,N\right\}$ MSPs with FL synchronous tasks [^6] That is, the update time of the global model is limited by the slowest MU, i.e., the AI engine of the MSP must wait to receive updated models from all MUs before performing model aggregation. The advantages of synchronous FL are high model accuracy and fast convergence [22, 23].. Each MSP initiates an FL synchronous task with a virtual deadline $\tau_{n}(n\in\mathcal{N})$. The MUs can join FL to generate local learning models and contribute to enhancing vehicular metaverse services. As shown in Fig. 4, MU $m$ provides learning models to different MSPs simultaneously for a given time period $T$, meaning multiple FL tasks can be processed in parallel. Moreover, the data collected at the $(r-1)$-th round is used as the local data for training at the $r$-th round. Without loss of generality, we consider that the trading decisions among MUs and MSPs are made at the beginning of the period $T$.
Figure 4: An illustration of the FL mechanism with $n$ tasks for MU $m$.
To be concrete, let $X_{mn}$ be the set of data samples for task $n$ of MU $m$ as a set of input-output pairs $\sum_{i=1}^{|X_{mn}|}(y_{i},z_{i})$, where $y_{i}$ is an input sample vector with $d$ features, $z_{i}$ is the labeled output (i.e., ground truth) value of sample $y_{i}$, and $|X_{mn}|$ is the size of set $X_{mn}$. The data, such as images of streets and pedestrians, can be generated by high-definition cameras inside and outside the vehicles. The steps involved in each iteration $r\in\left\{1,2,\ldots,R\right\}$ are as follows:
  • Publishing Tasks : The AI engine of MSP $n$ broadcasts its global model $\jmath_{n}^{(r)}$ (e.g., object detection models used in AR services for visible driving and intuitive road safety) to all MUs in the $r$-th round.
  • Local Training: Each MU, for example, MU $m$, trains $\jmath_{n}^{(r)}$ with the most recently collected data $X_{mn}^{(r)}$ by stochastic gradient descent (SGD) within certain local rounds. The number of local training rounds depends on its local accuracy threshold.
  • Uploading and Aggregation: MUs transmit their updated local model $\jmath_{mn}^{(r)}$ to the AI engine of MSP $n$. After receiving all the updated models, the MSP $n$’s AI engine aggregates them through the following weighted average:
    $\jmath_{n}^{(r+1)}$ $=\jmath^{(r)}_{n}+\sum_{m=1}^{M}\frac{|X_{mn}^{(r)}|}{|X_{n}^{(r)}|}(\jmath_{n}^{(r)}-\jmath_{mn}^{(r)}),$ (1)
    and then obtains the new global model $\jmath_{n}^{(r+1)}$ for the next iteration.
  • Model Utilization: After global aggregation, MSP $n$’s AR engine obtains an updated global model that can be used for corresponding 3D modeling and other functions [19]. The global rounds are iterated until a specific requirement is met, such as reaching a certain level of accuracy or a deadline.
III-C Immersion Metric of Learning Model
MSPs determine the rewards for MUs based on the immersion of the learning model (IoM) metric. Correspondingly, MUs need to decide the IoMs of the models provided to MSPs by allocating their computation and communication resources to maximize the benefits. This process highlights the key role of “IoM” in the interactions between the two parties, serving as both an evaluation criterion and a basis for decision-making. Unlike traditional FL and incentive mechanisms, in vehicular metaverse scenarios with immersion requirements, both the contribution and freshness of the learning model need to be considered [^7] For example, an MSP may receive fresher learning models but with lower contributions, or less fresh learning models but with higher contributions. , which will collectively affect the immersive experience of AR services, i.e., whether virtual objects can be placed in the physical world accurately and promptly. With the above consideration, we propose the IoM, a metric that jointly considers the contribution prediction $I_{mn}$ and the age of information (AoI) $\Delta_{mn}$ [24] to measure the value of a learning model that MU $m$ brought to the vehicular metaverse service of MSP $n$, denoted as
$V_{mn}=I_{mn}(\tau_{n}-\overline{\Delta}_{mn}),$ (2)
where the components are detailed as follows. In the rest of this paper, we refer to the trading engine of MSP as MSP.
III-C1 Contribution Prediction $I$
The contribution prediction $I_{mn}$ from MU $m$ to MSP $n$ is determined by the accuracy $\theta_{m}$ of the learning model, the total amount of training data $\lfloor\frac{T}{\tau_{n}}\rfloor|X_{mn}|$, and the potential value $\omega_{mn}$ of local data. Prior work [25] characterized the contribution of a learning model in terms of a logarithmic function of the amount of local training data. However, evaluating the contribution solely by the training data size is one-sided, as there may be a large amount of redundant data or the validity of the learning model cannot be guaranteed. It is more practical to comprehensively evaluate the amount of training data, model accuracy, and the potential value of local data; thus, we denote $I_{mn}$ as
$I_{mn}=\frac{\omega_{mn}\epsilon ln(1+\eta\lfloor\frac{T}{\tau_{n}}\rfloor|X_{mn}|)}{\theta_{m}},$ (3)
where $\lfloor\frac{T}{\tau_{n}}\rfloor$ is the number of iterations that the task of MSP $n$ can be performed within $T$. $\epsilon$ and $\eta$ are system parameters from experiments as obtained in [26]. $\theta_{m}\in\left(0,1\right)$ characterizes the accuracy of local training as decided by the MU. Here, $\theta\rightarrow 0$ means that a high-accuracy learning model is available, while $\theta\rightarrow 1$ implies that the accuracy of the learning model is low. The local data potential value $\omega_{mn}$ describes the difference between model predictions and true labels, defined as
$\omega_{mn}=\frac{1}{|X_{mn}|}\sum_{i=1}^{|X_{mn}|}(\hat{z}_{i}-z_{i})^{2},$ (4)
where $\hat{z}_{i}$ denotes the inferred label from the current global model and $z_{i}$ is its ground-truth label. If the value of $\omega_{mn}$ is large, it indicates that the current global model is not working well on local data or there are unseen samples. Therefore, updating the model by training on these data may lead to better performance of the global model. If the value of $\omega_{mn}$ is small, it means that the current data has little potential to further improve the global model since the knowledge of the data has already been learned or the data size is small. Note that to shorten the decision time, we use the potential value of the initial local data to be a proxy for the average potential value of all data within $T$. This is a reasonable approach when $T$ is not too long.
Figure 5: The AoI evolution of the models sent from MU $m$ to MSP $n$.
III-C2 Age of Information $\Delta$
The whole FL process involves model training, uploading, and waiting, which together constitute a complete cycle of model updating. From the perspective of MSP $n$, the age evolution of the model contributed by MU $m$ is shown in Fig. 5. Here, the filled circles denote the time instances when MU $m$ starts training the learning model. The intersections show the time instances when MSP $n$ receives the corresponding updated model, at which point the AoI drops to the lowest. We denote that at time $t$, the generation time of the last update received by MSP $n$ from MU $m$ is $s(t)$. Then, the instantaneous AoI of MU $m$’s learning model measured at MSP $n$, $\Delta_{mn}(t)$, is
$\Delta_{mn}(t)=t-s(t),$ (5)
which shows the elapsed time since the generation of the latest updated model. A smaller instantaneous AoI means that the latest model received is fresher at that time. To obtain a comprehensive understanding of the overall freshness of model updates over a certain period, we use average AoI to evaluate the freshness of the models. With the aid of Figure 5, the average AoI of the learning model from MU $m$ to MSP $n$ can be expressed as
[…]
$=\frac{1}{2}\tau_{n}+T_{mn}^{c}+T_{mn}^{t},$
where ${Q_{mn}}^{(r)}$ is the $r$-th trapezium under the curve. A lower average AoI indicates that the updated models are generally fresher over a long period.
Since MU $m$ constantly collects data throughout the FL process, the amount of training data collected for training in each round is $|X_{mn}|=x_{m}\tau_{n}$, where $x_{m}$ denotes the data (number of floats) collected per unit time. The cumulative time $T_{mn}^{c}$ for local training is determined by the amount of training data, computation resource $f_{mn}$ allocated to MSP $n$, and its local accuracy threshold $\theta_{m}$, which is equivalent to
$T_{mn}^{c}=\log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}.$ (7)
A smaller value of $\theta$ indicates higher accuracy, but leads to a higher MU cost, i.e., the number of local iterations, which is upper bounded by $\log(1/\theta_{m})$ [27].
Furthermore, the upload time $T_{mn}^{t}$ is directly related to the number of model parameters $b_{mn}$ and the resource $B_{mn}$ (i.e., uplink bandwidth) used to communicate between MU $m$ and MSP $n$, defined as
$T_{mn}^{t}=\frac{b_{mn}}{B_{mn}\log_{2}(1+\varsigma_{mn})},$ (8)
where $\varsigma_{mn}=\frac{p^{t}_{m}g_{mn}}{\sigma^{2}}$ denotes the signal-to-interference-plus-noise ratio (SINR) for the communication between MU $m$ and MSP $n$. $g_{mn}$ and $p^{t}_{m}$ are the corresponding channel gain and the transmission power of MU $m$, respectively. $\sigma^{2}$ is additive white Gaussian noise (AWGN).
Finally, by substituting (7) and (8) into (6), $\overline{\Delta}_{mn}$ can be expressed as
$\overline{\Delta}_{mn}=\frac{1}{2}\tau_{n}+\frac{x_{m}\tau_{n}\log(1/\theta_{m})}{f_{mn}}+\frac{b_{mn}}{B_{mn}\log_{2}(1+\varsigma_{mn})}.$ (9)
IV Game Formulation
We first define the utility functions of MU $m$ and MSP $n$ in Section IV-A and Section IV-B, respectively. Then, we formulate the interactions among MUs and MSPs as an equilibrium problem with equilibrium constraints (EPEC) in Section IV-C, where the equilibrium criterion exists at both the level of MUs and MSPs due to the conflicting interests between them.
IV-A Utility of MU
We define the utility of MU $m$ composed of the rewards from MSPs and the cost incurred for contribution, which is also affected by resource constraints and immersion requirements for consuming vehicular metaverse services. The cost incurred due to local training and model uploading, which can be expressed as
$C_{mn}=c_{m}^{f}\log(1/\theta_{m})f_{mn}+c_{m}^{B}B_{mn},$ (10)
where $c_{m}^{f}$ and $c_{m}^{B}$ denote the cost factors of computation and communication resources, respectively. Let $\bm{f}_{m}\triangleq\left(f_{m1},f_{m2},\ldots,f_{mn}\right)$ and $\bm{B}_{m}\triangleq\left(B_{m1},B_{m2},\ldots,B_{mn}\right)$ be the computation and communication allocation profiles of MU $m$, respectively. Then, the utility maximization problem for MU $m$ within $T$ can be formulated as
Problem 1
$\max\,\,\Phi_{m}(\bm{f}_{m},\bm{B}_{m})=\sum_{n=1}^{N}(p_{mn}V_{mn}-C_{mn}),$ (11)
$s.t.$ $C1:\sum_{n=1}^{N}f_{mn}<f_{m}^{max},\,\,\sum_{n=1}^{N}B_{mn}<B_{m}^{max},$
$C2:\log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}\log_{2}(1+\varsigma_{mn})}\leq\tau_{n},$
$C3:\frac{S_{m}}{f^{max}_{m}-\sum_{n=1}^{N}f_{mn}}<T^{req},$
where $S_{m}$ indicates the minimum computation resource required to enjoy other basic services. Constraints C1 and C3 indicate that the MU’s total computation and communication resources are limited and cannot be fully used to contribute learning models, given the MU’s demand for basic services. Constraint C2 ensures that the total time for local training and uploading cannot exceed the time constraint $\tau_{n}$ for each round. The goal of each MU is to choose the optimal allocation of computation $\bm{f}_{m}$ and communication $\bm{B}_{m}$ resources to maximize its utility. Note that $V_{mn}$ must be greater than 00. This is because if $V_{mn}\leq 0$ and MU $m$ is involved in contributing learning models, then the utility is calculated as $p_{mn}V_{mn}-C_{mn}\leq 0$, resulting in MU $m$ will not participate in the trading.
IV-B Utility of MSP
Each MSP has a gain function $\psi$ associated with the IoM to measure the benefits of all the learning models from MUs to the service. In this paper, we adopt $\psi_{n}=\mu_{n}ln(1+\sum_{m=1}^{M}V_{mn})$ as a monotonically increasing, differentiable, strictly concave IoM function, which is a simplification of the widely adopted function [28]. Let $\bm{p}_{n}\triangleq\left(p_{1n},\ldots,p_{mn}\right)$ and $\bm{p}_{\textrm{-}n}$ represent the reward profile of MSP $n$ and all other MSPs except MSP $n$, respectively. Then, the optimization problem for the MSP is defined as
Problem 2
$\max\,\,\Psi_{n}(\bm{p}_{n},\bm{p}_{\textrm{-}n})=\psi_{n}-\sum_{m=1}^{M}p_{mn}V_{mn},$ (12)
$s.t.\,\,p_{mn}>0,n\in\mathcal{N},m\in\mathcal{M},$
where $\mu_{n}$ refers to the profit conversion coefficient [29] from IoM, which is adjusted according to the AR services offered by the different MSPs. Note that the other MSPs’ decisions $\bm{p}_{\textrm{-}n}$ are captured by the MUs’ responses. Moreover, when MSP $n$ determine its reward $\bm{p}_{n}$ for different MUs, the MSP needs to consider the rewards offered by other MSPs (i.e., $\bm{p}_{\textrm{-}n}$ ) as well as the strategies of all MUs (i.e., $(\bm{f}_{m},\bm{B}_{m}),m\in\mathcal{M}$). This thereby leads to reward competition among MSPs. Therefore, the optimization among MSPs can be considered as a non-cooperative game, termed a multi-MSP rewarding game $\Omega$ as follows.
Definition 1.
A multi-MSP rewarding game $\Omega$ is a tuple $\Omega=\left\{\mathcal{N},\bm{p},\bm{\Psi}\right\}$ defined by
  • Players: The set of MSPs;
  • Strategies: The reward decisions $\bm{p}_{n}$ of any MSP $n$;
  • Utilities: The vector $\bm{\Psi}=\left\{\Psi_{1},\Psi_{2},\ldots,\Psi_{n}\right\}$ contains the utility functions of all the MSPs defined in (12).
Figure 6: The structure of EPEC.
IV-C Multi-MSP Multi-MU Game as EPEC
In such a multi-MSP and multi-MU scenario, each MU has the choice of multiple MSPs to contribute learning models, and each MSP can incentivize multiple MUs to earn high-value learning models and achieve more benefits. Accordingly, each MSP competes with others for the equilibrium, which is constrained by the equilibrium among the MUs. This leads to the EPEC formulation [30].
As shown in Fig. 6, MUs and MSPs negotiate allocation strategies of computation and communication resources and reward decisions to maximize their benefits. In the upper level, MSPs determine the rewards they are willing to offer by considering their cost, the responses of MUs, and the decisions of other MSPs. In the lower level, each MU m gives the optimal allocation response for computation and communication resources by considering its resource constraints, cost, and rewards from different MSPs. The objective of EPEC is to find the equilibria at two levels, i.e., the point at which the MSPs’ (leaders’) utilities are maximized given that the MUs (followers) will choose their best responses. For the proposed EPEC, the equilibria at two levels are defined as follows.
Definition 2.
Let $(f^{*}_{mn},B^{*}_{mn})$ and $p^{*}_{mn}$ denote the optimal computation and communication resource allocation of MU $m\in\mathcal{M}$ and the optimal reward decision of MSP $n\in\mathcal{N}$, respectively. Then, the points $(f^{*}_{mn},B^{*}_{mn})$ and $p^{*}_{mn}$ are the equilibria at two levels if the following conditions hold:
$\Phi_{m}((\bm{f}_{m}^{*},\bm{B}_{m}^{*}),\bm{p}_{n}^{*})\geq\Phi_{m}((\bm{f}_{m},\bm{B}_{m}),\bm{p}_{n}^{*}),\forall m\in\mathcal{M},$ (13)
$\Psi_{n}\left(\bm{f}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{B}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{p}_{n}^{*},\bm{p}_{-n}^{*}\right)\geq$
$\Psi_{n}\left(\bm{f}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{B}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{p}_{n},\bm{p}_{-n}^{*}\right),\forall n\in\mathcal{N}.$
In summary, the MSPs’ optimization problems are formulated as the following EPEC problems:
$\max_{p_{n}}\,\,\Psi_{n}=\psi_{n}-\sum_{m=1}^{M}\bigg{[}p_{mn}I_{mn}\bigg{(}\frac{1}{2}\tau_{n}$ (14)
$-\frac{x_{m}\tau_{n}\log(1/\theta_{m})}{f_{mn}^{*}}-\frac{b_{mn}}{B_{mn}^{*}\log_{2}(1+\varsigma_{mn})}\bigg{)}\bigg{]},$
[…]
To investigate the above EPEC, we address the lower level (Problem 1) and the upper level (Problem 2) by using the backward induction methods in the following section.
V Analysis and Solutions
In this section, we use backward induction to investigate the existence and uniqueness of the equilibria at two levels, i.e., the optimal computation and communication resource allocation of MUs at the lower level and the optimal reward decisions of MSPs at the upper level.
V-A Lower Level: Optimal Computation and Communication Resource Allocation Analysis
For any reward decisions $\bm{p}$ given by MSPs, MU $m$ would like to determine an optimal decision $\bm{f}_{m}^{*}$ and $\bm{B}_{m}^{*}$ to maximize its utility. We analyze and provide the solution for the optimal allocation of computation and communication resources as follows.
Definition 3.
If the problem is strictly concave for maximization and satisfies the following conditions, it will have a unique optimal solution [31]: (1)The Utility function is concave; (2) Constraints are all affine constraints; and (3) Strategy space is a compact and convex set.
Theorem 1.
Problem 1 has a unique optimal solution in the lower level, i.e., each MU has an optimal and unique allocation scheme for computation and communication resources.
Proof.
We first observe that the Hessian matrix of $\Phi_{m}(\bm{f}_{m},\bm{B}_{m})$ with respect to $(f_{mn},B_{mn})$ is
$\bm{H}_{m}=\begin{bmatrix}\bm{H}_{m}^{f}&0\\ 0&\bm{H}_{m}^{B}\end{bmatrix},$ (15)
where $\bm{H}^{f}_{m}$ can be obtained by the second-order derivation of $\Phi_{m}(\bm{f}_{m},\bm{B}_{m})$ with respect to $f_{mn}$
$\bm{H}^{f}_{m}$ $=\left[\frac{\partial^{2}\Phi_{m}(\bm{f}_{m},\bm{B}_{m})}{\partial f_{mn}\partial f_{mn^{\prime}}}\right]_{n,n^{\prime}\in\mathcal{N}}$ (16)
$=-diag\left(h^{f}_{m1},h^{f}_{m2}\ldots,h^{f}_{mn}\right)<\bm{0},$
where $h^{f}_{mn}=p_{mn}I_{mn}\frac{2x_{m}\tau_{n}}{f_{mn}^{3}}\log(1/\theta_{m})$.
Similarly, the $\bm{H}^{B}_{m}$ can be calculated as
$\bm{H}^{B}_{m}$ $=\left[\frac{\partial^{2}\Phi_{m}(\bm{f}_{m},\bm{B}_{m})}{\partial B_{mn}\partial B_{mn^{\prime}}}\right]_{n,n^{\prime}\in\mathcal{N}}$ (17)
$=-diag\left(h^{B}_{m1},h^{B}_{m2},\ldots,h^{B}_{mn}\right)<\bm{0},$
where $h^{B}_{mn}=\frac{2b_{mn}p_{mn}I_{mn}}{B_{mn}^{3}\log_{2}(1+\varsigma_{mn})}$.
We can easily show that $\bm{H}^{f}_{m}$ and $\bm{H}^{B}_{m}$ are negative semi-definite, and thus $\bm{H}_{m}$ is negative semi-definite and $\Phi_{m}(\bm{f}_{m},\bm{B}_{m})$ is concave. Moreover, according to the convexity condition of its constraints and Definition 3, it can be proved that Problem 1 has a unique optimal solution $(\bm{f}_{m}^{*},\bm{B}_{m}^{*})$. ∎
Next, we can obtain the best response function of MU $m$ by solving Problem 1. Based on the first-order condition and setting the derivative to zero, we can derive the following proposition.
Proposition 1. Given the digital currency price offered by the MSPs and the Lagrange multipliers induced by the constraints, the best response function $f_{mn}$ of MU $m$ satisfies:
$f_{mn}^{*}=\begin{cases}\sqrt{\frac{p_{mn}I_{mn}x_{m}\tau_{n}}{c_{m}^{f}}},&\text{ $p_{mn}\geq\frac{F^{2}}{I_{mn}}$},\\ F\sqrt{\frac{x_{m}\tau_{n}}{c_{m}^{f}}},&\text{\,\,$0<p_{mn}<\frac{F^{2}}{I_{mn}}$},\\ 0,&\,\,\text{otherwise},\end{cases}$ (18)
and the best response function $B_{mn}$ of MU $m$ satisfies:
$B_{mn}^{*}=\begin{cases}\sqrt{\frac{p_{mn}I_{mn}b_{mn}}{c_{m}^{B}log_{2}(1+\varsigma_{mn})}},&\text{$p_{mn}\geq\frac{F^{2}}{I_{mn}}$},\\ F\sqrt{\frac{b_{mn}}{c_{m}^{B}log_{2}(1+\varsigma_{mn})}},&\text{\,\,$0<p_{mn}<\frac{F^{2}}{I_{mn}}$},\\ 0,&\text{otherwise},\end{cases}$ (19)
where $F=log(1/\theta_{m})\sqrt{\frac{x_{m}c_{m}^{f}}{\tau_{n}}}+\frac{\sqrt{b_{mn}c_{m}^{B}}}{\tau_{n}\sqrt{log_{2}(1+\varsigma_{mn})}}$.
Proof.
First, we transform Problem 1 into an unconstrained optimization problem using the Lagrangian dual method. The Lagrangian function associated with MU $m$ is formulated as
$L_{m}\left(\bm{f}_{m},\bm{B}_{m},\bm{p}_{m},\lambda_{m},\beta_{m},\delta_{mn},\gamma_{m}\right)=\sum_{n}p_{mn}V_{mn}$ (20)
$-\sum_{n}\left[c_{m}^{f}f_{mn}log(1/\theta_{m})+c_{m}^{B}B_{mn}\right]$
$-\lambda_{m}(\sum_{n}f_{mn}-f_{m}^{max})-\beta_{m}(\sum_{n}B_{mn}-B_{m}^{max})$
$-\sum_{n}\delta_{mn}[log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}log_{2}(1+\varsigma_{mn})}-\tau_{n}]$
$-\gamma_{m}[S_{m}-T^{req}(f_{m}^{max}-\sum_{n}f_{mn})],$
where $\lambda_{m}\geq 0,\beta_{m}\geq 0$, $\delta_{mn}\geq 0$ and $\gamma_{m}\geq 0$ are the Lagrangian dual variables corresponding to constraints C1, C2 and C3, respectively. Then, the first-order derivative of $L_{m}$ with respect to $f_{mn}$ and $B_{mn}$ can be derived as
$\frac{\partial L_{m}}{\partial f_{mn}}$ $=p_{mn}I_{mn}\frac{x_{m}\tau_{n}log(1/\theta_{m})}{f_{mn}^{2}}-c_{m}^{f}log(1/\theta_{m})$ (21)
$-\lambda_{m}+\delta_{mn}log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}^{2}}-\gamma_{m}T^{req},$
and
$\frac{\partial L_{m}}{\partial B_{mn}}$ $=p_{mn}I_{mn}\frac{b_{mn}}{B_{mn}^{2}log_{2}(1+\varsigma_{mn})}-c_{m}^{B}-\beta_{m}$ (22)
$+\delta_{mn}\frac{b_{mn}}{B_{mn}^{2}log_{2}(1+\varsigma_{mn})}.$
From (21) and the constraints in (11), the KKT conditions are given by
$\frac{\partial L_{m}}{\partial f_{mn}}=0,\,\,\lambda_{m}(\sum_{n}f_{mn}-f_{m}^{max})=0,$ (23)
$\delta_{mn}[log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}log_{2}(1+\varsigma_{mn})}-\tau_{n}]=0,$
$\gamma_{m}[S_{m}-T^{req}(f_{m}^{max}-\sum_{n}f_{mn})]=0,$
$\lambda_{m},\delta_{mn},\gamma_{m}\geq 0,C1,C2,C3.$
Since $S_{m}<T^{req}\left(f_{m}^{max}-\sum_{n}f_{mn}\right)$ and $\sum_{n}f_{mn}<f_{m}^{max}$, we can obtain $\lambda_{m}=0$ and $\gamma_{m}=0$. Furthermore, $\frac{\partial L_{m}}{\partial f_{mn}}=0$ can be converted to
$(p_{mn}I_{mn}+\delta_{mn})\frac{x_{m}\tau_{n}}{f_{mn}^{2}}-c_{m}^{f}=0.$ (24)
Similarly, the KKT conditions for $B_{mn}$ are obtained as
$\frac{\partial L_{m}}{\partial B_{mn}}=0,\,\,\beta_{m}(\sum_{n}B_{mn}-B_{m}^{max})=0,$ (25)
$\delta_{mn}[log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}log_{2}(1+\varsigma_{mn})}-\tau_{n}]=0,$
$\beta_{m},\delta_{mn}\geq 0,C1,C2.$
Since $\sum_{n}B_{mn}<B_{m}^{max}$, we can obtain $\beta_{m}=0$. Furthermore, $\frac{\partial L_{m}}{\partial B_{mn}}=0$ can be converted to
$(p_{mn}I_{mn}+\delta_{mn})\frac{b_{mn}}{B_{mn}^{2}log_{2}(1+\varsigma_{mn})}-c_{m}^{B}=0.$ (26)
Then, based on the value of $\delta_{mn}$, the optimal allocation of computation and communication resources exists in the following two cases:
  • ($\delta_{mn}=0$): According to (24) and (26), we have
    $f_{mn}=\sqrt{\frac{p_{mn}I_{mn}x_{m}\tau_{n}}{c_{m}^{f}}},B_{mn}=\sqrt{\frac{p_{mn}I_{mn}b_{mn}}{c_{m}^{B}log_{2}(1+\varsigma_{mn})}}.$ (27)
  • ($\delta_{mn}>0$): Substituting $f_{mn}$ and $B_{mn}$ into $\delta_{mn}\left[log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}log_{2}(1+\varsigma_{mn})}-\tau_{n}\right]=0$, we can get
    $\delta_{mn}=F^{2}-p_{mn}I_{mn},$ (28)
    $F=log(1/\theta_{m})\sqrt{\frac{x_{m}c_{m}^{f}}{\tau_{n}}}+\frac{\sqrt{b_{mn}c_{m}^{B}}}{\tau_{n}\sqrt{log_{2}(1+\varsigma_{mn})}}.$
    Next, bring $\delta_{mn}$ into $f_{mn}$ and $B_{mn}$, we have
    $f_{mn}=F\sqrt{\frac{x_{m}\tau_{n}}{c_{m}^{f}}},B_{mn}=F\sqrt{\frac{b_{mn}}{c_{m}^{B}log_{2}(1+\varsigma_{mn})}}.$ (29)
According to Proposition 1, the optimal strategy for each MU $m$ is influenced by five main factors: $p_{mn}$, $I_{mn}$, $c_{m}^{f}$, $c_{m}^{B}$ and $\tau_{n}$. Specifically, MU $m$ focuses on the rewards paid by MSPs, and the MU tends to put more computation and communication resources into the contribution when $p$ increases. Moreover, MUs are more willing to invest more resources when the potential contribution prediction $I_{mn}$ is higher, resulting in greater benefits. In addition, the virtual deadline $\tau_{n}$ set by the MSP $n$ determines the minimum criteria for resource allocation.
V-B Upper Level: Optimal Reward Decision Equilibrium Analysis
Given the responses $(\bm{f},\bm{B})$ from all MUs, and other MSPs’ decisions $\bm{p}_{\textrm{-}n}$, the optimal reward decision $\bm{p}_{n}$ of MSP $n$ can be obtained by solving Problem 2, defined as follows.
Proposition 2. Given $\bm{p}_{\textrm{-}n}$, MSP $n$’s best response strategy is
$\bm{p}_{n}^{*}=arg\max_{p_{mn}>0}\Psi_{n}\left(\bm{f}_{n}^{*}(\bm{p}_{n},\bm{p}_{-n}),\bm{B}_{n}^{*}(\bm{p}_{n},\bm{p}_{-n}),\bm{p}_{-n}\right).$ (30)
Then, we analyze the existence and uniqueness of the optimal reward decision equilibrium in Theorem 2.
Theorem 2.
There exists a unique Nash equilibrium (NE) [32] in the multi-MSP rewarding game $\Omega$.
Proof.
We define the Hessian matrix of $\Psi_{n}$ with respect to $\bm{p}_{n}$ as $(\bm{\Lambda}_{n}+\bm{H}_{n})$. The matrix $\bm{\Lambda}_{n}=diag\left(\frac{\partial^{2}\Psi_{n}}{\partial p_{1n}^{2}},\ldots,\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}\right)$ and the second-order partial derivative matrix $\bm{H}_{n}$ is expressed by
$\bm{H}_{n}=\begin{bmatrix}0&\frac{\partial^{2}\Psi_{n}}{\partial p_{1n}\partial p_{2n}}&\cdots&\frac{\partial^{2}\Psi_{n}}{\partial p_{1n}\partial p_{mn}}\\ \frac{\partial^{2}\Psi_{n}}{\partial p_{2n}\partial p_{1n}}&0&\cdots&\frac{\partial^{2}\Psi_{n}}{\partial p_{2n}\partial p_{mn}}\\ \vdots&\vdots&\ddots&\vdots\\ \frac{\partial^{2}\Psi_{n}}{\partial p_{mn}\partial p_{1n}}&\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}\partial p_{2n}}&\cdots&0\end{bmatrix},$ (31)
where
$\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}$ $=\mu_{n}\frac{V_{mn}^{\prime\prime}(1+\sum_{m=1}^{M}V_{mn})-(V_{mn}^{\prime})^{2}}{(1+\sum_{m=1}^{M}V_{mn})^{2}}$ (32)
$-2V_{mn}^{\prime}-p_{mn}V_{mn}^{\prime\prime},\forall m\in\mathcal{M},$
$\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}\partial p_{m^{\prime}n}}=$ $-\mu_{n}\frac{V_{mn}^{\prime}V_{m^{\prime}n}^{\prime}}{(1+\sum_{m=1}^{M}V_{mn})^{2}},$ (33)
$\forall m\in\mathcal{M},m\neq m^{\prime}.$
To obtain the values of $V_{mn}^{\prime}$ and $V_{mn}^{\prime\prime}$, we need to obtain $\frac{\partial f_{mn}}{\partial p_{mn}}$, $\frac{\partial^{2}f_{mn}}{\partial p_{mn}^{2}}$, $\frac{\partial B_{mn}}{\partial p_{mn}}$ and $\frac{\partial^{2}B_{mn}}{\partial p_{mn}^{2}}$. From (24), we can obtain $\frac{\partial f_{mn}}{\partial p_{mn}}$ and $\frac{\partial^{2}f_{mn}}{\partial p_{mn}^{2}}$, the steps of which are shown as follows. The first-order derivative of $\frac{\partial L_{m}}{\partial f_{mn}}=0$ with respect to $p_{mn}$ is expressed as $I_{mn}\frac{x_{m}\tau_{n}}{f_{mn}^{2}}-2(p_{mn}I_{mn}+\delta_{mn})\frac{x_{m}\tau_{n}}{f_{mn}^{3}}\cdot\frac{\partial f_{mn}}{\partial p_{mn}}=0$. Consequently, we can obtain
$\frac{\partial f_{mn}}{\partial p_{mn}}=\frac{I_{mn}f_{mn}}{2(p_{mn}I_{mn}+\delta_{mn})}>0.$ (34)
The second-order derivative of $\frac{\partial L_{m}}{\partial f_{mn}}=0$ with respect to $p_{mn}$ is expressed as
$\frac{\partial^{2}f_{mn}}{\partial p_{mn}^{2}}=\frac{I_{mn}\frac{\partial f_{mn}}{\partial p_{mn}}(p_{mn}I_{mn}+\delta_{mn})-I_{mn}^{2}f_{mn}}{2(p_{mn}I_{mn}+\delta_{mn})^{2}},$ (35)
by substituting (34) into (35), we have
$\frac{\partial^{2}f_{mn}}{\partial p_{mn}^{2}}=\frac{-I_{mn}^{2}f_{mn}}{4(p_{mn}I_{mn}+\delta_{mn})^{2}}<0.$ (36)
Likewise, we derive $\frac{\partial B_{mn}}{\partial p_{mn}}$ and $\frac{\partial^{2}B_{mn}}{\partial p_{mn}^{2}}$ with the following steps based on (26). The first-order derivative of $\frac{\partial L_{m}}{\partial B_{mn}}=0$ with respect to $p_{mn}$ is expressed as $I_{mn}\frac{b_{mn}}{B_{mn}^{2}log_{2}(1+\varsigma_{mn})}-2(p_{mn}I_{mn}+\delta_{mn})\frac{b_{mn}}{B_{mn}^{3}log_{2}(1+\varsigma_{mn})}\cdot\frac{\partial B_{mn}}{\partial p_{mn}}=0$. Accordingly, we have
$\frac{\partial B_{mn}}{\partial p_{mn}}=\frac{I_{mn}B_{mn}}{2(p_{mn}I_{mn}+\delta_{mn})}>0.$ (37)
Similarly, we obtain the second-order derivative of $\frac{\partial L_{m}}{\partial B_{mn}}=0$ with respect to $p_{mn}$ as
$\frac{\partial^{2}B_{mn}}{\partial p_{mn}^{2}}=\frac{I_{mn}\frac{\partial B_{mn}}{\partial p_{mn}}(p_{mn}I_{mn}+\delta_{mn})-I_{mn}^{2}B_{mn}}{2(p_{mn}I_{mn}+\delta_{mn})^{2}},$ (38)
by substituting (37) into (38), we have
$\frac{\partial^{2}B_{mn}}{\partial p_{mn}^{2}}=\frac{-I_{mn}^{2}B_{mn}}{4(p_{mn}I_{mn}+\delta_{mn})^{2}}<0.$ (39)
Based on (34), (36), (37), and (39), we can easily infer that $V_{mn}^{\prime}>0$ and $V_{mn}^{\prime\prime}<0$, as shown in (40). Then, we can obtain $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}\partial p_{m^{\prime}n}}<0$. For $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}$, We substitute the specific expressions for $V_{mn}^{\prime}$ and $V_{mn}^{\prime\prime}$ into (32) and obtain the following result as shown in (41). Next, we randomly choose a vector $\bm{h}\in\mathbb{R}^{M\times 1}$ with elements not all 00. For $\bm{\Lambda}_{n}$, it is easy to see that $\bm{h}^{T}\bm{\Lambda}_{n}\bm{h}=\sum_{m}\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}(h^{m})^{2}<0$. Furthermore, $\bm{h}^{T}\bm{H}_{n}\bm{h}=\frac{-\mu_{n}}{(1+\sum_{m}V_{mn})^{2}}\left(\sum_{m}h^{m}V_{mn}^{\prime}\right)^{2}<0$. Thus, we have $\bm{h}^{T}(\bm{\Lambda}_{n}+\bm{H_{n}})\bm{h}=\bm{h}^{T}\bm{\Lambda}_{n}\bm{h}+\bm{h}^{T}\bm{H}_{n}\bm{h}<0$, indicating that the utility function $\Psi_{n}$ is strictly concave. So far, the uniqueness of the Nash equilibrium solution to the multi-MSP rewarding game $\Omega$ has been proved.
$V_{mn}^{\prime}=\frac{\partial V_{mn}}{\partial p_{mn}}=I_{mn}\left(\frac{x_{m}\tau_{n}log(1/\theta_{m})}{f_{mn}^{2}}\cdot\frac{\partial f_{mn}}{\partial p_{mn}}+\frac{b_{mn}}{B_{mn}^{2}log_{2}(1+\varsigma_{mn})}\cdot\frac{\partial B_{mn}}{\partial p_{mn}}\right)>0,$ (40)
$V_{mn}^{\prime\prime}=\frac{\partial^{2}V_{mn}}{\partial p_{mn}^{2}}=I_{mn}\left[\frac{-2x_{m}\tau_{n}log(1/\theta_{m})}{f_{mn}^{3}}\cdot(\frac{\partial f_{mn}}{\partial p_{mn}})^{2}+\frac{x_{m}\tau_{n}log(1/\theta_{m})}{f_{mn}^{2}}\cdot\frac{\partial^{2}f_{mn}}{\partial p_{mn}^{2}}\right.$
$\left.-\frac{2b_{mn}}{B_{mn}^{3}log_{2}(1+\varsigma_{mn})}\cdot(\frac{\partial B_{mn}}{\partial p_{mn}})^{2}+\frac{b_{mn}}{B_{mn}^{2}log_{2}(1+\varsigma_{mn})}\cdot\frac{\partial^{2}B_{mn}}{\partial p_{mn}^{2}}\right]<0.$
$\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}=-\left[\frac{4a_{1}I_{mn}^{2}\delta_{mn}+a_{1}I_{mn}^{3}p_{mn}}{4f_{mn}(\delta_{mn}+I_{mn}p_{mn})^{2}}+\frac{4b_{mn}I_{mn}^{2}\delta_{mn}+b_{mn}I_{mn}^{3}p_{mn}}{4a_{2}B_{mn}(\delta_{mn}+I_{mn}p_{mn})^{2}}+I_{mn}^{2}\frac{\left(\frac{a_{1}a_{2}I_{mn}B_{mn}+b_{mn}I_{mn}f_{mn}}{2a_{2}f_{mn}B_{mn}(\delta_{mn}+I_{mn}p_{mn})}\right)^{2}}{(1+\sum_{m}V_{mn})^{2}}\right.$ (41)
$\left.+\frac{3I_{mn}^{3}(a_{1}a_{2}B_{mn}+b_{mn}f_{mn})}{4a_{2}f_{mn}B_{mn}(\delta_{mn}+I_{mn}p_{mn})^{2}(1+\sum_{m}V_{mn})}\right]<0,\text{where}\,a_{1}=x_{m}\tau_{n}log(1/\theta_{m}),a_{2}=log_{2}(1+\varsigma_{mn}).$
1
2Input: Action threshold $[p_{min},p_{max}]$, learning rate;
3 Output: The reward decisions $\bm{p}$ of MSPs, resource allocation $[\bm{f},\bm{B}]$ of MUs;
4 Initialization: local observation $s_{n}$, actor network $\bm{\alpha}_{n}$, critic network $\bm{\beta}_{n}$ and episode buffer $\mathcal{D}_{n}$ for each MSP agent $n$ ;
5 # Trading process using MDDR:
6 for  Episode = $1,2,\ldots$ do
7      for Epoch $k=1,\ldots,|\mathcal{D}|$ do
8            foreach MSP agent $n\in\mathcal{N}$ do
9                  Observe space $s_{n}(k)$;
10                   Choose price $\bm{p}_{n}(k)$ within $[p_{min},p_{max}]$ by sampling from $\pi_{\alpha_{n}}(s_{n}(k))$ ;
11             end foreach
12            All MSP agents do:
13             Perform a joint reward $\bm{p}(k)=\left\{\bm{p}_{1}(k),\ldots,\bm{p}_{n}(k)\right\}\in\mathcal{A}$;
14             Broadcast reward decisions;
15             Interact with the environment and receive responses from the MUs ;
16             Calculate utility according to (12) ;
17             Store $s_{n}(k)=[V_{n}(k),\bm{p}_{n}(k),\mathcal{U}_{n}(k),V_{n}(k+1)]$ ;
18       end for
19      Update actor network $\bm{\alpha}$ and critic network $\bm{\beta}$ for each MSP agent as [33] ;
20       Clear up relay buffer $\mathcal{D}$ ;
21 end for
22# FL training guided by trading results within $T$:
23 foreach MSP agent $n\in\mathcal{N}$  do
24      foreach MU $m\in\mathcal{M}$ do
25             Based on the trading results, utilize $f_{mn}$ and $B_{mn}$ to locally train the task of MSP $n$, and subsequently upload the model to the MSP $n$.
26       end foreach
27      Aggregate and issue an updated global model;
28       Supply the updated model to the AR engine;
29 end foreach
Algorithm 1 Entire Process of the Framework.
We verify the existence and uniqueness of the equilibrium solution for EPEC through Theorem 1 and Theorem 2. However, balancing multiple and conflicting interests, i.e., the utilities of MSPs and MUs, inevitably complicates the optimization problem. To find a solution that optimizes the profits of MSPs while taking into account the utilities of MUs, we develop a reinforcement learning-based strategy in the following section.
VI Dynamic Reward Strategy for MSPs
According to the EPEC proved above, it is possible to obtain an optimal solution. However, the optimal approach faces the following practical difficulties: (i) under the time-varying network, the optimal approach is time-consuming due to the complexity of (12), which is non-linear with complicated form and the reward decisions of MSPs are tightly coupled. (ii) MSPs must know private information about MUs (e.g., cost factors) as a prior condition for determining their rewards, which violates the privacy of MUs. To overcome the above difficulties, we further formulate the multi-MSP rewarding game as a multi-agent Markov decision process (MAMDP) [34] to adapt to dynamic channel conditions, where each MSP acts as an individual agent to make intelligent reward decisions.
VI-A MAMDP for Multi-MSP Rewarding Game $\Omega$
We train the reward model based on the state-of-the-art policy gradient method proximal policy optimization (PPO) for the reasons described in [35]. The $MAMDP=\left\langle\mathcal{S}_{n},\mathcal{A}_{n},\mathcal{P}_{n},\mathcal{U}_{n}\right\rangle$ for multi-MSP rewarding game is composed of state space $\mathcal{S}_{n}\triangleq\left\{s_{n}\right\}$, action space $\mathcal{A}_{n}\triangleq\left\{\bm{p}_{n}\right\}$, state transition probability $\mathcal{P}_{n}\triangleq\left\{P_{n}\right\}$, and utility $\mathcal{U}_{n}\triangleq\left\{\Psi_{n}\right\}$. For the MSP agent, the local observation contains the channel information and the responses of MUs, all of which are captured in the IoM of (2). Thus, we set the local observation of MSP $n$ at the $k$-th stage game defined as $s_{n}(k)=[V_{1n}(k),V_{2n}(k),\ldots,V_{mn}(k)]$, and the state of the environment is $\mathcal{S}(k)=[s_{1}(k),\ldots,s_{n}(k)]$. At the $k$-th stage game, MSP $n$ observes a state $s_{n}(k)$ and determines an action $\bm{p}_{n}(k)$ within $[p_{min},p_{max}]$. When an action $\bm{p}_{n}(k)$ is applied to state $s_{n}(k)$, the agent $n$ receives a utility $\mathcal{U}_{n}(k)$ from the environment. Taking into account the competition among agents, we define the reward $\mathcal{U}_{n}$ for each MSP by (12). The state transition probability $P(s_{n}(k+1)\mid s_{n}(k),\bm{p}_{n}(k))$ leads to the new state $s_{n}(k+1)$ after executing an action $\bm{p}_{n}(k)$ at the state $s_{n}(k)$.
Figure 7: The utilities of MSPs.
(a) The utilities of MUs
(b) The reward decisions of MSP 1
(c) The reward decisions of MSP 2
(d) The reward decisions of MSP 3
Figure 8: The convergence processes of MDDR under dynamic networks.
VI-B The Multi-agent DRL-based Dynamic Reward (MDDR)
In our work, each MSP agent operates the DRL-based dynamic reward algorithm in a fully distributed manner. Algorithm 1 shows the pseudocode for model trading using MDDR under a dynamic environment (Lines 5-19) and the guidance for the FL process (Lines 22-27).
Trading process: When the game starts, each MSP $n$ initializes its actor $\bm{\alpha}_{n}$ and critic network $\bm{\beta}_{n}$, essential parameters and inference loss for trading (Line 3). During each game, each agent feeds the observed information into the policy network $\bm{\alpha}_{n}$ to derive its reward policy (Lines 8-9). Then, all the agents generate reward decisions and broadcast them to the MUs (Lines 12-13). After completing these steps, the MSP agents enter the interaction phase with the environment, receive the best feedbacks from the MUs, calculate their utilities, and store information into episode buffer $\mathcal{D}$ (Lines 14-16). The experience is collected through the above interactive process (Lines 8-16) until the buffer $\mathcal{D}$ is full. Then, the critic-network $\bm{\beta}_{n}$ evaluates the actor-network $\bm{\alpha}_{n}$ based on the experience and in turn adjusts it (Line 18). The game will terminate when the maximum number of training episodes is reached and then the trading results will be obtained.
FL training guided by trading results: After obtaining the trading results (MSPs’ reward decisions $\bm{p}$ and resource allocation $(\bm{f},\bm{B})$), MUs and MSPs enter the FL phase. MUs follow the resource allocation scheme for local training and model uploading (Line 24). Subsequently, the MSP is responsible for aggregating and updating the global model (Line 26). Finally, the updated model is provided for use by the AR engine (Line 27). This process iterates multiple times until the total FL time $T$ is reached. During real-time interactions, the AR service of the MSP can be enhanced in terms of immersion and service quality by the updated model.
VI-C The Complexity and Convergence of MDDR
The complexity of MDDR is mainly influenced by the structure of the actor-critic network. According to [36], the total computational complexity of the fully connected layer can be expressed as the number of multiplications $\mathcal{O}(\sum_{u=1}^{U}\varrho_{u}\cdot\varrho_{u-1})$, where $\varrho_{u}$ is the number of neurons in the $u$-th hidden layer. In this work, each agent has its actor-critic network, structured as $[5,256,10]$ and $[5,256,1]$, respectively. Therefore, its complexity can be expressed as $\mathcal{O}(5\times 256+256\times 10+5\times 256+256\times 1)$ without burden on the MSP. After several attempts, the MDDR can achieve convergence [37] (as per Fig. 7 and Table II) with a suitable set of hyper-parameters.
TABLE II: The Utilities of three reward approaches
MSP 1 MSP 2 MSP 3 Total
Optimal 2169.99 2564.56 3035.96 7770.51
MDDR 2169.11 2563.95 3034.94 7768.00
MAPPO 2169.23 2563.96 3035.26 7768.45
(a) MSP1
(b) MSP2
(c) MSP3
(d) The IoM for MSP 1
(e) The IoM for MSP 2
(f) The IoM for MSP 3
Figure 9: Evaluation results for MNIST.
(a) MSP1
(b) MSP2
(c) MSP3
(d) The IoM for MSP 1
(e) The IoM for MSP 2
(f) The IoM for MSP 3
Figure 10: Evaluation results for GTSRB.
VII Performance Evaluation
First, we verify the near-optimal performance of MDDR. Then, to evaluate the effectiveness of the proposed framework, we compare its performance with benchmark schemes in object detection and classification for AR services through FL.
VII-A Simulation Settings
[…]
VII-B Benchmark Schemes
To our knowledge, there are no comparable solutions to the model trading problem in the vehicular metaverse. Therefore, under the uniform rewards of MSPs, we develop and extend four benchmark schemes for multi-MSP and multi-MU scenarios while meeting FL deadlines and resource constraints of devices.
  • x_based: Inspired by the incentive mechanism of the Stackelberg game for federated learning [25], the MU allocates computation and communication resources based on the amount of data involved in local training.
  • w_based: The MU adjusts the resource allocation according to the inference loss [42], i.e., the potential value of local data in our work.
  • w_x_based: Combining the x_based and w_based designs, the MU allocates resources considering both the data size and the potential value of data.
  • fixed: Taking into account the differences in performance of the various devices that exist in reality, we randomly allocate a fixed set of computation and communication resources to reflect the heterogeneity of devices.
TABLE III: Time comparison of five schemes for achieving the same performance.
Schemes MNIST GTSRB
90% Accuracy 50% mAP 70% mAP50
MSP 1 MSP 2 MSP 3 MSP 1 MSP 2 MSP 3 MSP 1 MSP 2 MSP 3
immersion-aware 15.23s 13.09s 7.59s 513.16s 581.09s 623.89s 418.28s 447.69s 451.27s
x_based 17.79s 18.03s 19.80s 812.53s 898.59s 1061.50s 596.57s 679.91s 890.47s
w_based 30s+ 24.04s 10.28s 1064.83s 1175.58s 1200s+ 774.09s 953.50s 1013.89s
w_x_based 23.20s 19.08s 12.34s 901.61s 1190s 1194s 596.35s 848.68s 932.73s
fixed 30s+ 30s+ 30s+ 1200s+ 1200s+ 1200s+ 1200s+ 1200s+ 1200s+
  • The $1200$s+ and $30$s+ indicate that the guidance time for the current trading is exceeded.
VII-C Performance Evaluation of MDDR
First, we consider a relatively static communication environment (fixed channel conditions) to show the effectiveness of MDDR compared to the optimal solution and the general multi-agent depth approximation policy optimization (MAPPO) approach. We trained our proposed MDDR and MAPPO until convergence, and the results are shown in Fig. 7 and Table II. We find that both MDDR and MAPPO achieve good results, but MAPPO slightly outperforms MDDR. This is because MAPPO is a centrally controlled distributed execution [33] that can utilize more information for decision making. However, this approach is not suitable for non-cooperation due to the need to share parameters among MSPs. For the fully distributed MDDR, it is more suitable for non-cooperative scenarios. The reason is that in MDDR, each agent has a self-contained actor-critic network, and information is not required to be shared. This suggests that the approach can achieve considerable utility while preserving privacy and saving communication costs significantly. Then, we deploy the MDDR in dynamic communication environments where channel conditions change over time. The behaviors of each MU and MSP in the trading are shown in Figs. 8(a), 8(b), 8(c) and 8(d), reflecting that the MDDR can converge rapidly and show good stability in dynamic network environments.
VII-D Benchmark Comparison
Figs. 9(a), 9(b), and 9(c) show the performance improvement in classification accuracy over time, where the immersion-aware outperforms the other schemes. This is because immersion-aware utilizes IoM to incentivize MUs to allocate computation and communication resources for providing learning models to MSPs. The IoM metric not only evaluates the freshness and accuracy of the model but also considers the amount and potential value of raw data used for training. As a result, resource-limited MUs are motivated to contribute more resources, providing valuable learning models to MSPs at an accelerated pace. In contrast, baseline schemes like x_based, w_based and w_x_based struggle to offer adequate incentives for MUs to contribute their available resources due to uniform rewards. For example, if a given MU has already reached the maximum resource it can contribute when the reward is $p$, then even if $p$ is increased further, the MSP’s gains will be impaired rather than increased. Consequently, the uniform reward $p$ will not be increased, leading to a situation where other MUs with available resources will not contribute more resources for providing higher-value learning models.
In addition, Figs. 9(a), 9(b), and 9(c) reveal that MSP 3 performs better than MSP 1 and MSP 2. For example, at the $10$th second, the accuracy of MSP 3 is about $88\%$, which is $14\%$ and $3\%$ higher than those of MSP 1 and MSP 2, respectively. This is because the potential value $w$ is higher for MSP 3, leading to more rewards being given by MSP 3. As a result, the MUs will allocate more computation and communication resources to provide learning models for MSP 3 than for MSP 1 and MSP 2. For the three different MSPs, some fluctuations are normal for model training (e.g., at time $10$s). It can be seen that the effect of fixed is the worst due to synchronous FL, which is limited by the MU with the least resource. Meanwhile, it is clear from Figs. 9(d)9(e) and 9(f) that the nature of the effectiveness of the five schemes is influenced by IoM. With the higher IoM, the performance is improved faster, resulting in a better immersive experience. Yet, the IoM of w_based in Fig. 9(e) is higher than x_based and w_x_based but the performance improvement rate is not as fast as them. The reason is that the IoM here represents the sum of all IoM from MUs, whereas the rate of performance improvement is hampered by the worst MU. Additionally,from Figs. 9(b) and 9(c), it can be observed that both w_x_based and w_based perform better when compared to Fig. 9(a). The main reason is the higher potential value $w$ associated with MSP 2 and MSP 3.
For object detection on GTSRB, Figs. 10(a), 10(b) and 10(c) show the mean average precision (mAP) and Table III gives the mAP50 (The value of mAP when the intersection over union (IOU) threshold is larger than $0.5$) results. We observe that compared with MNIST, the performance enhancement for object detection is slower due to the complexity of the detection model. Specifically, it takes at least $513.16$ seconds to achieve a $50\%$ mAP using the immersion-aware method. Our research introduces the flexibility to adjust decision guidance time, making it adaptable to various task types. Meanwhile, Figs. 10(d)10(e)10(f) and Table III demonstrate that MSP 1 with greater IoM takes less time to achieve the same mAP, which indicates that IoM works well to capture immersion. For instance, MSP 1 has the highest IoM, and the time that it takes to reach $50\%$ mAP is $513.16$s, which is $67.93$s and $110.73$s less than those of MSP 2 and MSP 3, respectively. The above comparison validates the feasibility of the proposed trading framework and the advantages of IoM-based metric incentives.
VIII Conclusion
To facilitate the development of vehicular metaverse services, we propose an immersion-aware model trading framework that incorporates FL to incentivize MUs to become contributors and provide locally trained data for AR services, i.e., object detection and classification. Specifically, we construct an EPEC problem with MSPs as leaders and MUs as followers to achieve an equilibrium among their interests. A new metric called “IoM” is designed to comprehensively evaluate the enhancement brought by the learning models of MUs for AR services. In addition, considering the competitive relationship among MSPs and the dynamic network environment, we develop a fully distributed MDDR approach to obtain the reward decisions of MSPs. Through extensive simulations using AR-related vehicle and MNIST datasets, our results demonstrate that the proposed framework can efficiently provide MUs with better immersive experiences for AR services in the vehicular metaverse.

### v2

@@
 Abstract
-Updates of extensive Internet of Things (IoT) data are critical to the immersion of vehicular metaverse services.
-However, providing high-quality and sustainable data in unstable and resource-constrained vehicular networks remains a significant challenge.
-To address this problem, we put forth a novel immersion-aware model trading framework that incentivizes metaverse users (MUs) to contribute learning models trained by their latest local data for augmented reality (AR) services in the vehicular metaverse, while preserving their privacy through federated learning.
-To comprehensively evaluate the contribution of locally trained learning models provided by MUs to AR services, we design a new immersion metric that captures service immersion by considering the freshness and accuracy of learning models, as well as the amount and potential value of raw data used for training.
-We model the trading interactions between metaverse service providers (MSPs) and MUs as an equilibrium problem with equilibrium constraints (EPEC) to analyze and balance their costs and gains.
-Moreover, considering dynamic network conditions and privacy concerns, we formulate the reward decisions of MSPs as a multi-agent Markov decision process.
-Then, a fully distributed dynamic reward method based on deep reinforcement learning is presented, which operates without any private information about MUs and other MSPs.
-Experimental results demonstrate that the proposed framework can effectively provide higher-value models for object detection and classification in AR services on real AR-related vehicle datasets compared to benchmark schemes.
+Timely updating of Internet of Things (IoT) data is crucial for immersive vehicular metaverse services.
+However, challenges such as latency caused by massive data transmissions, privacy risks associated with user data, and computational burdens on metaverse service providers (MSPs) hinder continuous collection of high-quality data.
+To address these issues, we propose an immersion-aware model trading framework that facilitates data provision for services while ensuring privacy through federated learning (FL).
+Specifically, we first develop a novel multi-dimensional metric, the immersion of model (IoM), which assesses model value comprehensively by considering freshness and accuracy of learning models, as well as the amount and potential value of raw data used for training.
+Then, we design an incentive mechanism to incentivize metaverse users (MUs) to contribute high-value models under resource constraints.
+The trading interactions between MSPs and MUs are modeled as an equilibrium problem with equilibrium constraints (EPEC) to analyze and balance their costs and gains, where MSPs as leaders determine rewards, while MUs as followers optimize resource allocation.
+Furthermore, considering dynamic network conditions and privacy concerns, we formulate the reward decisions of MSPs as a multi-agent Markov decision process.
+To solve this, we develop a fully distributed dynamic reward algorithm based on deep reinforcement learning, without accessing any private information about MUs and other MSPs.
+Experimental results demonstrate that the proposed framework outperforms state-of-the-art benchmarks, achieving improvements in IoM of $38.3\%$ and $37.2\%$, and reductions in training time to reach the target accuracy of $43.5\%$ and $49.8\%$, on average, for the MNIST and GTSRB datasets, respectively.
 Index Terms:
@@
 I Introduction
-The metaverse, envisioned as one of the next evolution directions of the Internet, is designed to create a fully immersive, self-sustaining virtual environment in which people can engage in activities such as playing, working, and socializing [1].
-This vision is propelled by advances in the fifth-generation (5G) and the forthcoming sixth-generation (6G) communication technologies, renowned for their low latency and high data throughput.
+The metaverse, envisioned as a major evolutionary step for the Internet, aims to create a fully immersive, self-sustaining virtual environment for activities such as playing, working, and socializing [1].
+This vision is propelled by advances in fifth- and sixth-generation (5G/6G) communication technologies, which offer low latency and high data throughput.
 These technologies play a critical role in seamlessly integrating the Internet of Things (IoT) data into metaverse services, thus bringing the once-fictional concept of immersive experiences closer to reality.
+Figure 1: An example of AR services in the vehicular metaverse [^2] Image source: https://www.jasoren.com/ar-in-automotive/.
 Metaverse services are beginning to reveal their vast potential across a broad spectrum of industries, from gaming and autonomous driving to education and marketing.
-Notably, the application of vehicles within the metaverse has captured significant interest due to the enhanced safety features and immersive experiences offered by state-of-the-art augmented reality (AR) technologies.
-Figure 1: An example of AR services in the vehicular metaverse [^2] Image source: https://www.jasoren.com/ar-in-automotive/.
+Notably, the application of vehicles within the metaverse [2, 3] has attracted significant interest, particularly for the enhanced traffic safety enabled by state-of-the-art augmented reality (AR) technologies.
 Market report [4] forecasted that the global automotive metaverse market will grow from $1.9$ billion in 2022 to $6.5$ billion by 2030.
-Automakers like BMW have recently doubled down on AR technology in their projects.
+Automakers such as BMW have increasingly invested in AR technology.
 As shown in Fig.
@@
 Moreover, Nissan’s [5] upcoming technology utilizes a 3D AR interface that merges the real world with the virtual world to provide the driver with augmented information about the surrounding area.
-In addition, Vanarama has also implemented a visual parking assistance function using AR [4], which is expected to identify empty spaces in busy car parks and display useful information, such as parking fees, on the windscreen.
+In addition, Neuron mobility [^3] https://immersive-technology.com/augmentedreality/neuron-introduces-new-ar-parking-assistant/ has announced the launch of an innovative AR parking assistance system with the intention of improving passengers’ parking routines and trip-end experiences.
 The ability to capture information from the “real” world, particularly the ability to collect and process massive data from IoT devices, is the key to determining the success of immersive services (e.g., AR) in the vehicular metaverse.
@@
 Motivations.
-An existing widely-used data collection method, as adopted by Nissan [6], involves collecting massive amounts of data through vehicle sensors, cameras, and roadside devices, and then processing all the data centrally.
-However, when it comes to the situation with multiple metaverse users (MUs) and metaverse service providers (MSPs) on behalf of different companies, the centralized data collection approach is not applicable and may lead to the following issues.
+The widely used data collection method, as adopted by Nissan [7] and studied in [8], involves gathering massive amounts of data through vehicle sensors, cameras, and roadside devices, and all data being processed centrally.
+While this approach has been effective for various applications, when it comes to the situation with multiple metaverse users (MUs) and metaverse service providers (MSPs) associated with different companies, the centralized data collection approach is not applicable and may lead to the following issues.
 First, AR services in the vehicular metaverse need to be highly immersive so that MUs feel fully immersed in the rendered environment, such as visualized driving.
-However, providing a seamless AR experience is challenging due to the latency caused by massive real-time data updates under unstable and resource-limited communication conditions.
+However, the data synchronization will be hindered by the latency from massive real-time data updates under unstable and resource-limited communication conditions.
 Note that the value of real-time data diminishes over time [9].
 Also, delays can severely impact the MU’s experience and cause dizziness [10].
-Second, the data to be transmitted may be sensitive and private, such as location, movement, and biometrics, which can create a better immersive experience but may inevitably increase the privacy risk of MUs [9].
-Third, it is difficult to share data between MSPs due to competition.
-This leads to the same data being sampled multiple times by different MSPs in the physical world, resulting in wasted resources (e.g., sampling costs) and data redundancy.
-Moreover, tremendous data pressure and computational burden are imposed on MSPs.
-Therefore, an efficient method for collecting high-quality data while preserving privacy is needed to achieve immersive AR services.
-Challenges.
-To address the above issues, we propose an immersion-aware model trading framework designed to provide data support for AR services and ensure privacy by integrating federated learning (FL).
-Adopting FL enables MUs to contribute diverse learning models to multiple MSPs using local data, thus effectively protecting sensitive information [10].
-Moreover, learning models are uploaded to MSPs for AR services by individual MUs, rather than centrally transmitting large amounts of data, which significantly reduces the communication burden.
-Nonetheless, developing such a model trading framework still faces the following challenges:
-  • MUs are typically self-interested and are reluctant to share learning models with MSPs due to the additional computation, communication, and energy overhead.
-Thus, the challenge here is how to incentivize them to become contributors in providing high-value learning models that will benefit vehicular metaverse services.
-  • It is difficult to design an appropriate metric to comprehensively evaluate the value of learning models contributed by MUs due to the diverse aspects involved in model valuation.
-  • MUs have different sampling costs and limited computation and communication resources, while MSPs differ in their model preferences and compete with each other.
-Hence, the next challenge is to model the dynamic competitive interactions among MSPs and to achieve an equilibrium of interests among MUs and MSPs in the model trading.
+Second, the data to be transmitted may be sensitive and private, such as location, movement, and biometrics, which can create a better immersive experience but may inevitably increase the privacy risk of MUs.
+Federated learning (FL) has been adopted in prior work [11, 12, 6, 13] to enable collaborative model training without sharing raw data (e.g., sensor/imaging data from vehicles).
+In this approach, learning models are uploaded by individual MUs to MSPs for AR services, rather than transmitting large volumes of data centrally, thus significantly reducing the communication burden.
+However, MUs are typically self-interested and are reluctant to share learning models with MSPs due to the additional computation, communication, and energy overhead.
+To overcome this, incentive mechanisms using strategies such as contract theory [14], Stackelberg game [15, 16] and multi-winner sealed-bid auction [17], have been proposed to encourage MUs to contribute models.
+However, existing studies fail to explicitly assess the model value from multiple dimensions, making it difficult for both MSPs and MUs to quantify the benefits of the model for MSPs.
+Furthermore, most of these solutions are designed for a single MSP and do not consider the joint optimization of MUs’ limited computational and communication resources.
+As a result, an efficient and privacy-preserving framework for data synchronization is needed to improve the immersive AR experience in the vehicular metaverse.
+To address the above research gaps, we propose an immersion-aware model trading framework for a multi-MSP, multi-MU vehicular metaverse.
+This framework is designed to incentivize MUs to become active contributors by providing models tailored to the specific needs of MSPs.
+However, MUs have different sampling costs and limited computational and communication resources, while MSPs differ in their model preferences and compete with each other.
+These inherent asymmetries and competitive dynamics make it challenging to establish an efficient model trading ecosystem.
+This leads to two key research questions that underpin our work:
+  • How can MUs be effectively incentivized to contribute high-value learning models that benefit vehicular metaverse services for diverse MSPs?
+  • How can the dynamic competitive interactions among MSPs be modeled to achieve a trade-off in gains between MUs and MSPs, ensuring sustainable and stable trading of learning models in the vehicular metaverse?
 Figure 2: The outline of an immersion-aware framework for FL-assisted vehicular metaverse.
-Contributions.
-To address these challenges, we model the dynamic competitive trading interactions as an equilibrium problem with equilibrium constraints (EPEC) involving multiple leaders (MSPs) and multiple followers (MUs), which is a hierarchical optimization problem with equilibria at two levels [11].
-Furthermore, to evaluate the value of learning models contributed by MUs to MSPs, we design a new metric called the immersion of the learning model (IoM).
-The metric jointly considers the freshness and accuracy of the learning model, as well as the amount and potential value of raw data used for training.
+Proposed Framework and Contributions.
+As depicted in Fig.
+2, our model trading framework consists of four components: FL mechanism, metric design, game modeling, and dynamic algorithms.
+Specifically, we first design a new metric called the immersion of the learning model (IoM) to evaluate the value of learning models contributed by MUs to MSPs.
+This metric jointly considers the freshness and accuracy of the learning model, as well as the amount and potential value of raw training data.
+Building on this metric, we propose an immersion-aware incentive mechanism that aligns the interests of both MUs and MSPs.
+Then, we model the dynamic competitive trading interactions as an equilibrium problem with equilibrium constraints (EPEC), which is a hierarchical optimization problem with equilibria at two levels [18].
 Moreover, given the dynamic networks and the privacy concerns of MSPs, we formulate the reward decisions of MSPs as a multi-agent Markov decision process (MAMDP) and develop a multi-agent deep reinforcement learning (DRL)-based dynamic reward (MDDR) approach to obtain the reward decisions in a fully distributed manner.
-The outline of the proposed framework is shown in Fig.
-2, and the main contributions of this paper are summarized as follows:
-  • We propose an immersion-aware trading framework for learning models with FL assistance, incentivizing MUs to contribute high-value learning models for MSPs in a privacy-preserving and efficient manner.
-Moreover, a novel metric called “IoM” is designed to quantify the immersion enhancement provided by MUs for AR services.
-  • We model the dynamic competitive trading interactions as an EPEC and provide theoretical support for the existence and uniqueness of equilibria at two levels.
-Moreover, a fully distributed MDDR approach is developed to adapt to dynamic environments and address complex reward decisions of MSPs without accessing any private information of MUs or MSPs.
+In summary, the key contributions of this work are as follows.
+  • Incentive mechanism design for immersion-aware model trading.
+From the perspectives of both MUs and MSPs, we propose an incentive mechanism that encourages MUs to contribute high-value learning models tailored to the specific demands of MSPs.
+To our knowledge, this is the first study focusing on incentive mechanism design for efficient and privacy-preserving data synchronization in multi-MU, multi-MSP vehicular metaverse environments.
+  • Novel design of multi-dimensional metric.
+To quantify immersion enhancement provided by MUs for AR services, we design an immersion metric of the learning model integrating four critical dimensions: the freshness and accuracy of the learning model, as well as the amount and potential value of raw training data.
+Freshness is captured through age of information (AoI), while potential value is evaluated by the difference between model predictions and true labels.
+This metric enables fine-grained evaluation of model value under resource constraints and provides a basis for incentive and decision-making strategies in model trading.
+  • Theoretical analysis and algorithm designs.
+Given multiple resource-constrained MUs and competing MSPs, we model their interactions as an equilibrium problem with equilibrium constraints (EPEC).
+We theoretically prove the existence and uniqueness of the equilibrium, and further develop a fully distributed MDDR approach that adapts to dynamic environments and operates without accessing any private information of MUs or MSPs.
+  • Performance evaluation.
 We conduct extensive numerical simulations based on AR-related vehicle datasets to validate the efficacy and efficiency of MDDR and the proposed immersion-aware model trading framework.
-Compared to benchmark schemes, the framework better motivates MUs to provide higher-value learning models for object detection and classification functions of AR services effectively.
+Numerical results show that our proposed mechanism improves the IoM by $38.3\%$ and $37.2\%$, and reduces the training time to reach the target accuracy by $43.5\%$ and $49.8\%$, on average, for the MNIST and GTSRB datasets, respectively, compared with state-of-the-art benchmarks.
 The rest of this paper is organized as follows.
@@
 II Related Work
-Incentive Mechanism.
-Incentive Mechanisms have been widely applied to metaverse services [12, 13, 14, 15, 16, 17].
-For example, Xu et al.
-[12] proposed an incentive framework for learning-based virtual reality (VR) services and designed a perceptual quality to measure VR users’ immersion in the virtual world.
-Sun et al.
-[13] investigated dynamic digital twins in air-assisted vehicular networking and designed a two-stage resource allocation incentive mechanism based on the Stackelberg game.
-Jiang et al.
-[14] developed a hierarchical game-theoretic coded distributed computing (CDC) framework for metaverse services.
-The framework employs coalition and Stackelberg games to select reliable CDC workers for MSPs and incentivize reliable CDC workers to participate in metaverse services.
-Xi et al.
-[16] proposed an incentive-based congestion control scheme for stochastic demand response in digital twin edge networks, using Lyapunov optimization and contract theory.
-Nguyen et al.
-[17] introduced MetaChain, an innovative blockchain-based framework designed to address the emerging challenges in developing metaverse applications.
-The framework integrates game theory to motivate MUs to contribute to the metaverse.
-However, the above incentives for metaverse services mainly focus on digital twins, blockchain, or virtual reality, while ignoring the scenarios of multiple MSPs and multiple MUs in the vehicular metaverse.
-In this scenario, the complexity of interactions among MSPs and MUs requires novel incentive mechanisms to better facilitate the development and innovation of metaverse services.
-AR and FL.
-The AR applications, characterized by smoothness and immersion requirements, provide immersive and interactive experiences for MUs by seamlessly integrating virtual content with the real world.
-For a better immersion experience, Cozzolino et al.
-[18] developed a latency-energy efficient task offloading for AR services and designed a best-effort resource provisioning algorithm to solve the problem of heterogeneous resource competition.
-Further, to reduce latency while protecting data privacy, Chen et al.
+In this section, we survey work related to our study in terms of edge-enabled vehicular metaverse services, FL for AR, and incentive mechanisms for data synchronization in the metaverse.
+II-A Edge-Enabled Vehicular Metaverse Services
+Vehicular metaverse services require ultra-low latency and high reliability to enable immersive experiences such as AR navigation and real-time digital twins.
+To meet these demands, researchers have extensively explored edge-based strategies for task offloading and resource management.
+For instance, Feng et al.
+[19] proposed a resource allocation framework for AR-powered vehicular edge metaverses that aims to maximize overall system utility.
+Similarly, Khan et al.
+[20] introduced a cooperative framework integrating task offloading, sensing, learning, and communication to reduce transmission energy consumption and latency in resource-constrained environments.
+Beyond static offloading, Chen et al.
+[21] developed a multi-agent deep reinforcement learning approach for dynamic avatar task migration.
+In terms of security and privacy, Kang et al.
+[22] presented a cross-metaverse empowered dual pseudonym management framework to mitigate privacy leakage risks during the dynamic communications among vehicular edge metaverses.
+Moreover, an optimization framework for AR-assisted driving was proposed in [23] to enhance the robustness of metaverse maps against adversarial attacks.
+This framework maximizes the mean average precision (mAP) of adversarial patch detection while minimizing AR scene uplink latency and worst-case battery consumption.
+Recent studies have also explored applying FL to vehicular metaverse services as a means of preserving user privacy during distributed model training.
+These efforts, particularly in AR-enhanced metaverse environments, are discussed in the next section.
+To ensure service continuity and prevent user disengagement during unexpected failures, Qiu et al.
+[24] proposed deploying redundant backup virtual vehicle services and keeping the age of backup information to improve the reliability of services.
+These works, from diverse perspectives, have significantly advanced the security, reliability, and performance of vehicular metaverse services, providing valuable insights for future research.
+Notably, distinct from existing studies, our work focuses on designing an incentive mechanism that transforms MUs into active contributors.
+These contributors efficiently provide high-value models to MSPs, thereby enhancing the overall service experience.
+II-B Federated Learning for AR in the Metaverse
+Privacy preservation is a critical concern in vehicular metaverse services, particularly when training models using sensitive user data.
+Federated learning (FL) has emerged as a promising technique to enable collaborative model training without sharing raw data.
+For example, Chen et al.
 [11] proposed a framework that integrates mobile edge computing and FL to solve the computational efficiency and low-latency object classification problem for AR services.
-In FL, MUs can train AI models on their local devices, thus enabling privacy-preserving collaborative machine learning while fully leveraging the computation resources of those MUs [20].
-Furthermore, Lu et al.
-[21] investigated communication-efficient FL for digital twin edge networks in industrial IoT, which provides virtual and real synchronization while protecting privacy.
+However, they did not address how communication and computational resources should be allocated within the FL framework.
+To overcome this limitation, Zhou et al.
+[12] incorporated FL into mobile AR systems in the metaverse to enable collaborative model training.
+Their approach considers the trade-offs among energy consumption, execution latency, and model accuracy to meet varying demands across different application scenarios.
+To further improve learning efficiency, an adaptive and resource-efficient FL algorithm tailored for AR applications was proposed in [6].
+This approach mitigates the effects of non-IID data distributions, reduces resource usage, and enhances the quality of experience.
+Additionally, Hazarika et al.
+[13] explored the integration of quantum computing with FL to provide a cost-efficient and adaptive solution for the dynamic nature of vehicular environments.
 Incorporating FL into AR services brings the benefits of reduced communication latency and privacy protection.
 However, the existing studies fail to consider the integration of FL into AR for the multi-MSP and multi-MU metaverse scenario, while overlooking the selfishness of MUs, i.e., whether they are willing to participate in FL learning.
-Inspired by the above existing works, we design an immersion-aware framework integrated FL to provide a learning model trading platform for MSPs and MUs in the vehicular metaverse service markets, enabling MUs to contribute high-value learning models for AR services.
-Figure 3: The workflow of the immersion-aware model trading framework [^4] Image source: https://mixed-news.com/en/google-transforms-google-maps-into-the-backbone-of-our-ar-future/, https://about.facebook.com/meta, https://www.vanarama.com/blog/cars/4-ways-augmented-reality-will-revolutionise-the-automotive-industry..
+II-C Incentive Mechanism for Data Synchronization in the Metaverse.
+The quality of metaverse services depends heavily on massive sensing data from the real world.
+To facilitate this, incentive mechanisms are widely employed to motivate MUs to contribute data and participate in model construction, thereby enhancing their service experience.
+For instance, Zhang et al.
+[25] proposed a vehicle-assisted data sensing framework that incentivizes vehicles to contribute sensing data to metaverse service, improving the driving experience for vehicle users.
+Similarly, Lin et al.
+[26] introduced a contract-based incentive design aimed at optimizing the long-term profits of digital twin service providers, addressing service congestion caused by stochastic demand responses.
+Xu et al.
+[27] developed a metaverse-based unmanned aerial vehicle (UAV) swarm system that uses digital twins and semantic communication to enhance data synchronization and reduce communication latency.
+Moreover, Lotfi et al.
+[28] introduced VMGuard, a novel four-layer security framework for the vehicular metaverse that protects against data poisoning attacks by using a reputation-based incentive mechanism.
+The framework leveraged user feedback and subjective logic modeling to assess the trustworthiness of devices.
+However, while these works successfully incentivize MUs to contribute raw data, they also pose a risk to user privacy due to potential data leakage.
+To address the privacy concern, Kang et al.
+[14] developed a cross-chain FL framework with an AoI-based contract theory model under prospect theory to incentivize data sharing.
+They utilized a hierarchical cross-chain architecture with a main chain and multiple subchains to perform decentralized, privacy-preserving, and secure data training in both virtual and physical spaces.
+Baccour et al.
+[15] proposed a dual game-theoretic framework for federated meta-learning (FML) in metaverse services.
+They introduced a reputation system and a Stackelberg game-based incentive mechanism to enhance data privacy, minimize energy costs, and improve FML efficacy, outperforming traditional clustering methods in training performance and metaverse utility.
+Li et al.
+[16] designed a satisfaction metric that considers data size, AoI, and training latency.
+They integrated this metric into utility functions to incentivize node participation in FL while optimizing resource allocation.
+However, these works are designed for a single MSP scenario, limiting scalability and applicability in multi-server scenarios commonly encountered in large-scale metaverse environments.
+Building on the concept of multiple MSPs, Zhang et al.
+[17] proposed a privacy-preserving auction mechanism for learning model as an NFT in Blockchain-driven Metaverse.
+They employed an imperfect information Stackelberg game to optimize the strategies of MUs and MSPs, and introduced a multi-winner sealed-bid auction mechanism to enhance the quality of FL-NFT.
+However, this approach primarily focuses on modeling NFTs and overlooks the joint optimization of computational and communication resources.
+Moreover, it fails to evaluate the model’s value from multiple dimensions.
+Inspired by the above existing works, this study focuses on multi-MSP and multi-MU scenarios in vehicular metaverse services.
+By assigning MUs the dual roles of service consumers and model contributors, we build a low-cost, privacy-preserving data support framework.
+This allows MUs to enjoy immersive in-vehicle services (e.g., AR navigation, digital twin interactions) while participating as distributed model contributors in service optimization.
+This framework continuously incentivizes MUs to contribute high-value learning models for metaverse services, thereby enhancing the overall service experience.
+Figure 3: Workflow of the immersion-aware model trading framework [^5] Image source: https://mixed-news.com/en/google-transforms-google-maps-into-the-backbone-of-our-ar-future/, https://about.facebook.com/meta, https://www.vanarama.com/blog/cars/4-ways-augmented-reality-will-revolutionise-the-automotive-industry..
 III System Model
 We first present the system overview of the immersion-aware model trading framework in Section III-A.
-We next introduce the FL mechanism adopted by our framework in Section III-B and design the immersion metric of the learning model in Section III-C.
+Next, we introduce the FL mechanism adopted by our framework in Section III-B and design the immersion metric of the learning model in Section III-C.
 III-A System Overview
-The workflow of the immersion-aware model trading framework is illustrated in Fig.
-4, aiming to provide MUs with immersive and interactive AR service experiences.
+Fig.
+5 illustrates the workflow of our immersion-aware model trading framework, designed to enhance the immersive and interactive AR experiences of MUs.
 MUs can contribute learning models to the AR services of MSPs for better immersive experiences.
@@
 The framework mainly consists of MSPs owned by different companies (e.g., Meta, Nissan, and Google), an infrastructure layer, and an interaction layer with the MUs that enjoy vehicular metaverse services.
-MSPs are fueled by many technologies that work together to support vehicular metaverse services, including AI engines, AR/VR engines, metaverse trading engines, etc.
+MSPs are powered by various technologies supporting vehicular metaverse services, such as AI engines, AR/VR engines, metaverse trading engines.
 The infrastructure layer with base stations (equipped with caches and edge servers) provides the basis of 5G/6G communication services for the interactions among MUs and MSPs.
 MUs play essential roles as both consumers and contributors to vehicular metaverse services.
-TABLE I: Parameter Definitions and Notations
+TABLE I: Definitions of Notation
 Notation Definition
-$M$ Number of MUs
-$N$ Number of MSPs
+$m,M$ Index of an MU, Number of MUs
+$n,N$ Index of an MSP, Number of MSPs
 $\overline{\Delta}_{mn}$ The average AoI of MU $m$’s model provided for MSP $n$
 $V_{mn}$ The IoM contributed by MU $m$ to MSP $n$
-$f_{mn}$ MU $m$’s computation resource used for MSP $n$
+$f_{mn}$ MU $m$’s computational resource used for MSP $n$
 $B_{mn}$ MU $m$’s communication resource used for MSP $n$
@@
 The framework is a multi-engine configuration, and this paper focuses on two components: the trading engine and the AI engine of the MSP.
-The trading engine is responsible for incentivizing MUs to participate in mutually beneficial collaborations with MSPs, providing them with a trading platform to make their decisions (i.e., computation and communication resource allocation of MUs and reward decisions of MSPs).
+The trading engine is responsible for incentivizing MUs to participate in mutually beneficial collaborations with MSPs, providing them with a trading platform to make their decisions (i.e., computational and communication resource allocation of MUs and reward decisions of MSPs).
 Then, each MU deciding to contribute utilizes the FL mechanisms integrated into AI engines to provide learning models for MSPs under the strategic guidance of the trading results.
@@
 Note that it is feasible to train on multiple tasks with the same local data.
-For example, for the same image, intuitive road safety focuses on pedestrians and vehicles, while intelligent terrain mapping focuses on landmarks and routes.
+For instance, for the same image, intuitive road safety focuses on pedestrians and vehicles, while intelligent terrain mapping focuses on landmarks and routes.
 (① in Fig.
-4).
+5).
 Then, the trading engines of MSPs determine their digital currency prices per IoM, i.e., rewards to MUs, and broadcast them to MUs (②).
-Next, MUs determine their allocation of computation and communication resources for contributing learning models to various MSPs based on the IoM, costs, and rewards given by MSPs.
-(③).
-Based on the responses from MUs, the trading engines of MSPs adjust the digital currency prices to maximize their benefits.
+Next, MUs determine their allocation of computational and communication resources for contributing learning models to various MSPs based on the IoM, costs, and rewards given by MSPs (③).
+Based on the responses from MUs, the trading engines of MSPs adjust the digital currency prices to maximize their own utilities.
 Steps ② and ③ are repeated until an agreement is reached among MUs and MSPs.
-Phase II (FL process): Each MU, guided by decisions obtained through tradings, allocates computation resources to local training based on different global models from MSPs and generates local learning models for different MSPs (④).
+Phase II (FL process): Each MU, guided by decisions obtained through tradings, allocates computational resources to local training based on different global models from MSPs and generates local learning models for different MSPs (④).
 Then, MUs upload their updated learning models to the corresponding AI engines of MSPs by consuming communication resources (⑤).
 Following this, AI engines aggregate models from all MUs to update their global models and transmit them to their AR/VR engines for utilization (⑥).
-Finally, the AR/VR engines enable better AR services with the assistance of the AI engines (⑦).
+Finally, the AR/VR engines provide enhanced AR services to MUs (⑦).
 Steps ④-⑦ are repeated until the guidance time $T$ ends.
@@
 Here, we consider a set of =$\mathcal{M}=\left\{1,\ldots,M\right\}$ MUs with the capacities of local computation and a set of $\mathcal{N}=\left\{1,\ldots,N\right\}$ MSPs with FL synchronous tasks [^7] That is, the update time of the global model is limited by the slowest MU, i.e., the AI engine of the MSP must wait to receive updated models from all MUs before performing model aggregation.
-The advantages of synchronous FL are high model accuracy and fast convergence [22, 23]..
+The advantages of synchronous FL are high model accuracy and fast convergence [29]..
 Each MSP initiates an FL synchronous task with a virtual deadline $\tau_{n}(n\in\mathcal{N})$.
@@
 Without loss of generality, we consider that the trading decisions among MUs and MSPs are made at the beginning of the period $T$.
-Figure 4: An illustration of the FL mechanism with $n$ tasks for MU $m$.
-To be concrete, let $X_{mn}$ be the set of data samples for task $n$ of MU $m$ as a set of input-output pairs $\sum_{i=1}^{|X_{mn}|}(y_{i},z_{i})$, where $y_{i}$ is an input sample vector with $d$ features, $z_{i}$ is the labeled output (i.e., ground truth) value of sample $y_{i}$, and $|X_{mn}|$ is the size of set $X_{mn}$.
+Figure 4: Illustration of the FL mechanism with $n$ tasks for MU $m$.
+In detail, let $X_{mn}$ be the set of data samples for task $n$ of MU $m$ as a set of input-output pairs $\sum_{i=1}^{|X_{mn}|}(y_{i},z_{i})$, where $y_{i}$ is an input sample vector with $d$ features, $z_{i}$ is the labeled output (i.e., ground truth) value of sample $y_{i}$, and $|X_{mn}|$ is the size of set $X_{mn}$.
 The data, such as images of streets and pedestrians, can be generated by high-definition cameras inside and outside the vehicles.
@@
 After receiving all the updated models, the MSP $n$’s AI engine aggregates them through the following weighted average:
-    $\jmath_{n}^{(r+1)}$ $=\jmath^{(r)}_{n}+\sum_{m=1}^{M}\frac{|X_{mn}^{(r)}|}{|X_{n}^{(r)}|}(\jmath_{n}^{(r)}-\jmath_{mn}^{(r)}),$ (1)
+    $\jmath_{n}^{(r+1)}$ $=\jmath^{(r)}_{n}+\sum_{m=1}^{M}\frac{|X_{mn}^{(r)}|}{|X_{n}^{(r)}|}(\jmath_{mn}^{(r)}-\jmath_{n}^{(r)}),$ (1)
     and then obtains the new global model $\jmath_{n}^{(r+1)}$ for the next iteration.
@@
 MSPs determine the rewards for MUs based on the immersion of the learning model (IoM) metric.
-Correspondingly, MUs need to decide the IoMs of the models provided to MSPs by allocating their computation and communication resources to maximize the benefits.
+Correspondingly, MUs need to decide the IoMs of the models provided to MSPs by allocating their computational and communication resources to maximize the benefits.
 This process highlights the key role of “IoM” in the interactions between the two parties, serving as both an evaluation criterion and a basis for decision-making.
@@
 It is more practical to comprehensively evaluate the amount of training data, model accuracy, and the potential value of local data; thus, we denote $I_{mn}$ as
-$I_{mn}=\frac{\omega_{mn}\epsilon ln(1+\eta\lfloor\frac{T}{\tau_{n}}\rfloor|X_{mn}|)}{\theta_{m}},$ (3)
+$I_{mn}=\frac{\omega_{mn}\epsilon\ln(1+\eta\lfloor\frac{T}{\tau_{n}}\rfloor|X_{mn}|)}{\theta_{m}},$ (3)
 where $\lfloor\frac{T}{\tau_{n}}\rfloor$ is the number of iterations that the task of MSP $n$ can be performed within $T$.
@@
 To obtain a comprehensive understanding of the overall freshness of model updates over a certain period, we use average AoI to evaluate the freshness of the models.
-With the aid of Figure 5, the average AoI of the learning model from MU $m$ to MSP $n$ can be expressed as
+With the aid of Figure 5, the time-averaged AoI of the learning model from MU $m$ to MSP $n$ can be expressed as
 […]
@@
 Since MU $m$ constantly collects data throughout the FL process, the amount of training data collected for training in each round is $|X_{mn}|=x_{m}\tau_{n}$, where $x_{m}$ denotes the data (number of floats) collected per unit time.
-The cumulative time $T_{mn}^{c}$ for local training is determined by the amount of training data, computation resource $f_{mn}$ allocated to MSP $n$, and its local accuracy threshold $\theta_{m}$, which is equivalent to
+The cumulative time $T_{mn}^{c}$ for local training is determined by the amount of training data, computational resource $f_{mn}$ allocated to MSP $n$, and its local accuracy threshold $\theta_{m}$, which is equivalent to
 $T_{mn}^{c}=\log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}.$ (7)
@@
 $T_{mn}^{t}=\frac{b_{mn}}{B_{mn}\log_{2}(1+\varsigma_{mn})},$ (8)
-where $\varsigma_{mn}=\frac{p^{t}_{m}g_{mn}}{\sigma^{2}}$ denotes the signal-to-interference-plus-noise ratio (SINR) for the communication between MU $m$ and MSP $n$.
+where $\varsigma_{mn}=\frac{p^{t}_{m}g_{mn}}{\sigma^{2}}$ denotes the signal-to-noise ratio (SNR) [34] for the communication between MU $m$ and MSP $n$.
 $g_{mn}$ and $p^{t}_{m}$ are the corresponding channel gain and the transmission power of MU $m$, respectively.
@@
 We define the utility of MU $m$ composed of the rewards from MSPs and the cost incurred for contribution, which is also affected by resource constraints and immersion requirements for consuming vehicular metaverse services.
-The cost incurred due to local training and model uploading, which can be expressed as
+The cost incurred due to local training and model uploading can be expressed as
 $C_{mn}=c_{m}^{f}\log(1/\theta_{m})f_{mn}+c_{m}^{B}B_{mn},$ (10)
-where $c_{m}^{f}$ and $c_{m}^{B}$ denote the cost factors of computation and communication resources, respectively.
-Let $\bm{f}_{m}\triangleq\left(f_{m1},f_{m2},\ldots,f_{mn}\right)$ and $\bm{B}_{m}\triangleq\left(B_{m1},B_{m2},\ldots,B_{mn}\right)$ be the computation and communication allocation profiles of MU $m$, respectively.
+where $c_{m}^{f}$ and $c_{m}^{B}$ denote the cost factors of computational and communication resources, respectively.
+Let $\bm{f}_{m}\triangleq\left(f_{m1},f_{m2},\ldots,f_{mn}\right)$ and $\bm{B}_{m}\triangleq\left(B_{m1},B_{m2},\ldots,B_{mn}\right)$ be the computational and communication allocation profiles of MU $m$, respectively.
 Then, the utility maximization problem for MU $m$ within $T$ can be formulated as
-Problem 1
+Problem 1.
 $\max\,\,\Phi_{m}(\bm{f}_{m},\bm{B}_{m})=\sum_{n=1}^{N}(p_{mn}V_{mn}-C_{mn}),$ (11)
@@
 $C3:\frac{S_{m}}{f^{max}_{m}-\sum_{n=1}^{N}f_{mn}}<T^{req},$
-where $S_{m}$ indicates the minimum computation resource required to enjoy other basic services.
-Constraints C1 and C3 indicate that the MU’s total computation and communication resources are limited and cannot be fully used to contribute learning models, given the MU’s demand for basic services.
+where $S_{m}$ indicates the minimum computational resource required to enjoy other basic services.
+Constraints C1 and C3 indicate that the MU’s total computational and communication resources are limited and cannot be fully used to contribute learning models, given the MU’s demand for basic services.
 Constraint C2 ensures that the total time for local training and uploading cannot exceed the time constraint $\tau_{n}$ for each round.
-The goal of each MU is to choose the optimal allocation of computation $\bm{f}_{m}$ and communication $\bm{B}_{m}$ resources to maximize its utility.
+The goal of each MU is to choose the optimal allocation of computational $\bm{f}_{m}$ and communication $\bm{B}_{m}$ resources to maximize its utility.
 Note that $V_{mn}$ must be greater than 00.
-This is because if $V_{mn}\leq 0$ and MU $m$ is involved in contributing learning models, then the utility is calculated as $p_{mn}V_{mn}-C_{mn}\leq 0$, resulting in MU $m$ will not participate in the trading.
+This is because if $V_{mn}\leq 0$ and MU $m$ is involved in contributing learning models, the utility is calculated as $p_{mn}V_{mn}-C_{mn}\leq 0$, resulting in MU $m$ choosing not to participate in the trading.
 IV-B Utility of MSP
 Each MSP has a gain function $\psi$ associated with the IoM to measure the benefits of all the learning models from MUs to the service.
-In this paper, we adopt $\psi_{n}=\mu_{n}ln(1+\sum_{m=1}^{M}V_{mn})$ as a monotonically increasing, differentiable, strictly concave IoM function, which is a simplification of the widely adopted function [28].
+In this paper, we adopt $\psi_{n}=\mu_{n}\ln(1+\sum_{m=1}^{M}V_{mn})$ as a monotonically increasing, differentiable, strictly concave IoM function, which is a simplification of the widely adopted function [35].
 Let $\bm{p}_{n}\triangleq\left(p_{1n},\ldots,p_{mn}\right)$ and $\bm{p}_{\textrm{-}n}$ represent the reward profile of MSP $n$ and all other MSPs except MSP $n$, respectively.
 Then, the optimization problem for the MSP is defined as
-Problem 2
+Problem 2.
 $\max\,\,\Psi_{n}(\bm{p}_{n},\bm{p}_{\textrm{-}n})=\psi_{n}-\sum_{m=1}^{M}p_{mn}V_{mn},$ (12)
@@
 Note that the other MSPs’ decisions $\bm{p}_{\textrm{-}n}$ are captured by the MUs’ responses.
-Moreover, when MSP $n$ determine its reward $\bm{p}_{n}$ for different MUs, the MSP needs to consider the rewards offered by other MSPs (i.e., $\bm{p}_{\textrm{-}n}$ ) as well as the strategies of all MUs (i.e., $(\bm{f}_{m},\bm{B}_{m}),m\in\mathcal{M}$).
+Moreover, when MSP $n$ determines its reward $\bm{p}_{n}$ for different MUs, the MSP needs to consider the rewards offered by other MSPs (i.e., $\bm{p}_{\textrm{-}n}$ ) as well as the strategies of all MUs (i.e., $(\bm{f}_{m},\bm{B}_{m}),m\in\mathcal{M}$).
 This thereby leads to reward competition among MSPs.
@@
   • Utilities: The vector $\bm{\Psi}=\left\{\Psi_{1},\Psi_{2},\ldots,\Psi_{n}\right\}$ contains the utility functions of all the MSPs defined in (12).
-Figure 6: The structure of EPEC.
+Figure 6: The hierarchical structure of equilibrium problem with equilibrium constraints (EPEC).
 IV-C Multi-MSP Multi-MU Game as EPEC
-In such a multi-MSP and multi-MU scenario, each MU has the choice of multiple MSPs to contribute learning models, and each MSP can incentivize multiple MUs to earn high-value learning models and achieve more benefits.
-Accordingly, each MSP competes with others for the equilibrium, which is constrained by the equilibrium among the MUs.
-This leads to the EPEC formulation [30].
 As shown in Fig.
-6, MUs and MSPs negotiate allocation strategies of computation and communication resources and reward decisions to maximize their benefits.
+6, MUs and MSPs negotiate allocation strategies of computational and communication resources and reward decisions to maximize their benefits.
 In the upper level, MSPs determine the rewards they are willing to offer by considering their cost, the responses of MUs, and the decisions of other MSPs.
-In the lower level, each MU m gives the optimal allocation response for computation and communication resources by considering its resource constraints, cost, and rewards from different MSPs.
+In the lower level, each MU $m$ gives the optimal allocation response for computational and communication resources by considering its resource constraints, cost, and rewards from different MSPs.
 The objective of EPEC is to find the equilibria at two levels, i.e., the point at which the MSPs’ (leaders’) utilities are maximized given that the MUs (followers) will choose their best responses.
@@
 Definition 2.
-Let $(f^{*}_{mn},B^{*}_{mn})$ and $p^{*}_{mn}$ denote the optimal computation and communication resource allocation of MU $m\in\mathcal{M}$ and the optimal reward decision of MSP $n\in\mathcal{N}$, respectively.
+Let $(f^{*}_{mn},B^{*}_{mn})$ and $p^{*}_{mn}$ denote the optimal computational and communication resource allocation of MU $m\in\mathcal{M}$ and the optimal reward decision of MSP $n\in\mathcal{N}$, respectively.
 Then, the points $(f^{*}_{mn},B^{*}_{mn})$ and $p^{*}_{mn}$ are the equilibria at two levels if the following conditions hold:
@@
 $\Psi_{n}\left(\bm{f}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{B}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{p}_{n}^{*},\bm{p}_{-n}^{*}\right)\geq$
-$\Psi_{n}\left(\bm{f}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{B}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{p}_{n},\bm{p}_{-n}^{*}\right),\forall n\in\mathcal{N}.$
+$\Psi_{n}\left(\bm{f}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{B}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{p}_{n},\bm{p}_{-n}^{*}\right),\forall n\in\mathcal{N},$
+where $\bm{p}_{-n}^{*}$ denotes the optimal rewards offered by other MSPs.
 In summary, the MSPs’ optimization problems are formulated as the following EPEC problems:
@@
 To investigate the above EPEC, we address the lower level (Problem 1) and the upper level (Problem 2) by using the backward induction methods in the following section.
-V Analysis and Solutions
-In this section, we use backward induction to investigate the existence and uniqueness of the equilibria at two levels, i.e., the optimal computation and communication resource allocation of MUs at the lower level and the optimal reward decisions of MSPs at the upper level.
-V-A Lower Level: Optimal Computation and Communication Resource Allocation Analysis
-For any reward decisions $\bm{p}$ given by MSPs, MU $m$ would like to determine an optimal decision $\bm{f}_{m}^{*}$ and $\bm{B}_{m}^{*}$ to maximize its utility.
-We analyze and provide the solution for the optimal allocation of computation and communication resources as follows.
+V EPEC Analysis and Solutions
+In this section, we utilize backward induction to analyze the EPEC formulated in (14).
+Specifically, we prove the existence and uniqueness of the equilibrium solutions at two hierarchical levels: (1) the optimal allocation of computational and communication resources for MUs (lower level), and (2) the optimal reward decisions for MSPs (upper level).
+V-A Lower Level: Optimal Resource Allocation for MUs
+In the lower level of EPEC, for any reward decisions $\bm{p}$ given by MSPs, MU $m$ aims to solve Problem 1 in (11) to determine its optimal computational and communication resource allocations, i.e., $\bm{f}_{m}^{*}$ and $\bm{B}_{m}^{*}$, to maximize its utility.
+Below, we analyze and derive the unique optimal allocation.
 Definition 3.
-If the problem is strictly concave for maximization and satisfies the following conditions, it will have a unique optimal solution [31]: (1)The Utility function is concave; (2) Constraints are all affine constraints; and (3) Strategy space is a compact and convex set.
+A maximization problem has a unique global optimum if the following conditions are satisfied [37]: (1) The objective function is strictly concave; (2) All constraints (equality and inequality) are affine; and (3) The strategy space is compact and convex.
 Theorem 1.
-Problem 1 has a unique optimal solution in the lower level, i.e., each MU has an optimal and unique allocation scheme for computation and communication resources.
+Problem 1 is strictly concave and has a unique globally optimal solution in the lower level.
+That is, for each MU $m$, there exists a unique resource allocation tuple $(\bm{f}_{m}^{*},\bm{B}_{m}^{*})$ that maximizes its utility.
 Proof.
-We first observe that the Hessian matrix of $\Phi_{m}(\bm{f}_{m},\bm{B}_{m})$ with respect to $(f_{mn},B_{mn})$ is
-$\bm{H}_{m}=\begin{bmatrix}\bm{H}_{m}^{f}&0\\ 0&\bm{H}_{m}^{B}\end{bmatrix},$ (15)
-where $\bm{H}^{f}_{m}$ can be obtained by the second-order derivation of $\Phi_{m}(\bm{f}_{m},\bm{B}_{m})$ with respect to $f_{mn}$
+We first examine the Hessian matrix of MU $m$’s utility $\Phi_{m}(\bm{f}_{m},\bm{B}_{m})$ with respect to $(f_{mn},B_{mn})$.
+Let this Hessian matrix be denoted by $\bm{H}_{m}$, which can be block-diagonalized as
+$\bm{H}_{m}=\begin{bmatrix}\bm{H}_{m}^{f}&0\\ 0&\bm{H}_{m}^{B}\end{bmatrix}.$ (15)
+The block matrix $\bm{H}^{f}_{m}$ can be computed as the second-order partial derivative of $\Phi_{m}(\bm{f}_{m},\bm{B}_{m})$ with respect to $f_{mn}$, i.e.,
 $\bm{H}^{f}_{m}$ $=\left[\frac{\partial^{2}\Phi_{m}(\bm{f}_{m},\bm{B}_{m})}{\partial f_{mn}\partial f_{mn^{\prime}}}\right]_{n,n^{\prime}\in\mathcal{N}}$ (16)
 $=-diag\left(h^{f}_{m1},h^{f}_{m2}\ldots,h^{f}_{mn}\right)<\bm{0},$
-where $h^{f}_{mn}=p_{mn}I_{mn}\frac{2x_{m}\tau_{n}}{f_{mn}^{3}}\log(1/\theta_{m})$.
-Similarly, the $\bm{H}^{B}_{m}$ can be calculated as
+where $h^{f}_{mn}=p_{mn}I_{mn}\frac{2x_{m}\tau_{n}}{f_{mn}^{3}}\log(1/\theta_{m})$. Similarly, the block matrix $\bm{H}^{B}_{m}$ can be calculated by
 $\bm{H}^{B}_{m}$ $=\left[\frac{\partial^{2}\Phi_{m}(\bm{f}_{m},\bm{B}_{m})}{\partial B_{mn}\partial B_{mn^{\prime}}}\right]_{n,n^{\prime}\in\mathcal{N}}$ (17)
@@
 where $h^{B}_{mn}=\frac{2b_{mn}p_{mn}I_{mn}}{B_{mn}^{3}\log_{2}(1+\varsigma_{mn})}$.
-We can easily show that $\bm{H}^{f}_{m}$ and $\bm{H}^{B}_{m}$ are negative semi-definite, and thus $\bm{H}_{m}$ is negative semi-definite and $\Phi_{m}(\bm{f}_{m},\bm{B}_{m})$ is concave.
-Moreover, according to the convexity condition of its constraints and Definition 3, it can be proved that Problem 1 has a unique optimal solution $(\bm{f}_{m}^{*},\bm{B}_{m}^{*})$.
+We find that $\bm{H}^{f}_{m}$ and $\bm{H}^{B}_{m}$ are both negative semi-definite.
+Hence, $\bm{H}_{m}$ is negative semi-definite, implying $\Phi_{m}(\bm{f}_{m},\bm{B}_{m})$ is strictly concave.
+Since the constraints of Problem 1 are affine and the strategy space is closed and bounded, Definition 3 ensures the existence of a unique global optimum $(\bm{f}_{m}^{*},\bm{B}_{m}^{*})$.
 ∎
-Next, we can obtain the best response function of MU $m$ by solving Problem 1. Based on the first-order condition and setting the derivative to zero, we can derive the following proposition.
-Proposition 1. Given the digital currency price offered by the MSPs and the Lagrange multipliers induced by the constraints, the best response function $f_{mn}$ of MU $m$ satisfies:
+Consequently, MU $m$ has a unique best-response resource allocation for each MSP by solving Problem 1. From the Karush–Kuhn–Tucker (KKT) conditions, we establish the following proposition.
+Proposition 1.
+Given the digital currency price offered by the MSPs and the Lagrange multipliers induced by the constraints, the optimal computational resource $f_{mn}$ allocated for MSP $n$ by MU $m$ satisfies:
 $f_{mn}^{*}=\begin{cases}\sqrt{\frac{p_{mn}I_{mn}x_{m}\tau_{n}}{c_{m}^{f}}},&\text{ $p_{mn}\geq\frac{F^{2}}{I_{mn}}$},\\ F\sqrt{\frac{x_{m}\tau_{n}}{c_{m}^{f}}},&\text{\,\,$0<p_{mn}<\frac{F^{2}}{I_{mn}}$},\\ 0,&\,\,\text{otherwise},\end{cases}$ (18)
-and the best response function $B_{mn}$ of MU $m$ satisfies:
-$B_{mn}^{*}=\begin{cases}\sqrt{\frac{p_{mn}I_{mn}b_{mn}}{c_{m}^{B}log_{2}(1+\varsigma_{mn})}},&\text{$p_{mn}\geq\frac{F^{2}}{I_{mn}}$},\\ F\sqrt{\frac{b_{mn}}{c_{m}^{B}log_{2}(1+\varsigma_{mn})}},&\text{\,\,$0<p_{mn}<\frac{F^{2}}{I_{mn}}$},\\ 0,&\text{otherwise},\end{cases}$ (19)
-where $F=log(1/\theta_{m})\sqrt{\frac{x_{m}c_{m}^{f}}{\tau_{n}}}+\frac{\sqrt{b_{mn}c_{m}^{B}}}{\tau_{n}\sqrt{log_{2}(1+\varsigma_{mn})}}$.
-Proof.
-First, we transform Problem 1 into an unconstrained optimization problem using the Lagrangian dual method. The Lagrangian function associated with MU $m$ is formulated as
-$L_{m}\left(\bm{f}_{m},\bm{B}_{m},\bm{p}_{m},\lambda_{m},\beta_{m},\delta_{mn},\gamma_{m}\right)=\sum_{n}p_{mn}V_{mn}$ (20)
-$-\sum_{n}\left[c_{m}^{f}f_{mn}log(1/\theta_{m})+c_{m}^{B}B_{mn}\right]$
-$-\lambda_{m}(\sum_{n}f_{mn}-f_{m}^{max})-\beta_{m}(\sum_{n}B_{mn}-B_{m}^{max})$
-$-\sum_{n}\delta_{mn}[log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}log_{2}(1+\varsigma_{mn})}-\tau_{n}]$
-$-\gamma_{m}[S_{m}-T^{req}(f_{m}^{max}-\sum_{n}f_{mn})],$
-where $\lambda_{m}\geq 0,\beta_{m}\geq 0$, $\delta_{mn}\geq 0$ and $\gamma_{m}\geq 0$ are the Lagrangian dual variables corresponding to constraints C1, C2 and C3, respectively.
-Then, the first-order derivative of $L_{m}$ with respect to $f_{mn}$ and $B_{mn}$ can be derived as
-$\frac{\partial L_{m}}{\partial f_{mn}}$ $=p_{mn}I_{mn}\frac{x_{m}\tau_{n}log(1/\theta_{m})}{f_{mn}^{2}}-c_{m}^{f}log(1/\theta_{m})$ (21)
-$-\lambda_{m}+\delta_{mn}log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}^{2}}-\gamma_{m}T^{req},$
-and
-$\frac{\partial L_{m}}{\partial B_{mn}}$ $=p_{mn}I_{mn}\frac{b_{mn}}{B_{mn}^{2}log_{2}(1+\varsigma_{mn})}-c_{m}^{B}-\beta_{m}$ (22)
-$+\delta_{mn}\frac{b_{mn}}{B_{mn}^{2}log_{2}(1+\varsigma_{mn})}.$
-From (21) and the constraints in (11), the KKT conditions are given by
-$\frac{\partial L_{m}}{\partial f_{mn}}=0,\,\,\lambda_{m}(\sum_{n}f_{mn}-f_{m}^{max})=0,$ (23)
-$\delta_{mn}[log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}log_{2}(1+\varsigma_{mn})}-\tau_{n}]=0,$
-$\gamma_{m}[S_{m}-T^{req}(f_{m}^{max}-\sum_{n}f_{mn})]=0,$
-$\lambda_{m},\delta_{mn},\gamma_{m}\geq 0,C1,C2,C3.$
-Since $S_{m}<T^{req}\left(f_{m}^{max}-\sum_{n}f_{mn}\right)$ and $\sum_{n}f_{mn}<f_{m}^{max}$, we can obtain $\lambda_{m}=0$ and $\gamma_{m}=0$.
-Furthermore, $\frac{\partial L_{m}}{\partial f_{mn}}=0$ can be converted to
-$(p_{mn}I_{mn}+\delta_{mn})\frac{x_{m}\tau_{n}}{f_{mn}^{2}}-c_{m}^{f}=0.$ (24)
-Similarly, the KKT conditions for $B_{mn}$ are obtained as
-$\frac{\partial L_{m}}{\partial B_{mn}}=0,\,\,\beta_{m}(\sum_{n}B_{mn}-B_{m}^{max})=0,$ (25)
-$\delta_{mn}[log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}log_{2}(1+\varsigma_{mn})}-\tau_{n}]=0,$
-$\beta_{m},\delta_{mn}\geq 0,C1,C2.$
-Since $\sum_{n}B_{mn}<B_{m}^{max}$, we can obtain $\beta_{m}=0$. Furthermore, $\frac{\partial L_{m}}{\partial B_{mn}}=0$ can be converted to
-$(p_{mn}I_{mn}+\delta_{mn})\frac{b_{mn}}{B_{mn}^{2}log_{2}(1+\varsigma_{mn})}-c_{m}^{B}=0.$ (26)
-Then, based on the value of $\delta_{mn}$, the optimal allocation of computation and communication resources exists in the following two cases:
-  • ($\delta_{mn}=0$): According to (24) and (26), we have
-    $f_{mn}=\sqrt{\frac{p_{mn}I_{mn}x_{m}\tau_{n}}{c_{m}^{f}}},B_{mn}=\sqrt{\frac{p_{mn}I_{mn}b_{mn}}{c_{m}^{B}log_{2}(1+\varsigma_{mn})}}.$ (27)
-  • ($\delta_{mn}>0$): Substituting $f_{mn}$ and $B_{mn}$ into $\delta_{mn}\left[log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}log_{2}(1+\varsigma_{mn})}-\tau_{n}\right]=0$, we can get
-    $\delta_{mn}=F^{2}-p_{mn}I_{mn},$ (28)
-    $F=log(1/\theta_{m})\sqrt{\frac{x_{m}c_{m}^{f}}{\tau_{n}}}+\frac{\sqrt{b_{mn}c_{m}^{B}}}{\tau_{n}\sqrt{log_{2}(1+\varsigma_{mn})}}.$
-    Next, bring $\delta_{mn}$ into $f_{mn}$ and $B_{mn}$, we have
-    $f_{mn}=F\sqrt{\frac{x_{m}\tau_{n}}{c_{m}^{f}}},B_{mn}=F\sqrt{\frac{b_{mn}}{c_{m}^{B}log_{2}(1+\varsigma_{mn})}}.$ (29)
+and the optimal communication bandwidth $B_{mn}$ is:
+$B_{mn}^{*}=\begin{cases}\sqrt{\frac{p_{mn}I_{mn}b_{mn}}{c_{m}^{B}\log_{2}(1+\varsigma_{mn})}},&\text{$p_{mn}\geq\frac{F^{2}}{I_{mn}}$},\\ F\sqrt{\frac{b_{mn}}{c_{m}^{B}\log_{2}(1+\varsigma_{mn})}},&\text{\,\,$0<p_{mn}<\frac{F^{2}}{I_{mn}}$},\\ 0,&\text{otherwise},\end{cases}$ (19)
+where $F=\log(1/\theta_{m})\sqrt{\frac{x_{m}c_{m}^{f}}{\tau_{n}}}+\frac{\sqrt{b_{mn}c_{m}^{B}}}{\tau_{n}\sqrt{\log_{2}(1+\varsigma_{mn})}}$.
+For the detailed proof of Proposition 1, please refer to Appendix A.
 According to Proposition 1, the optimal strategy for each MU $m$ is influenced by five main factors: $p_{mn}$, $I_{mn}$, $c_{m}^{f}$, $c_{m}^{B}$ and $\tau_{n}$.
-Specifically, MU $m$ focuses on the rewards paid by MSPs, and the MU tends to put more computation and communication resources into the contribution when $p$ increases.
+Specifically, MU $m$ focuses on the rewards paid by MSPs, and the MU tends to put more computational and communication resources into the contribution when $p$ increases.
 Moreover, MUs are more willing to invest more resources when the potential contribution prediction $I_{mn}$ is higher, resulting in greater benefits.
 In addition, the virtual deadline $\tau_{n}$ set by the MSP $n$ determines the minimum criteria for resource allocation.
-V-B Upper Level: Optimal Reward Decision Equilibrium Analysis
-Given the responses $(\bm{f},\bm{B})$ from all MUs, and other MSPs’ decisions $\bm{p}_{\textrm{-}n}$, the optimal reward decision $\bm{p}_{n}$ of MSP $n$ can be obtained by solving Problem 2, defined as follows.
-Proposition 2. Given $\bm{p}_{\textrm{-}n}$, MSP $n$’s best response strategy is
+V-B Upper Level: Optimal Reward Equilibrium among MSPs
+In the upper level of the EPEC, each MSP $n$ competes with other MSPs and determines its reward vector $\bm{p}_{n}$.
+Specifically, given the responses $(\bm{f},\bm{B})$ from all MUs, and other MSPs’ decisions $\bm{p}_{\textrm{-}n}$, the optimal reward decision $\bm{p}_{n}$ of MSP $n$ can be obtained by solving Problem 2, defined as follows.
+Proposition 2. Given other MSPs’ reward vectors $\bm{p}_{\textrm{-}n}$, the optimal strategy of MSP $n$ is
 $\bm{p}_{n}^{*}=arg\max_{p_{mn}>0}\Psi_{n}\left(\bm{f}_{n}^{*}(\bm{p}_{n},\bm{p}_{-n}),\bm{B}_{n}^{*}(\bm{p}_{n},\bm{p}_{-n}),\bm{p}_{-n}\right).$ (20)
@@
 Theorem 2.
-There exists a unique Nash equilibrium (NE) [32] in the multi-MSP rewarding game $\Omega$.
+There exists a unique Nash equilibrium in the multi-MSP rewarding game $\Omega$, ensuring a unique optimal set of reward decisions $\{\bm{p}_{n}^{*}\}_{n\in\mathcal{N}}$.
 Proof.
-We define the Hessian matrix of $\Psi_{n}$ with respect to $\bm{p}_{n}$ as $(\bm{\Lambda}_{n}+\bm{H}_{n})$.
+We define the Hessian matrix of $\Psi_{n}$ with respect to its reward vector $\bm{p}_{n}$ as $(\bm{\Lambda}_{n}+\bm{H}_{n})$.
 The matrix $\bm{\Lambda}_{n}=diag\left(\frac{\partial^{2}\Psi_{n}}{\partial p_{1n}^{2}},\ldots,\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}\right)$ and the second-order partial derivative matrix $\bm{H}_{n}$ is expressed by
@@
 $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}$ $=\mu_{n}\frac{V_{mn}^{\prime\prime}(1+\sum_{m=1}^{M}V_{mn})-(V_{mn}^{\prime})^{2}}{(1+\sum_{m=1}^{M}V_{mn})^{2}}$ (22)
-$-2V_{mn}^{\prime}-p_{mn}V_{mn}^{\prime\prime},\forall m\in\mathcal{M},$
-$\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}\partial p_{m^{\prime}n}}=$ $-\mu_{n}\frac{V_{mn}^{\prime}V_{m^{\prime}n}^{\prime}}{(1+\sum_{m=1}^{M}V_{mn})^{2}},$ (33)
+$-2V_{mn}^{\prime}-p_{mn}V_{mn}^{\prime\prime}<0,\forall m\in\mathcal{M},$
+$\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}\partial p_{m^{\prime}n}}=$ $-\mu_{n}\frac{V_{mn}^{\prime}V_{m^{\prime}n}^{\prime}}{(1+\sum_{m=1}^{M}V_{mn})^{2}}<0,$ (23)
 $\forall m\in\mathcal{M},m\neq m^{\prime}.$
-To obtain the values of $V_{mn}^{\prime}$ and $V_{mn}^{\prime\prime}$, we need to obtain $\frac{\partial f_{mn}}{\partial p_{mn}}$, $\frac{\partial^{2}f_{mn}}{\partial p_{mn}^{2}}$, $\frac{\partial B_{mn}}{\partial p_{mn}}$ and $\frac{\partial^{2}B_{mn}}{\partial p_{mn}^{2}}$.
-From (24), we can obtain $\frac{\partial f_{mn}}{\partial p_{mn}}$ and $\frac{\partial^{2}f_{mn}}{\partial p_{mn}^{2}}$, the steps of which are shown as follows.
-The first-order derivative of $\frac{\partial L_{m}}{\partial f_{mn}}=0$ with respect to $p_{mn}$ is expressed as $I_{mn}\frac{x_{m}\tau_{n}}{f_{mn}^{2}}-2(p_{mn}I_{mn}+\delta_{mn})\frac{x_{m}\tau_{n}}{f_{mn}^{3}}\cdot\frac{\partial f_{mn}}{\partial p_{mn}}=0$.
-Consequently, we can obtain
-$\frac{\partial f_{mn}}{\partial p_{mn}}=\frac{I_{mn}f_{mn}}{2(p_{mn}I_{mn}+\delta_{mn})}>0.$ (34)
-The second-order derivative of $\frac{\partial L_{m}}{\partial f_{mn}}=0$ with respect to $p_{mn}$ is expressed as
-$\frac{\partial^{2}f_{mn}}{\partial p_{mn}^{2}}=\frac{I_{mn}\frac{\partial f_{mn}}{\partial p_{mn}}(p_{mn}I_{mn}+\delta_{mn})-I_{mn}^{2}f_{mn}}{2(p_{mn}I_{mn}+\delta_{mn})^{2}},$ (35)
-by substituting (34) into (35), we have
-$\frac{\partial^{2}f_{mn}}{\partial p_{mn}^{2}}=\frac{-I_{mn}^{2}f_{mn}}{4(p_{mn}I_{mn}+\delta_{mn})^{2}}<0.$ (36)
-Likewise, we derive $\frac{\partial B_{mn}}{\partial p_{mn}}$ and $\frac{\partial^{2}B_{mn}}{\partial p_{mn}^{2}}$ with the following steps based on (26).
-The first-order derivative of $\frac{\partial L_{m}}{\partial B_{mn}}=0$ with respect to $p_{mn}$ is expressed as $I_{mn}\frac{b_{mn}}{B_{mn}^{2}log_{2}(1+\varsigma_{mn})}-2(p_{mn}I_{mn}+\delta_{mn})\frac{b_{mn}}{B_{mn}^{3}log_{2}(1+\varsigma_{mn})}\cdot\frac{\partial B_{mn}}{\partial p_{mn}}=0$.
-Accordingly, we have
-$\frac{\partial B_{mn}}{\partial p_{mn}}=\frac{I_{mn}B_{mn}}{2(p_{mn}I_{mn}+\delta_{mn})}>0.$ (37)
-Similarly, we obtain the second-order derivative of $\frac{\partial L_{m}}{\partial B_{mn}}=0$ with respect to $p_{mn}$ as
-$\frac{\partial^{2}B_{mn}}{\partial p_{mn}^{2}}=\frac{I_{mn}\frac{\partial B_{mn}}{\partial p_{mn}}(p_{mn}I_{mn}+\delta_{mn})-I_{mn}^{2}B_{mn}}{2(p_{mn}I_{mn}+\delta_{mn})^{2}},$ (38)
-by substituting (37) into (38), we have
-$\frac{\partial^{2}B_{mn}}{\partial p_{mn}^{2}}=\frac{-I_{mn}^{2}B_{mn}}{4(p_{mn}I_{mn}+\delta_{mn})^{2}}<0.$ (39)
-Based on (34), (36), (37), and (39), we can easily infer that $V_{mn}^{\prime}>0$ and $V_{mn}^{\prime\prime}<0$, as shown in (40).
-Then, we can obtain $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}\partial p_{m^{\prime}n}}<0$.
-For $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}$, We substitute the specific expressions for $V_{mn}^{\prime}$ and $V_{mn}^{\prime\prime}$ into (32) and obtain the following result as shown in (41).
-Next, we randomly choose a vector $\bm{h}\in\mathbb{R}^{M\times 1}$ with elements not all 00.
+The proof of $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}<0$ and $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}\partial p_{m^{\prime}n}}<0$ can be found in Appendix B.
+We randomly choose a vector $\bm{h}\in\mathbb{R}^{M\times 1}$ with elements not all 00.
 For $\bm{\Lambda}_{n}$, it is easy to see that $\bm{h}^{T}\bm{\Lambda}_{n}\bm{h}=\sum_{m}\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}(h^{m})^{2}<0$.
@@
 Thus, we have $\bm{h}^{T}(\bm{\Lambda}_{n}+\bm{H_{n}})\bm{h}=\bm{h}^{T}\bm{\Lambda}_{n}\bm{h}+\bm{h}^{T}\bm{H}_{n}\bm{h}<0$, indicating that the utility function $\Psi_{n}$ is strictly concave.
-So far, the uniqueness of the Nash equilibrium solution to the multi-MSP rewarding game $\Omega$ has been proved.
-$V_{mn}^{\prime}=\frac{\partial V_{mn}}{\partial p_{mn}}=I_{mn}\left(\frac{x_{m}\tau_{n}log(1/\theta_{m})}{f_{mn}^{2}}\cdot\frac{\partial f_{mn}}{\partial p_{mn}}+\frac{b_{mn}}{B_{mn}^{2}log_{2}(1+\varsigma_{mn})}\cdot\frac{\partial B_{mn}}{\partial p_{mn}}\right)>0,$ (40)
-$V_{mn}^{\prime\prime}=\frac{\partial^{2}V_{mn}}{\partial p_{mn}^{2}}=I_{mn}\left[\frac{-2x_{m}\tau_{n}log(1/\theta_{m})}{f_{mn}^{3}}\cdot(\frac{\partial f_{mn}}{\partial p_{mn}})^{2}+\frac{x_{m}\tau_{n}log(1/\theta_{m})}{f_{mn}^{2}}\cdot\frac{\partial^{2}f_{mn}}{\partial p_{mn}^{2}}\right.$
-$\left.-\frac{2b_{mn}}{B_{mn}^{3}log_{2}(1+\varsigma_{mn})}\cdot(\frac{\partial B_{mn}}{\partial p_{mn}})^{2}+\frac{b_{mn}}{B_{mn}^{2}log_{2}(1+\varsigma_{mn})}\cdot\frac{\partial^{2}B_{mn}}{\partial p_{mn}^{2}}\right]<0.$
-$\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}=-\left[\frac{4a_{1}I_{mn}^{2}\delta_{mn}+a_{1}I_{mn}^{3}p_{mn}}{4f_{mn}(\delta_{mn}+I_{mn}p_{mn})^{2}}+\frac{4b_{mn}I_{mn}^{2}\delta_{mn}+b_{mn}I_{mn}^{3}p_{mn}}{4a_{2}B_{mn}(\delta_{mn}+I_{mn}p_{mn})^{2}}+I_{mn}^{2}\frac{\left(\frac{a_{1}a_{2}I_{mn}B_{mn}+b_{mn}I_{mn}f_{mn}}{2a_{2}f_{mn}B_{mn}(\delta_{mn}+I_{mn}p_{mn})}\right)^{2}}{(1+\sum_{m}V_{mn})^{2}}\right.$ (41)
-$\left.+\frac{3I_{mn}^{3}(a_{1}a_{2}B_{mn}+b_{mn}f_{mn})}{4a_{2}f_{mn}B_{mn}(\delta_{mn}+I_{mn}p_{mn})^{2}(1+\sum_{m}V_{mn})}\right]<0,\text{where}\,a_{1}=x_{m}\tau_{n}log(1/\theta_{m}),a_{2}=log_{2}(1+\varsigma_{mn}).$
-1
-2Input: Action threshold $[p_{min},p_{max}]$, learning rate;
-3 Output: The reward decisions $\bm{p}$ of MSPs, resource allocation $[\bm{f},\bm{B}]$ of MUs;
-4 Initialization: local observation $s_{n}$, actor network $\bm{\alpha}_{n}$, critic network $\bm{\beta}_{n}$ and episode buffer $\mathcal{D}_{n}$ for each MSP agent $n$ ;
+So far, the uniqueness of the Nash equilibrium [37] solution to the multi-MSP rewarding game $\Omega$ has been proved.
+∎
+1 Input: Action threshold $[p_{min},p_{max}]$ and number of episodes $\mathcal{T}$;
+2 Initialization: local observation $s_{n}$, actor network $\bm{\alpha}_{n}$, critic network $\bm{\beta}_{n}$, and episode buffer $\mathcal{D}_{n}$ for each MSP agent $n\in\mathcal{N}$;
 3 # Trading process using MDDR:
-6 for  Episode = $1,2,\ldots$ do
-7      for Epoch $k=1,\ldots,|\mathcal{D}|$ do
-8            foreach MSP agent $n\in\mathcal{N}$ do
+4 for episode = $1,2,\ldots,\mathcal{T}$ do
+5      Concurrently for each MSP agent $n\in\mathcal{N}$:
+6       for epoch $k=1,\ldots,|\mathcal{D}_{n}|$ do
 7            Observe space $s_{n}(k)$;
-10                   Choose price $\bm{p}_{n}(k)$ within $[p_{min},p_{max}]$ by sampling from $\pi_{\alpha_{n}}(s_{n}(k))$ ;
-11             end foreach
-12            All MSP agents do:
-13             Perform a joint reward $\bm{p}(k)=\left\{\bm{p}_{1}(k),\ldots,\bm{p}_{n}(k)\right\}\in\mathcal{A}$;
-14             Broadcast reward decisions;
+8             Choose price $\bm{p}_{n}(k)\in[p_{min},p_{max}]$ by sampling from its current policy $\pi_{\alpha_{n}}(s_{n}(k))$;
+9             Broadcast reward decisions to MUs;
 10             Interact with the environment and receive responses from the MUs;
 11             Calculate utility according to (12) ;
-17             Store $s_{n}(k)=[V_{n}(k),\bm{p}_{n}(k),\mathcal{U}_{n}(k),V_{n}(k+1)]$ ;
+12             Store transition $e_{n}(k)=[s_{n}(k),\bm{p}_{n}(k),\mathcal{U}_{n}(k),s_{n}(k+1)]$;
 13       end for
-19      Update actor network $\bm{\alpha}$ and critic network $\bm{\beta}$ for each MSP agent as [33] ;
-20       Clear up relay buffer $\mathcal{D}$ ;
+14      Update actor network $\bm{\alpha}_{n}$ and critic network $\bm{\beta}_{n}$ for each MSP agent $n$;
+15       Clear the episode buffer $\mathcal{D}_{n}$ for each MSP agent $n$;
 16 end for
-22# FL training guided by trading results within $T$:
+17Return: Reward decisions $\bm{p}$ of MSPs and resource allocation $[\bm{f},\bm{B}]$ of MUs;
+18 # FL training guided by trading results within time $T$:
 19 foreach MSP agent $n\in\mathcal{N}$ do
 20      foreach MU $m\in\mathcal{M}$ do
-25             Based on the trading results, utilize $f_{mn}$ and $B_{mn}$ to locally train the task of MSP $n$, and subsequently upload the model to the MSP $n$.
+21             Based on the trading results, utilize $f_{mn}$ and $B_{mn}$ to perform the local FL task for MSP $n$, and then upload the model to MSP $n$.
 22       end foreach
-27      Aggregate and issue an updated global model;
+23      Aggregate received local models and issue an updated global model;
 24       Supply the updated model to the AR engine;
@@
 Algorithm 1 Entire Process of the Framework.
-We verify the existence and uniqueness of the equilibrium solution for EPEC through Theorem 1 and Theorem 2.
-However, balancing multiple and conflicting interests, i.e., the utilities of MSPs and MUs, inevitably complicates the optimization problem.
-To find a solution that optimizes the profits of MSPs while taking into account the utilities of MUs, we develop a reinforcement learning-based strategy in the following section.
+Utilizing Theorem 1 and Theorem 2, we prove the existence and uniqueness of the hierarchical equilibrium in the proposed EPEC model.
+Concretely, each MU has a unique best-response resource allocation strategy in reaction to any given MSP rewards, while each MSP determines a unique optimal pricing to maximize its own utility given the rewards of other MSPs.
+Although the backward induction approach theoretically ensures an equilibrium solution to (14), solving it in real-world vehicular metaverse environments can be challenging due to dynamic network conditions and limited system knowledge (e.g., private cost parameters).
+In the following section, we address these practical issues by introducing a fully distributed reinforcement learning approach, enabling MSPs to adaptively and privately optimize their reward decisions in real time.
 VI Dynamic Reward Strategy for MSPs
-According to the EPEC proved above, it is possible to obtain an optimal solution.
-However, the optimal approach faces the following practical difficulties: (i) under the time-varying network, the optimal approach is time-consuming due to the complexity of (12), which is non-linear with complicated form and the reward decisions of MSPs are tightly coupled.
-(ii) MSPs must know private information about MUs (e.g., cost factors) as a prior condition for determining their rewards, which violates the privacy of MUs.
-To overcome the above difficulties, we further formulate the multi-MSP rewarding game as a multi-agent Markov decision process (MAMDP) [34] to adapt to dynamic channel conditions, where each MSP acts as an individual agent to make intelligent reward decisions.
+According to the EPEC proven above, it is possible to obtain an optimal solution.
+However, the optimal approach faces the following practical challenges: (i) In a time-varying network, the optimal approach is time-consuming due to the complexity of (12), which is a non-linear problem with a complicated structure.
+Additionally, the reward decisions of MSPs are tightly coupled.
+(ii) MSPs must have prior knowledge of private information about MUs (e.g., cost factors) to determine their rewards, which raises privacy concerns for MUs.
+To address these challenges, we further formulate the multi-MSP rewarding game as a multi-agent Markov decision process (MAMDP) [35], which adapts to dynamic channel conditions.
+In this framework, each MSP is modeled as an individual agent that makes intelligent reward decisions.
 VI-A MAMDP for Multi-MSP Rewarding Game $\Omega$
@@
 In our work, each MSP agent operates the DRL-based dynamic reward algorithm in a fully distributed manner.
-Algorithm 1 shows the pseudocode for model trading using MDDR under a dynamic environment (Lines 5-19) and the guidance for the FL process (Lines 22-27).
-Trading process: When the game starts, each MSP $n$ initializes its actor $\bm{\alpha}_{n}$ and critic network $\bm{\beta}_{n}$, essential parameters and inference loss for trading (Line 3).
-During each game, each agent feeds the observed information into the policy network $\bm{\alpha}_{n}$ to derive its reward policy (Lines 8-9).
-Then, all the agents generate reward decisions and broadcast them to the MUs (Lines 12-13).
-After completing these steps, the MSP agents enter the interaction phase with the environment, receive the best feedbacks from the MUs, calculate their utilities, and store information into episode buffer $\mathcal{D}$ (Lines 14-16).
-The experience is collected through the above interactive process (Lines 8-16) until the buffer $\mathcal{D}$ is full.
-Then, the critic-network $\bm{\beta}_{n}$ evaluates the actor-network $\bm{\alpha}_{n}$ based on the experience and in turn adjusts it (Line 18).
+Algorithm 1 shows the pseudocode for model trading using MDDR under a dynamic environment (Lines 4-17) and the guidance for the FL process (Lines 19-25).
+Trading process: At the start of the game, each MSP $n$ initializes its actor $\bm{\alpha}_{n}$ and critic network $\bm{\beta}_{n}$, and episode buffer $\mathcal{D}_{n}$ for each MSP agent $n$ (Line 2).
+Each agent $n$ feeds the observed information into the policy network $\bm{\alpha}_{n}$ to derive its reward policy (Lines 7-8).
+Subsequently, each MSP agent $n$ broadcasts its reward decisions, interacts with the environment, receives feedback from the MUs, calculates its utility, and stores the experience in the episode buffer $\mathcal{D}_{n}$ (Lines 9-12).
+The experience is collected through this interactive process (Lines 7-12) until the buffer $\mathcal{D}_{n}$ is full.
+Once the buffer is full, the critic network $\bm{\beta}_{n}$ evaluates the actor network $\bm{\alpha}_{n}$ based on the collected experience and updates it accordingly (Line 14).
+Finally, the episode buffer is cleared in preparation for the next episode (Line 15).
 The game will terminate when the maximum number of training episodes is reached and then the trading results will be obtained.
 FL training guided by trading results: After obtaining the trading results (MSPs’ reward decisions $\bm{p}$ and resource allocation $(\bm{f},\bm{B})$), MUs and MSPs enter the FL phase.
-MUs follow the resource allocation scheme for local training and model uploading (Line 24).
-Subsequently, the MSP is responsible for aggregating and updating the global model (Line 26).
-Finally, the updated model is provided for use by the AR engine (Line 27).
+MUs follow the resource allocation scheme for local training and model uploading (Line 21).
+Then, the MSP is responsible for aggregating and updating the global model (Line 23).
+Finally, the updated model is provided for use by the AR engine (Line 24).
 This process iterates multiple times until the total FL time $T$ is reached.
-During real-time interactions, the AR service of the MSP can be enhanced in terms of immersion and service quality by the updated model.
+During real-time interactions, the MSP’s AR service immersion and quality can be enhanced by utilizing the updated model.
 VI-C The Complexity and Convergence of MDDR
@@
 Therefore, under the uniform rewards of MSPs, we develop and extend four benchmark schemes for multi-MSP and multi-MU scenarios while meeting FL deadlines and resource constraints of devices.
-  • x_based: Inspired by the incentive mechanism of the Stackelberg game for federated learning [25], the MU allocates computation and communication resources based on the amount of data involved in local training.
+  • x_based: Inspired by the incentive mechanism of the Stackelberg game for federated learning [31], the MU allocates computational and communication resources based on the amount of data involved in local training.
   • w_based: The MU adjusts the resource allocation according to the inference loss [43], i.e., the potential value of local data in our work.
   • w_x_based: Combining the x_based and w_based designs, the MU allocates resources considering both the data size and the potential value of data.
-  • fixed: Taking into account the differences in performance of the various devices that exist in reality, we randomly allocate a fixed set of computation and communication resources to reflect the heterogeneity of devices.
+  • fixed: Taking into account the differences in performance of the various devices that exist in reality, we randomly allocate a fixed set of computational and communication resources to reflect the heterogeneity of devices.
 TABLE III: Time comparison of five schemes for achieving the same performance.
 Schemes MNIST GTSRB
-90% Accuracy 50% mAP 70% mAP50
+Training Time to reach 90% Accuracy (s) Training Time to reach 50% mAP (s) Training Time to reach 70% mAP@5050{50}50 (s)
 MSP 1 MSP 2 MSP 3 MSP 1 MSP 2 MSP 3 MSP 1 MSP 2 MSP 3
-immersion-aware 15.23s 13.09s 7.59s 513.16s 581.09s 623.89s 418.28s 447.69s 451.27s
-x_based 17.79s 18.03s 19.80s 812.53s 898.59s 1061.50s 596.57s 679.91s 890.47s
-w_based 30s+ 24.04s 10.28s 1064.83s 1175.58s 1200s+ 774.09s 953.50s 1013.89s
-w_x_based 23.20s 19.08s 12.34s 901.61s 1190s 1194s 596.35s 848.68s 932.73s
-fixed 30s+ 30s+ 30s+ 1200s+ 1200s+ 1200s+ 1200s+ 1200s+ 1200s+
-  • The $1200$s+ and $30$s+ indicate that the guidance time for the current trading is exceeded.
+immersion-aware 15.23 13.09 7.59 513.16 581.09 623.89 418.28 447.69 451.27
+x_based 17.79 18.03 19.80 812.53 898.59 1061.50 596.57 679.91 890.47
+w_based 30+ 24.04 10.28 1064.83 1175.58 1200+ 774.09 953.50 1013.89
+w_x_based 23.20 19.08 12.34 901.61 1190 1194 596.35 848.68 932.73
+fixed 30+ 30+ 30+ 1200+ 1200+ 1200+ 1200+ 1200+ 1200+
+  • Note: $30$+ or $1200$+ indicates that the scheme required more than $30$ seconds or $1200$ seconds, respectively, to achieve the target accuracy, exceeding the guideline time for the trading results.
 VII-C Performance Evaluation of MDDR
@@
 9(a), 9(b), and 9(c) show the performance improvement in classification accuracy over time, where the immersion-aware outperforms the other schemes.
-This is because immersion-aware utilizes IoM to incentivize MUs to allocate computation and communication resources for providing learning models to MSPs.
+This is because immersion-aware utilizes IoM to incentivize MUs to allocate computational and communication resources for providing learning models to MSPs.
 The IoM metric not only evaluates the freshness and accuracy of the model but also considers the amount and potential value of raw data used for training.
@@
 For example, at the $10$th second, the accuracy of MSP 3 is about $88\%$, which is $14\%$ and $3\%$ higher than those of MSP 1 and MSP 2, respectively.
-This is because the potential value $w$ is higher for MSP 3, leading to more rewards being given by MSP 3.
-As a result, the MUs will allocate more computation and communication resources to provide learning models for MSP 3 than for MSP 1 and MSP 2.
+This is because the potential value $\omega$ is higher for MSP 3, leading to more rewards being given by MSP 3.
+As a result, the MUs will allocate more computational and communication resources to provide learning models for MSP 3 than for MSP 1 and MSP 2.
 For the three different MSPs, some fluctuations are normal for model training (e.g., at time $10$s).
@@
 Meanwhile, it is clear from Figs.
-9(d)9(e) and 9(f) that the nature of the effectiveness of the five schemes is influenced by IoM.
+9(d)9(e) and 9(f) that the effectiveness of the five schemes is influenced by IoM.
 With the higher IoM, the performance is improved faster, resulting in a better immersive experience.
@@
 9(a).
-The main reason is the higher potential value $w$ associated with MSP 2 and MSP 3.
+The main reason is the higher potential value $\omega$ associated with MSP 2 and MSP 3.
 For object detection on GTSRB, Figs.
-10(a), 10(b) and 10(c) show the mean average precision (mAP) and Table III gives the mAP50 (The value of mAP when the intersection over union (IOU) threshold is larger than $0.5$) results.
+10(a), 10(b) and 10(c) show the mean average precision (mAP) and Table III gives the mAP@5050{50}50 (The value of mAP when the intersection over union (IOU) threshold is larger than $0.5$) results.
 We observe that compared with MNIST, the performance enhancement for object detection is slower due to the complexity of the detection model.
@@
 For instance, MSP 1 has the highest IoM, and the time that it takes to reach $50\%$ mAP is $513.16$s, which is $67.93$s and $110.73$s less than those of MSP 2 and MSP 3, respectively.
-The above comparison validates the feasibility of the proposed trading framework and the advantages of IoM-based metric incentives.
+The comparison results validate the advantages of the proposed trading framework, which outperforms other schemes by improving IoM by $38.3\%$ and $37.2\%$ and reducing the training time to reach the target accuracy by $43.5\%$ and $49.8\%$, on average, for the MNIST and GTSRB datasets, respectively.
 VIII Conclusion
-To facilitate the development of vehicular metaverse services, we propose an immersion-aware model trading framework that incorporates FL to incentivize MUs to become contributors and provide locally trained data for AR services, i.e., object detection and classification.
+To facilitate the development of vehicular metaverse services, we propose an immersion-aware model trading framework that incorporates FL to incentivize MUs to contribute locally trained data for AR services such as object detection and classification.
 Specifically, we construct an EPEC problem with MSPs as leaders and MUs as followers to achieve an equilibrium among their interests.
@@
 In addition, considering the competitive relationship among MSPs and the dynamic network environment, we develop a fully distributed MDDR approach to obtain the reward decisions of MSPs.
-Through extensive simulations using AR-related vehicle and MNIST datasets, our results demonstrate that the proposed framework can efficiently provide MUs with better immersive experiences for AR services in the vehicular metaverse.
+Extensive simulations on AR-related vehicle and MNIST datasets demonstrate that the proposed framework enables more efficient and immersive AR services in the vehicular metaverse.
+Appendix A Proof of Proposition 1
+Proof.
+First, we transform Problem 1 into an unconstrained optimization problem using the Lagrangian dual method [45, 46]. The Lagrangian function associated with MU $m$ is formulated as
+$L_{m}\left(\bm{f}_{m},\bm{B}_{m},\bm{p}_{m},\lambda_{m},\beta_{m},\delta_{mn},\gamma_{m}\right)=\sum_{n}p_{mn}V_{mn}$ (24)
+$-\sum_{n}\left[c_{m}^{f}f_{mn}\log(1/\theta_{m})+c_{m}^{B}B_{mn}\right]$
+$-\lambda_{m}(\sum_{n}f_{mn}-f_{m}^{max})-\beta_{m}(\sum_{n}B_{mn}-B_{m}^{max})$
+$-\sum_{n}\delta_{mn}[\log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}\log_{2}(1+\varsigma_{mn})}-\tau_{n}]$
+$-\gamma_{m}[S_{m}-T^{req}(f_{m}^{max}-\sum_{n}f_{mn})],$
+where $\lambda_{m}\geq 0,\beta_{m}\geq 0$, $\delta_{mn}\geq 0$ and $\gamma_{m}\geq 0$ are the Lagrangian dual variables corresponding to constraints C1, C2 and C3, respectively.
+Then, the first-order derivative of $L_{m}$ with respect to $f_{mn}$ and $B_{mn}$ can be derived as
+$\frac{\partial L_{m}}{\partial f_{mn}}$ $=p_{mn}I_{mn}\frac{x_{m}\tau_{n}\log(1/\theta_{m})}{f_{mn}^{2}}-c_{m}^{f}\log(1/\theta_{m})$ (25)
+$-\lambda_{m}+\delta_{mn}\log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}^{2}}-\gamma_{m}T^{req},$
+and
+$\frac{\partial L_{m}}{\partial B_{mn}}$ $=p_{mn}I_{mn}\frac{b_{mn}}{B_{mn}^{2}\log_{2}(1+\varsigma_{mn})}-c_{m}^{B}-\beta_{m}$ (26)
+$+\delta_{mn}\frac{b_{mn}}{B_{mn}^{2}\log_{2}(1+\varsigma_{mn})}.$
+From (25) and the constraints in (11), the Karush-Kuhn-Tucker (KKT) [47] conditions are given by
+$\frac{\partial L_{m}}{\partial f_{mn}}=0,\,\,\lambda_{m}(\sum_{n}f_{mn}-f_{m}^{max})=0,$ (27)
+$\delta_{mn}[\log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}\log_{2}(1+\varsigma_{mn})}-\tau_{n}]=0,$
+$\gamma_{m}[S_{m}-T^{req}(f_{m}^{max}-\sum_{n}f_{mn})]=0,$
+$\lambda_{m},\delta_{mn},\gamma_{m}\geq 0,C1,C2,C3.$
+Since $S_{m}<T^{req}\left(f_{m}^{max}-\sum_{n}f_{mn}\right)$ and $\sum_{n}f_{mn}<f_{m}^{max}$, we can obtain $\lambda_{m}=0$ and $\gamma_{m}=0$.
+Furthermore, $\frac{\partial L_{m}}{\partial f_{mn}}=0$ can be converted to
+$(p_{mn}I_{mn}+\delta_{mn})\frac{x_{m}\tau_{n}}{f_{mn}^{2}}-c_{m}^{f}=0.$ (28)
+Similarly, the KKT conditions for $B_{mn}$ are obtained as
+$\frac{\partial L_{m}}{\partial B_{mn}}=0,\,\,\beta_{m}(\sum_{n}B_{mn}-B_{m}^{max})=0,$ (29)
+$\delta_{mn}[\log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}\log_{2}(1+\varsigma_{mn})}-\tau_{n}]=0,$
+$\beta_{m},\delta_{mn}\geq 0,C1,C2.$
+Since $\sum_{n}B_{mn}<B_{m}^{max}$, we can obtain $\beta_{m}=0$. Furthermore, $\frac{\partial L_{m}}{\partial B_{mn}}=0$ can be converted to
+$(p_{mn}I_{mn}+\delta_{mn})\frac{b_{mn}}{B_{mn}^{2}\log_{2}(1+\varsigma_{mn})}-c_{m}^{B}=0.$ (30)
+Then, based on the value of $\delta_{mn}$, the optimal allocation of computation and communication resources exists in the following two cases:
+  • ($\delta_{mn}=0$): According to (28) and (30), we have
+    $f_{mn}=\sqrt{\frac{p_{mn}I_{mn}x_{m}\tau_{n}}{c_{m}^{f}}},B_{mn}=\sqrt{\frac{p_{mn}I_{mn}b_{mn}}{c_{m}^{B}\log_{2}(1+\varsigma_{mn})}}.$ (31)
+  • ($\delta_{mn}>0$): Substituting $f_{mn}$ and $B_{mn}$ into $\delta_{mn}\left[\log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}\log_{2}(1+\varsigma_{mn})}-\tau_{n}\right]=0$, we can get
+    $\delta_{mn}=F^{2}-p_{mn}I_{mn},$ (32)
+    $F=\log(1/\theta_{m})\sqrt{\frac{x_{m}c_{m}^{f}}{\tau_{n}}}+\frac{\sqrt{b_{mn}c_{m}^{B}}}{\tau_{n}\sqrt{\log_{2}(1+\varsigma_{mn})}}.$
+    Next, bring $\delta_{mn}$ into $f_{mn}$ and $B_{mn}$, we have
+    $f_{mn}=F\sqrt{\frac{x_{m}\tau_{n}}{c_{m}^{f}}},B_{mn}=F\sqrt{\frac{b_{mn}}{c_{m}^{B}\log_{2}(1+\varsigma_{mn})}}.$ (33)
+Appendix B Proof of $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}<0$ and $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}\partial p_{m^{\prime}n}}<0$
+Proof.
+First, we need to compute $\frac{\partial f_{mn}}{\partial p_{mn}}$, $\frac{\partial^{2}f_{mn}}{\partial p_{mn}^{2}}$, $\frac{\partial B_{mn}}{\partial p_{mn}}$, and $\frac{\partial^{2}B_{mn}}{\partial p_{mn}^{2}}$ to obtain the values of $V_{mn}^{\prime}$ and $V_{mn}^{\prime\prime}$.
+From (28), we can obtain $\frac{\partial f_{mn}}{\partial p_{mn}}$ and $\frac{\partial^{2}f_{mn}}{\partial p_{mn}^{2}}$, the steps of which are shown as follows.
+The first-order derivative of $\frac{\partial L_{m}}{\partial f_{mn}}=0$ with respect to $p_{mn}$ is expressed as $I_{mn}\frac{x_{m}\tau_{n}}{f_{mn}^{2}}-2(p_{mn}I_{mn}+\delta_{mn})\frac{x_{m}\tau_{n}}{f_{mn}^{3}}\cdot\frac{\partial f_{mn}}{\partial p_{mn}}=0$.
+Consequently, we can obtain
+$\frac{\partial f_{mn}}{\partial p_{mn}}=\frac{I_{mn}f_{mn}}{2(p_{mn}I_{mn}+\delta_{mn})}>0.$ (34)
+The second-order derivative of $\frac{\partial L_{m}}{\partial f_{mn}}=0$ with respect to $p_{mn}$ is expressed as
+$\frac{\partial^{2}f_{mn}}{\partial p_{mn}^{2}}=\frac{I_{mn}\frac{\partial f_{mn}}{\partial p_{mn}}(p_{mn}I_{mn}+\delta_{mn})-I_{mn}^{2}f_{mn}}{2(p_{mn}I_{mn}+\delta_{mn})^{2}},$ (35)
+by substituting (34) into (35), we have
+$\frac{\partial^{2}f_{mn}}{\partial p_{mn}^{2}}=\frac{-I_{mn}^{2}f_{mn}}{4(p_{mn}I_{mn}+\delta_{mn})^{2}}<0.$ (36)
+Likewise, we derive $\frac{\partial B_{mn}}{\partial p_{mn}}$ and $\frac{\partial^{2}B_{mn}}{\partial p_{mn}^{2}}$ with the following steps based on (30).
+The first-order derivative of $\frac{\partial L_{m}}{\partial B_{mn}}=0$ with respect to $p_{mn}$ is expressed as $I_{mn}\frac{b_{mn}}{B_{mn}^{2}\log_{2}(1+\varsigma_{mn})}-2(p_{mn}I_{mn}+\delta_{mn})\frac{b_{mn}}{B_{mn}^{3}\log_{2}(1+\varsigma_{mn})}\cdot\frac{\partial B_{mn}}{\partial p_{mn}}=0$.
+Accordingly, we have
+$\frac{\partial B_{mn}}{\partial p_{mn}}=\frac{I_{mn}B_{mn}}{2(p_{mn}I_{mn}+\delta_{mn})}>0.$ (37)
+Similarly, we obtain the second-order derivative of $\frac{\partial L_{m}}{\partial B_{mn}}=0$ with respect to $p_{mn}$ as
+$\frac{\partial^{2}B_{mn}}{\partial p_{mn}^{2}}=\frac{I_{mn}\frac{\partial B_{mn}}{\partial p_{mn}}(p_{mn}I_{mn}+\delta_{mn})-I_{mn}^{2}B_{mn}}{2(p_{mn}I_{mn}+\delta_{mn})^{2}},$ (38)
+by substituting (37) into (38), we have
+$\frac{\partial^{2}B_{mn}}{\partial p_{mn}^{2}}=\frac{-I_{mn}^{2}B_{mn}}{4(p_{mn}I_{mn}+\delta_{mn})^{2}}<0.$ (39)
+Based on (34), (36), (37), and (39), we can easily infer that $V_{mn}^{\prime}>0$ and $V_{mn}^{\prime\prime}<0$, as shown in (40).
+Then, we can obtain $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}\partial p_{m^{\prime}n}}<0$.
+For $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}$, We substitute the specific expressions for $V_{mn}^{\prime}$ and $V_{mn}^{\prime\prime}$ into (22) and obtain the following result as shown in (41).
+Next, we randomly choose a vector $\bm{h}\in\mathbb{R}^{M\times 1}$ with elements not all 00.
+For $\bm{\Lambda}_{n}$, it is easy to see that $\bm{h}^{T}\bm{\Lambda}_{n}\bm{h}=\sum_{m}\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}(h^{m})^{2}<0$.
+Furthermore, $\bm{h}^{T}\bm{H}_{n}\bm{h}=\frac{-\mu_{n}}{(1+\sum_{m}V_{mn})^{2}}\left(\sum_{m}h^{m}V_{mn}^{\prime}\right)^{2}<0$.
+Thus, we have $\bm{h}^{T}(\bm{\Lambda}_{n}+\bm{H_{n}})\bm{h}=\bm{h}^{T}\bm{\Lambda}_{n}\bm{h}+\bm{h}^{T}\bm{H}_{n}\bm{h}<0$, indicating that the utility function $\Psi_{n}$ is strictly concave.
+So far, the uniqueness of the Nash equilibrium [37] solution to the multi-MSP rewarding game $\Omega$ has been proved.
+$V_{mn}^{\prime}=\frac{\partial V_{mn}}{\partial p_{mn}}=I_{mn}\left(\frac{x_{m}\tau_{n}\log(1/\theta_{m})}{f_{mn}^{2}}\cdot\frac{\partial f_{mn}}{\partial p_{mn}}+\frac{b_{mn}}{B_{mn}^{2}\log_{2}(1+\varsigma_{mn})}\cdot\frac{\partial B_{mn}}{\partial p_{mn}}\right)>0,$ (40)
+$V_{mn}^{\prime\prime}=\frac{\partial^{2}V_{mn}}{\partial p_{mn}^{2}}=I_{mn}\left[\frac{-2x_{m}\tau_{n}\log(1/\theta_{m})}{f_{mn}^{3}}\cdot(\frac{\partial f_{mn}}{\partial p_{mn}})^{2}+\frac{x_{m}\tau_{n}\log(1/\theta_{m})}{f_{mn}^{2}}\cdot\frac{\partial^{2}f_{mn}}{\partial p_{mn}^{2}}\right.$
+$\left.-\frac{2b_{mn}}{B_{mn}^{3}\log_{2}(1+\varsigma_{mn})}\cdot(\frac{\partial B_{mn}}{\partial p_{mn}})^{2}+\frac{b_{mn}}{B_{mn}^{2}\log_{2}(1+\varsigma_{mn})}\cdot\frac{\partial^{2}B_{mn}}{\partial p_{mn}^{2}}\right]<0.$
+$\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}=-\left[\frac{4a_{1}I_{mn}^{2}\delta_{mn}+a_{1}I_{mn}^{3}p_{mn}}{4f_{mn}(\delta_{mn}+I_{mn}p_{mn})^{2}}+\frac{4b_{mn}I_{mn}^{2}\delta_{mn}+b_{mn}I_{mn}^{3}p_{mn}}{4a_{2}B_{mn}(\delta_{mn}+I_{mn}p_{mn})^{2}}+I_{mn}^{2}\frac{\left(\frac{a_{1}a_{2}I_{mn}B_{mn}+b_{mn}I_{mn}f_{mn}}{2a_{2}f_{mn}B_{mn}(\delta_{mn}+I_{mn}p_{mn})}\right)^{2}}{(1+\sum_{m}V_{mn})^{2}}\right.$ (41)
+$\left.+\frac{3I_{mn}^{3}(a_{1}a_{2}B_{mn}+b_{mn}f_{mn})}{4a_{2}f_{mn}B_{mn}(\delta_{mn}+I_{mn}p_{mn})^{2}(1+\sum_{m}V_{mn})}\right]<0,\text{where}\,a_{1}=x_{m}\tau_{n}\log(1/\theta_{m}),a_{2}=\log_{2}(1+\varsigma_{mn}).$

### v3

@@
 Abstract
-Timely updating of Internet of Things (IoT) data is crucial for immersive vehicular metaverse services.
-However, challenges such as latency caused by massive data transmissions, privacy risks associated with user data, and computational burdens on metaverse service providers (MSPs) hinder continuous collection of high-quality data.
-To address these issues, we propose an immersion-aware model trading framework that facilitates data provision for services while ensuring privacy through federated learning (FL).
-Specifically, we first develop a novel multi-dimensional metric, the immersion of model (IoM), which assesses model value comprehensively by considering freshness and accuracy of learning models, as well as the amount and potential value of raw data used for training.
-Then, we design an incentive mechanism to incentivize metaverse users (MUs) to contribute high-value models under resource constraints.
+Timely updating of Internet of Things (IoT) data is crucial for achieving immersion in vehicular metaverse services.
+However, challenges such as latency caused by massive data transmissions, privacy risks associated with user data, and computational burdens on metaverse service providers (MSPs) hinder the continuous collection of high-quality data.
+To address these challenges, we propose an immersion-aware model trading framework that enables efficient and privacy-preserving data provisioning through federated learning (FL).
+Specifically, we first develop a novel multi-dimensional evaluation metric for the immersion of models (IoM).
+The metric considers i) the freshness and accuracy of the local model, and ii) the amount and potential value of raw training data.
+Building on the IoM, we design an incentive mechanism to encourage metaverse users (MUs) to participate in FL by providing local updates to MSPs under resource constraints.
 The trading interactions between MSPs and MUs are modeled as an equilibrium problem with equilibrium constraints (EPEC) to analyze and balance their costs and gains, where MSPs as leaders determine rewards, while MUs as followers optimize resource allocation.
-Furthermore, considering dynamic network conditions and privacy concerns, we formulate the reward decisions of MSPs as a multi-agent Markov decision process.
-To solve this, we develop a fully distributed dynamic reward algorithm based on deep reinforcement learning, without accessing any private information about MUs and other MSPs.
-Experimental results demonstrate that the proposed framework outperforms state-of-the-art benchmarks, achieving improvements in IoM of $38.3\%$ and $37.2\%$, and reductions in training time to reach the target accuracy of $43.5\%$ and $49.8\%$, on average, for the MNIST and GTSRB datasets, respectively.
-Index Terms:
-Equilibrium problem with equilibrium constraints, immersion-aware, incentive mechanism, resource allocation, vehicular metaverse.
+To ensure privacy and adapt to dynamic network conditions, we develop a distributed dynamic reward algorithm based on deep reinforcement learning, without acquiring any private information from MUs and other MSPs.
+Experimental results show that the proposed framework outperforms state-of-the-art benchmarks, achieving improvements in IoM of $38.3\%$ and $37.2\%$, and reductions in training time to reach the target accuracy of $43.5\%$ and $49.8\%$, on average, for the MNIST and GTSRB datasets, respectively.
+These findings validate the effectiveness of our approach in incentivizing MUs to contribute high-value local models to MSPs, providing a flexible and adaptive scheme for data provisioning in vehicular metaverse services.
 I Introduction
@@
 Metaverse services are beginning to reveal their vast potential across a broad spectrum of industries, from gaming and autonomous driving to education and marketing.
-Notably, the application of vehicles within the metaverse [2, 3] has attracted significant interest, particularly for the enhanced traffic safety enabled by state-of-the-art augmented reality (AR) technologies.
-Market report [4] forecasted that the global automotive metaverse market will grow from $1.9$ billion in 2022 to $6.5$ billion by 2030.
-Automakers such as BMW have increasingly invested in AR technology.
+Notably, the application of vehicles within the metaverse has attracted significant interest, particularly for the enhanced traffic safety enabled by state-of-the-art augmented reality (AR) technologies.
+Market report [2] projected that the global automotive metaverse market would grow from \$$1.9$ billion in 2022 to \$$6.5$ billion by 2030.
+Automakers such as BMW are actively investing in AR technologies.
 As shown in Fig.
-2, augmented information makes driving safer by showing potential hazards hidden behind the vehicle in front of you.
+2, augmented information improves driving safety by displaying potential hazards obscured by the vehicle ahead.
 Moreover, Nissan’s [3] upcoming technology utilizes a 3D AR interface that merges the real world with the virtual world to provide the driver with augmented information about the surrounding area.
-In addition, Neuron mobility [^3] https://immersive-technology.com/augmentedreality/neuron-introduces-new-ar-parking-assistant/ has announced the launch of an innovative AR parking assistance system with the intention of improving passengers’ parking routines and trip-end experiences.
-The ability to capture information from the “real” world, particularly the ability to collect and process massive data from IoT devices, is the key to determining the success of immersive services (e.g., AR) in the vehicular metaverse.
-Meanwhile, the data must be processed and presented in a meaningful, responsive, and appropriately protected way.
-Technically, a high-quality experience with AR services relies on accurate detection and classification of real-world objects (e.g., cars and pedestrians) under complex conditions [6].
-To achieve this goal, sufficient valid data needs to be collected and processed in depth to detect and classify objects accurately.
+In addition, Neuron mobility [^3] https://immersive-technology.com/augmentedreality/neuron-introduces-new-ar-parking-assistant/ has announced the launch of an innovative AR parking assistance system to improve passengers’ parking routines and trip-end experiences.
+The ability to capture information from the “real” world, particularly the ability to collect and process massive amounts of data from IoT devices, is the key to determining the success of immersive services (e.g., AR) in the vehicular metaverse.
+Meanwhile, the data must be processed and presented in a meaningful and responsive manner with adequate privacy protection.
+Technically, a high-quality AR experience relies on accurately detecting and classifying real-world objects (e.g., cars and pedestrians) under complex conditions [4].
+To achieve this goal, sufficient valid data need to be collected and deeply processed to detect and classify objects accurately.
 Therefore, it is essential to focus on effectively collecting, processing, and protecting the data that supports a safer and more enjoyable driving experience.
@@
 The widely used data collection method, as adopted by Nissan [5] and studied in [6], involves gathering massive amounts of data through vehicle sensors, cameras, and roadside devices, and all data being processed centrally.
-While this approach has been effective for various applications, when it comes to the situation with multiple metaverse users (MUs) and metaverse service providers (MSPs) associated with different companies, the centralized data collection approach is not applicable and may lead to the following issues.
+While this approach is effective for various applications, when it comes to the situation with multiple metaverse users (MUs) and metaverse service providers (MSPs) associated with different companies, the centralized data collection approach is not applicable and may lead to the following issues.
 First, AR services in the vehicular metaverse need to be highly immersive so that MUs feel fully immersed in the rendered environment, such as visualized driving.
-However, the data synchronization will be hindered by the latency from massive real-time data updates under unstable and resource-limited communication conditions.
+However, data synchronization is hindered by the latency from massive real-time data updates under unstable and resource-limited communication conditions.
 Note that the value of real-time data diminishes over time [7].
@@
 Federated learning (FL) has been adopted in prior work [9, 10, 4, 11] to enable collaborative model training without sharing raw data (e.g., sensor/imaging data from vehicles).
-In this approach, learning models are uploaded by individual MUs to MSPs for AR services, rather than transmitting large volumes of data centrally, thus significantly reducing the communication burden.
-However, MUs are typically self-interested and are reluctant to share learning models with MSPs due to the additional computation, communication, and energy overhead.
-To overcome this, incentive mechanisms using strategies such as contract theory [14], Stackelberg game [15, 16] and multi-winner sealed-bid auction [17], have been proposed to encourage MUs to contribute models.
-However, existing studies fail to explicitly assess the model value from multiple dimensions, making it difficult for both MSPs and MUs to quantify the benefits of the model for MSPs.
+In this approach, local updates are uploaded by individual MUs to MSPs for AR services, rather than transmitting large volumes of data centrally, thus significantly reducing the communication burden.
+However, MUs are typically self-interested and are reluctant to share local models with MSPs due to the additional computation, communication, and energy overhead.
+To overcome this issue, incentive mechanisms using strategies such as contract theory [12], Stackelberg game [13, 14] and multi-winner sealed-bid auction [15], have been proposed to encourage MUs to contribute local models.
+However, existing studies fail to explicitly assess the model value from multiple dimensions, making it difficult for both MSPs and MUs to quantify the benefits of the models for MSPs.
 Furthermore, most of these solutions are designed for a single MSP and do not consider the joint optimization of MUs’ limited computational and communication resources.
@@
 To address the above research gaps, we propose an immersion-aware model trading framework for a multi-MSP, multi-MU vehicular metaverse.
-This framework is designed to incentivize MUs to become active contributors by providing models tailored to the specific needs of MSPs.
+This framework is designed to incentivize MUs to become active contributors by providing local models tailored to the specific needs of MSPs.
 However, MUs have different sampling costs and limited computational and communication resources, while MSPs differ in their model preferences and compete with each other.
@@
 This leads to two key research questions that underpin our work:
-  • How can MUs be effectively incentivized to contribute high-value learning models that benefit vehicular metaverse services for diverse MSPs?
-  • How can the dynamic competitive interactions among MSPs be modeled to achieve a trade-off in gains between MUs and MSPs, ensuring sustainable and stable trading of learning models in the vehicular metaverse?
+  • How can MUs be effectively incentivized to contribute high-value local models that benefit vehicular metaverse services for diverse MSPs?
+  • How can the dynamic competitive interactions among MSPs be modeled to achieve a trade-off in gains between MUs and MSPs, ensuring sustainable and stable trading of local models in the vehicular metaverse?
 Figure 2: The outline of an immersion-aware framework for FL-assisted vehicular metaverse.
@@
 2, our model trading framework consists of four components: FL mechanism, metric design, game modeling, and dynamic algorithms.
-Specifically, we first design a new metric called the immersion of the learning model (IoM) to evaluate the value of learning models contributed by MUs to MSPs.
-This metric jointly considers the freshness and accuracy of the learning model, as well as the amount and potential value of raw training data.
+Specifically, this work first designs a new metric called the immersion of the learning model (IoM) to evaluate the value of local models contributed by MUs to MSPs.
+This metric jointly considers the freshness and accuracy of the local model, as well as the amount and potential value of raw training data.
 Building on this metric, we propose an immersion-aware incentive mechanism that aligns the interests of both MUs and MSPs.
@@
   • Incentive mechanism design for immersion-aware model trading.
-From the perspectives of both MUs and MSPs, we propose an incentive mechanism that encourages MUs to contribute high-value learning models tailored to the specific demands of MSPs.
-To our knowledge, this is the first study focusing on incentive mechanism design for efficient and privacy-preserving data synchronization in multi-MU, multi-MSP vehicular metaverse environments.
+From the perspectives of both MUs and MSPs, we propose an incentive mechanism that encourages MUs to contribute high-value local models tailored to the specific demands of MSPs.
+To our knowledge, this is the first study focusing on incentive mechanism design for efficient and privacy-preserving data synchronization in multi-MU and multi-MSP vehicular metaverse environments.
   • Novel design of multi-dimensional metric.
-To quantify immersion enhancement provided by MUs for AR services, we design an immersion metric of the learning model integrating four critical dimensions: the freshness and accuracy of the learning model, as well as the amount and potential value of raw training data.
+To quantify immersion enhancement provided by MUs for AR services, we design an immersion metric of the local model integrating four critical dimensions: the freshness and accuracy of the local model, as well as the amount and potential value of raw training data.
 Freshness is captured through age of information (AoI), while potential value is evaluated by the difference between model predictions and true labels.
@@
   • Performance evaluation.
-We conduct extensive numerical simulations based on AR-related vehicle datasets to validate the efficacy and efficiency of MDDR and the proposed immersion-aware model trading framework.
+Extensive numerical simulations are conducted based on AR-related vehicle datasets to validate the efficacy and efficiency of MDDR and the proposed immersion-aware model trading framework.
 Numerical results show that our proposed mechanism improves the IoM by $38.3\%$ and $37.2\%$, and reduces the training time to reach the target accuracy by $43.5\%$ and $49.8\%$, on average, for the MNIST and GTSRB datasets, respectively, compared with state-of-the-art benchmarks.
@@
 Section II discusses the related work.
-In Section III, we present the system overview and design the immersion metric of the learning model.
+In Section III, we present the system overview and design the immersion metric of the local model.
 Section IV gives the game formulation, and Section V analyzes the existence of the equilibria at two levels.
 In Section VI, we give the detailed design of MDDR.
-Section VII shows numerical experiments to evaluate the framework performance, and finally Section VIII concludes the paper.
+Section VII shows numerical experiments to evaluate the framework performance, and finally, Section VIII concludes the paper.
 II Related Work
-In this section, we survey work related to our study in terms of edge-enabled vehicular metaverse services, FL for AR, and incentive mechanisms for data synchronization in the metaverse.
+In this section, we discuss the work related to our study in terms of edge-enabled vehicular metaverse services, FL for AR, and incentive mechanisms for data synchronization in the metaverse.
 II-A Edge-Enabled Vehicular Metaverse Services
-Vehicular metaverse services require ultra-low latency and high reliability to enable immersive experiences such as AR navigation and real-time digital twins.
+Vehicular metaverse services require ultra-low latency and high reliability to enable immersive experiences, such as AR navigation and real-time digital twins.
 To meet these demands, researchers have extensively explored edge-based strategies for task offloading and resource management.
 For instance, Feng et al.
-[19] proposed a resource allocation framework for AR-powered vehicular edge metaverses that aims to maximize overall system utility.
+[17] proposed an innovative resource allocation framework for AR-empowered vehicular edge metaverse, which significantly improves data utility and reduces vehicle energy consumption through multi-dimensional optimization.
 Similarly, Khan et al.
-[20] introduced a cooperative framework integrating task offloading, sensing, learning, and communication to reduce transmission energy consumption and latency in resource-constrained environments.
-Beyond static offloading, Chen et al.
-[21] developed a multi-agent deep reinforcement learning approach for dynamic avatar task migration.
-In terms of security and privacy, Kang et al.
+[18] introduced a cooperative framework integrating task offloading, sensing, learning, and communication to reduce both transmission energy consumption and latency in resource-constrained environments.
+To cope with dynamic network conditions, Tong et al.
+[19] proposed an attribute-aware auction mechanism for resource allocation, and applied a GPT-based deep reinforcement learning algorithm to adjust the auction clocks.
+Regarding security and privacy, Kang et al.
 [20] presented a cross-metaverse empowered dual pseudonym management framework to mitigate privacy leakage risks during the dynamic communications among vehicular edge metaverses.
-Moreover, an optimization framework for AR-assisted driving was proposed in [23] to enhance the robustness of metaverse maps against adversarial attacks.
-This framework maximizes the mean average precision (mAP) of adversarial patch detection while minimizing AR scene uplink latency and worst-case battery consumption.
-Recent studies have also explored applying FL to vehicular metaverse services as a means of preserving user privacy during distributed model training.
-These efforts, particularly in AR-enhanced metaverse environments, are discussed in the next section.
-To ensure service continuity and prevent user disengagement during unexpected failures, Qiu et al.
-[24] proposed deploying redundant backup virtual vehicle services and keeping the age of backup information to improve the reliability of services.
+Moreover, a novel four-layer security framework was designed to protect vehicular metaverse systems from data poisoning attacks [21].
+In addition to theoretical models and optimization frameworks for integrating autonomous vehicles into the metaverse, several studies have proposed practical decision-making approaches to evaluate and select among alternative implementation strategies.
+For example, Deveci et al.
+[22] proposed a framework that combines a fuzzy logarithmic weighting method with a similarity-based ranking technique to assess four traffic safety solutions enabled by the metaverse.
+Their results show that public transportation is the most practical and scalable option.
+Moreover, they applied a decision-making method based on Q-rung orthopair fuzzy sets to rank personal mobility alternatives, offering practical guidance for system planning in [23].
+Gokasar et al.
+[6] introduced a two-stage fuzzy decision-making framework to evaluate integration options for self-powered sensors in autonomous vehicles.
+The study concludes that using these sensors for real-time traffic management in the metaverse is the most effective solution.
 These works, from diverse perspectives, have significantly advanced the security, reliability, and performance of vehicular metaverse services, providing valuable insights for future research.
 Notably, distinct from existing studies, our work focuses on designing an incentive mechanism that transforms MUs into active contributors.
-These contributors efficiently provide high-value models to MSPs, thereby enhancing the overall service experience.
+These contributors efficiently provide high-value local models to MSPs, thereby enhancing the overall service experience.
 II-B Federated Learning for AR in the Metaverse
-Privacy preservation is a critical concern in vehicular metaverse services, particularly when training models using sensitive user data.
+Privacy preservation is critical in vehicular metaverse services, particularly when training models using sensitive user data.
 Federated learning (FL) has emerged as a promising technique to enable collaborative model training without sharing raw data.
@@
 To overcome this limitation, Zhou et al.
-[12] incorporated FL into mobile AR systems in the metaverse to enable collaborative model training.
-Their approach considers the trade-offs among energy consumption, execution latency, and model accuracy to meet varying demands across different application scenarios.
+[10] developed an FL-assisted MAR system via non-orthogonal multiple access for the metaverse.
+Furthermore, a resource allocation algorithm was designed to balance accuracy and energy consumption.
 To further improve learning efficiency, an adaptive and resource-efficient FL algorithm tailored for AR applications was proposed in [4].
@@
 [11] explored the integration of quantum computing with FL to provide a cost-efficient and adaptive solution for the dynamic nature of vehicular environments.
+Nadimi et al.
+[24] proposed a multi-modal, multi-task federated foundation model (FedFMs) architecture.
+It combines the privacy-preserving benefits of FL with the representation capabilities of foundation models to address sensor diversity, personalized interactions, and device heterogeneity in extended reality systems.
 Incorporating FL into AR services brings the benefits of reduced communication latency and privacy protection.
 However, the existing studies fail to consider the integration of FL into AR for the multi-MSP and multi-MU metaverse scenario, while overlooking the selfishness of MUs, i.e., whether they are willing to participate in FL learning.
+TABLE I: Incentive Mechanisms for Data Synchronization
+Ref.
+MSPs
+MUs
+Theory
+Method
+Main
+Metric
+Data
+Privacy
+[12] $\times$
+Contract theory
+under PT
+AoI
+[13] $\times$
+Dual
+game-theoretic
+Reputation
+scores
+[14] $\times$ Stackelberg game
+Model quality
+and latency
+[15] IISG and PMS-AM
+Model quality
+and freshness
+[25]
+Coalition formation,
+Stackelberg game
+Volume of
+sensing data
+$\times$
+[26]
+Deep learning
+based auction
+Quality,
+reliability, latency
+$\times$
+[27] Contract theory Computing delay $\times$
+[28] Stackelberg game
+Energy
+consumption
+$\times$
+[29]
+Multi-objective
+optimization
+Peak signal-
+to-noise ratio
+Our
+work
+EPEC IoM
 II-C Incentive Mechanism for Data Synchronization in the Metaverse.
-The quality of metaverse services depends heavily on massive sensing data from the real world.
-To facilitate this, incentive mechanisms are widely employed to motivate MUs to contribute data and participate in model construction, thereby enhancing their service experience.
+Immersive metaverse services depend on high-quality real-world sensing data.
+However, the process of collecting such data is often costly and may raise privacy concerns.
+To this end, various incentive mechanisms have been proposed to encourage MUs to contribute data and participate in model training.
 For instance, Zhang et al.
-[25] proposed a vehicle-assisted data sensing framework that incentivizes vehicles to contribute sensing data to metaverse service, improving the driving experience for vehicle users.
-Similarly, Lin et al.
-[26] introduced a contract-based incentive design aimed at optimizing the long-term profits of digital twin service providers, addressing service congestion caused by stochastic demand responses.
+[25] introduced a vehicle-assisted data sensing framework that incentivizes vehicles to upload sensing data, enhancing the driving experience.
 Xu et al.
-[27] developed a metaverse-based unmanned aerial vehicle (UAV) swarm system that uses digital twins and semantic communication to enhance data synchronization and reduce communication latency.
-Moreover, Lotfi et al.
-[28] introduced VMGuard, a novel four-layer security framework for the vehicular metaverse that protects against data poisoning attacks by using a reputation-based incentive mechanism.
-The framework leveraged user feedback and subjective logic modeling to assess the trustworthiness of devices.
-However, while these works successfully incentivize MUs to contribute raw data, they also pose a risk to user privacy due to potential data leakage.
-To address the privacy concern, Kang et al.
-[14] developed a cross-chain FL framework with an AoI-based contract theory model under prospect theory to incentivize data sharing.
+[26] developed a UAV swarm system using digital twins and semantic communication to improve data synchronization and reduce latency.
+Lin et al.
+[27] designed an incentive-based congestion control scheme for digital twin edge networks, which addresses the stochastic service demand and the long-term provider profit by combining Lyapunov optimization and contract theory.
+Zhao et al.
+[28] proposed a physical reality incentive mechanism based on a two-stage Stackelberg game to motivate MUs to collect physical reality sensing data for the metaverse.
+Although these studies successfully encourage raw data contributions, they also pose a risk to user privacy due to potential data leakage.
+To address privacy concerns, Kang et al.
+[12] developed a cross-chain FL framework with an AoI-based contract theory model under prospect theory (PT) to incentivize data sharing.
 They utilized a hierarchical cross-chain architecture with a main chain and multiple subchains to perform decentralized, privacy-preserving, and secure data training in both virtual and physical spaces.
 Baccour et al.
-[15] proposed a dual game-theoretic framework for federated meta-learning (FML) in metaverse services.
-They introduced a reputation system and a Stackelberg game-based incentive mechanism to enhance data privacy, minimize energy costs, and improve FML efficacy, outperforming traditional clustering methods in training performance and metaverse utility.
+[13] designed a dual game-theoretic framework for federated meta-learning (FML), incorporating a reputation system and Stackelberg game to enhance privacy, reduce energy consumption, and improve model performance.
 Li et al.
-[16] designed a satisfaction metric that considers data size, AoI, and training latency.
+[14] designed a new metric named “Satisfaction” to balance training latency and model quality.
 They integrated this metric into utility functions to incentivize node participation in FL while optimizing resource allocation.
 However, these works are designed for a single MSP scenario, limiting scalability and applicability in multi-server scenarios commonly encountered in large-scale metaverse environments.
-Building on the concept of multiple MSPs, Zhang et al.
-[17] proposed a privacy-preserving auction mechanism for learning model as an NFT in Blockchain-driven Metaverse.
-They employed an imperfect information Stackelberg game to optimize the strategies of MUs and MSPs, and introduced a multi-winner sealed-bid auction mechanism to enhance the quality of FL-NFT.
-However, this approach primarily focuses on modeling NFTs and overlooks the joint optimization of computational and communication resources.
-Moreover, it fails to evaluate the model’s value from multiple dimensions.
-Inspired by the above existing works, this study focuses on multi-MSP and multi-MU scenarios in vehicular metaverse services.
-By assigning MUs the dual roles of service consumers and model contributors, we build a low-cost, privacy-preserving data support framework.
-This allows MUs to enjoy immersive in-vehicle services (e.g., AR navigation, digital twin interactions) while participating as distributed model contributors in service optimization.
-This framework continuously incentivizes MUs to contribute high-value learning models for metaverse services, thereby enhancing the overall service experience.
-Figure 3: Workflow of the immersion-aware model trading framework [^5] Image source: https://mixed-news.com/en/google-transforms-google-maps-into-the-backbone-of-our-ar-future/, https://about.facebook.com/meta, https://www.vanarama.com/blog/cars/4-ways-augmented-reality-will-revolutionise-the-automotive-industry..
+To fill this gap, Zhang et al.
+[15] employed an imperfect information Stackelberg game (IISG) to optimize the strategies of MUs and MSPs, and introduced a privacy-preserving multi-winner sealed-bid auction mechanism (PMS-AM) to enhance the quality of non-fungible token (NFT) with FL assistance.
+Esmail et al.
+[29] proposed a metaverse FL and semantic communication (MFSC) framework that matches IoT device capabilities to training tasks, optimizing digital twin quality through adaptive resource allocation.
+Nonetheless, existing incentive criteria are often simplistic and fail to evaluate the values of contributed models comprehensively.
+Unlike these prior works, our study focuses on vehicular metaverse services in multi-MSP and multi-MU scenarios.
+We propose a novel metric, IoM, which evaluates model value from multiple dimensions.
+Based on this metric, we design an incentive mechanism that better motivates MUs to contribute high-value local models, thereby improving the quality and user experience of metaverse services.
+A detailed comparison of these approaches is provided in Table I, highlighting the key contributions of this study.
+Figure 3: Workflow of the immersion-aware model trading framework [^5] Image source: https://www.vanarama.com/blog/cars/4-ways-augmented-reality-will-revolutionise-the-automotive-industry..
 III System Model
 We first present the system overview of the immersion-aware model trading framework in Section III-A.
-Next, we introduce the FL mechanism adopted by our framework in Section III-B and design the immersion metric of the learning model in Section III-C.
+Next, we introduce the FL mechanism adopted by our framework in Section III-B and design the immersion metric of the local model in Section III-C.
 III-A System Overview
@@
 5 illustrates the workflow of our immersion-aware model trading framework, designed to enhance the immersive and interactive AR experiences of MUs.
-MUs can contribute learning models to the AR services of MSPs for better immersive experiences.
-For example, services such as object detection can be improved by training on images of streets, pedestrians, and so on, captured by vehicles’ cameras.
-In this way, AR services can be better applied to visible driving, with more accurate and timely hazard warnings from windscreens.
+MUs can contribute local models to the AR services of MSPs for better immersive experiences.
+For example, services such as object detection can be improved by training on street and pedestrian images captured by vehicles’ cameras.
+In this way, AR services can be applied more effectively to visible driving, with more accurate and timely hazard warnings from windscreens.
 The framework mainly consists of MSPs owned by different companies (e.g., Meta, Nissan, and Google), an infrastructure layer, and an interaction layer with the MUs that enjoy vehicular metaverse services.
-MSPs are powered by various technologies supporting vehicular metaverse services, such as AI engines, AR/VR engines, metaverse trading engines.
+MSPs are powered by various technologies supporting vehicular metaverse services, such as AI engines, AR/VR engines, and metaverse trading engines.
 The infrastructure layer with base stations (equipped with caches and edge servers) provides the basis of 5G/6G communication services for the interactions among MUs and MSPs.
 MUs play essential roles as both consumers and contributors to vehicular metaverse services.
-TABLE I: Definitions of Notation
+TABLE II: Definitions of Notation
 Notation Definition
@@
 $\theta_{m}$ MU $m$’s local accuracy threshold
-The framework is a multi-engine configuration, and this paper focuses on two components: the trading engine and the AI engine of the MSP.
+The framework adopts a multi-engine architecture, and this paper focuses on two components: the trading engine and the AI engine of the MSP.
 The trading engine is responsible for incentivizing MUs to participate in mutually beneficial collaborations with MSPs, providing them with a trading platform to make their decisions (i.e., computational and communication resource allocation of MUs and reward decisions of MSPs).
-Then, each MU deciding to contribute utilizes the FL mechanisms integrated into AI engines to provide learning models for MSPs under the strategic guidance of the trading results.
+Then, each MU deciding to contribute utilizes the FL mechanisms integrated into AI engines to provide local models for MSPs under the strategic guidance of the trading results.
 The specific workflow of the framework involves the following two phases.
@@
 Then, the trading engines of MSPs determine their digital currency prices per IoM, i.e., rewards to MUs, and broadcast them to MUs (②).
-Next, MUs determine their allocation of computational and communication resources for contributing learning models to various MSPs based on the IoM, costs, and rewards given by MSPs (③).
+Next, each MU determines its allocation of computational and communication resources for contributing local models to various MSPs based on the IoM, costs, and rewards given by MSPs (③).
 Based on the responses from MUs, the trading engines of MSPs adjust the digital currency prices to maximize their own utilities.
-Steps ② and ③ are repeated until an agreement is reached among MUs and MSPs.
-Phase II (FL process): Each MU, guided by decisions obtained through tradings, allocates computational resources to local training based on different global models from MSPs and generates local learning models for different MSPs (④).
-Then, MUs upload their updated learning models to the corresponding AI engines of MSPs by consuming communication resources (⑤).
-Following this, AI engines aggregate models from all MUs to update their global models and transmit them to their AR/VR engines for utilization (⑥).
+Steps ② and ③ iterate until an agreement is reached between MUs and MSPs.
+Phase II (FL process): Each MU, guided by decisions obtained through tradings, allocates computational resources to local training based on different global models from MSPs and generates local models for different MSPs (④).
+Then, MUs upload their local updates to the corresponding AI engines of MSPs by consuming communication resources (⑤).
+Following this, AI engines aggregate local updates from all MUs to update their global models and transmit them to their AR/VR engines for utilization (⑥).
 Finally, the AR/VR engines provide enhanced AR services to MUs (⑦).
 Steps ④-⑦ are repeated until the guidance time $T$ ends.
-To ease the presentation, we summarize some important notations in Table I.
+To ease the presentation, we summarize some important notations in Table II.
 III-B FL Framework Mechanism Adopted
-Here, we consider a set of =$\mathcal{M}=\left\{1,\ldots,M\right\}$ MUs with the capacities of local computation and a set of $\mathcal{N}=\left\{1,\ldots,N\right\}$ MSPs with FL synchronous tasks [^7] That is, the update time of the global model is limited by the slowest MU, i.e., the AI engine of the MSP must wait to receive updated models from all MUs before performing model aggregation.
+We consider a set of $\mathcal{M}=\left\{1,\ldots,M\right\}$ MUs equipped with local computational capabilities and a set of $\mathcal{N}=\left\{1,\ldots,N\right\}$ MSPs with FL synchronous tasks [^7] That is, the update time of the global model is limited by the slowest MU, i.e., the AI engine of the MSP must wait to receive local updates from all MUs before performing model aggregation.
 The advantages of synchronous FL are high model accuracy and fast convergence [30]..
 Each MSP initiates an FL synchronous task with a virtual deadline $\tau_{n}\ (n\in\mathcal{N})$.
-The MUs can join FL to generate local learning models and contribute to enhancing vehicular metaverse services.
+The MUs can participate in FL to generate local models and contribute to enhancing vehicular metaverse services.
 As shown in Fig.
-4, MU $m$ provides learning models to different MSPs simultaneously for a given time period $T$, meaning multiple FL tasks can be processed in parallel.
-Moreover, the data collected at the $(r-1)$-th round is used as the local data for training at the $r$-th round.
-Without loss of generality, we consider that the trading decisions among MUs and MSPs are made at the beginning of the period $T$.
+4, MU $m$ provides local models to different MSPs simultaneously for a given time period $T$, meaning multiple FL tasks can be processed in parallel.
+Moreover, the data collected at the (r1)(r-1)-th round are used for local training at the $r$-th round.
+Without loss of generality, we consider that the trading decisions among MUs and MSPs are made at the beginning of period $T$.
 Figure 4: Illustration of the FL mechanism with $n$ tasks for MU $m$.
-In detail, let $X_{mn}$ be the set of data samples for task $n$ of MU $m$ as a set of input-output pairs $\sum_{i=1}^{|X_{mn}|}(y_{i},z_{i})$, where $y_{i}$ is an input sample vector with $d$ features, $z_{i}$ is the labeled output (i.e., ground truth) value of sample $y_{i}$, and $|X_{mn}|$ is the size of set $X_{mn}$.
+Let $X_{mn}=\{(y_{i},z_{i})\}_{i=1}^{|X_{mn}|}$ be the set of input-output pairs sampled from MU $m$ for task $n$, where $y_{i}$ is the input with $d$ features and $z_{i}$ is its corresponding ground-truth label.
+Here, $|X_{mn}|$ denotes the size of the set $X_{mn}$.
 The data, such as images of streets and pedestrians, can be generated by high-definition cameras inside and outside the vehicles.
@@
 The number of local training rounds depends on its local accuracy threshold.
-  • Uploading and Aggregation: MUs transmit their updated local model $\jmath_{mn}^{(r)}$ to the AI engine of MSP $n$.
-After receiving all the updated models, the MSP $n$’s AI engine aggregates them through the following weighted average:
+  • Uploading and Aggregation: MUs transmit their local updates $\jmath_{mn}^{(r)}$ to the AI engine of MSP $n$.
+Upon receiving all the local updates, the MSP $n$’s AI engine aggregates them through the following weighted average:
     $\jmath_{n}^{(r+1)}$ $=\jmath^{(r)}_{n}+\sum_{m=1}^{M}\frac{|X_{mn}^{(r)}|}{|X_{n}^{(r)}|}(\jmath_{mn}^{(r)}-\jmath_{n}^{(r)}),$ (1)
@@
 The global rounds are iterated until a specific requirement is met, such as reaching a certain level of accuracy or a deadline.
-III-C Immersion Metric of Learning Model
-MSPs determine the rewards for MUs based on the immersion of the learning model (IoM) metric.
-Correspondingly, MUs need to decide the IoMs of the models provided to MSPs by allocating their computational and communication resources to maximize the benefits.
+III-C Immersion Metric of Local Model
+MSPs determine the rewards for MUs based on the immersion of the local model (IoM) metric.
+Correspondingly, MUs need to decide the IoMs of local models provided to MSPs by allocating their computational and communication resources to maximize the benefits.
 This process highlights the key role of “IoM” in the interactions between the two parties, serving as both an evaluation criterion and a basis for decision-making.
-Unlike traditional FL and incentive mechanisms, in vehicular metaverse scenarios with immersion requirements, both the contribution and freshness of the learning model need to be considered [^8] For example, an MSP may receive fresher learning models but with lower contributions, or less fresh learning models but with higher contributions.
+Unlike traditional FL and incentive mechanisms, vehicular metaverse scenarios with immersion requirements necessitate consideration of both the contribution and freshness of the local model [^8] For example, an MSP may receive fresher local updates but with lower contributions, or less fresh local updates but with higher contributions.
 , which will collectively affect the immersive experience of AR services, i.e., whether virtual objects can be placed in the physical world accurately and promptly.
-With the above consideration, we propose the IoM, a metric that jointly considers the contribution prediction $I_{mn}$ and the age of information (AoI) $\Delta_{mn}$ [30] to measure the value of a learning model that MU $m$ brought to the vehicular metaverse service of MSP $n$, denoted as
+Driven by the above considerations, we design the IoM based on four key dimensions: the freshness and accuracy of the local model, as well as the amount and potential value of raw training data.
+The freshness is captured through age of information (AoI) $\Delta_{mn}$ [31], while the other three dimensions are jointly reflected in contribution prediction $I_{mn}$.
+To measure the value of a local model that MU $m$ brought to the vehicular metaverse service of MSP $n$, we define the IoM as
 $V_{mn}=I_{mn}(\tau_{n}-\overline{\Delta}_{mn}),$ (2)
 where the components are detailed as follows. In the rest of this paper, we refer to the trading engine of MSP as MSP.
-III-C1 Contribution Prediction $I$
-The contribution prediction $I_{mn}$ from MU $m$ to MSP $n$ is determined by the accuracy $\theta_{m}$ of the learning model, the total amount of training data $\lfloor\frac{T}{\tau_{n}}\rfloor|X_{mn}|$, and the potential value $\omega_{mn}$ of local data.
-Prior work [31] characterized the contribution of a learning model in terms of a logarithmic function of the amount of local training data.
-However, evaluating the contribution solely by the training data size is one-sided, as there may be a large amount of redundant data or the validity of the learning model cannot be guaranteed.
+III-C1 Contribution Prediction II
+The contribution prediction $I_{mn}$ from MU $m$ to MSP $n$ is determined by the accuracy $\theta_{m}$ of the local model, the total amount of training data $\lfloor\frac{T}{\tau_{n}}\rfloor|X_{mn}|$, and the potential value $\omega_{mn}$ of local data.
+Prior work [32] characterized model contributions using a logarithmic function of the amount of local training data.
+However, evaluating the contribution solely by the training data size is one-sided, as there may be a large amount of redundant data or the validity of the local model cannot be guaranteed.
 It is more practical to comprehensively evaluate the amount of training data, model accuracy, and the potential value of local data; thus, we denote $I_{mn}$ as
@@
 $\theta_{m}\in\left(0,1\right)$ characterizes the accuracy of local training as decided by the MU.
-Here, $\theta\rightarrow 0$ means that a high-accuracy learning model is available, while $\theta\rightarrow 1$ implies that the accuracy of the learning model is low.
+Here, $\theta\rightarrow 0$ means that a high-accuracy local model is available, while $\theta\rightarrow 1$ implies that the accuracy of the local model is low.
 The local data potential value $\omega_{mn}$ describes the difference between model predictions and true labels, defined as
@@
 Therefore, updating the model by training on these data may lead to better performance of the global model.
-If the value of $\omega_{mn}$ is small, it means that the current data has little potential to further improve the global model since the knowledge of the data has already been learned or the data size is small.
+If the value of $\omega_{mn}$ is small, the current data has little potential to improve the global model further since the knowledge of the data has already been learned or the data size is small.
 Note that to shorten the decision time, we use the potential value of the initial local data to be a proxy for the average potential value of all data within $T$.
 This is a reasonable approach when $T$ is not too long.
-Figure 5: The AoI evolution of the models sent from MU $m$ to MSP $n$.
+Figure 5: The AoI evolution of the local updates sent from MU $m$ to MSP $n$.
 III-C2 Age of Information $\Delta$
@@
 5.
-Here, the filled circles denote the time instances when MU $m$ starts training the learning model.
-The intersections show the time instances when MSP $n$ receives the corresponding updated model, at which point the AoI drops to the lowest.
+Here, the filled circles denote the time instances when MU $m$ starts training the local model.
+The intersections show the time instances when MSP $n$ receives the corresponding local updates, at which point the AoI drops to the lowest.
 We denote that at time $t$, the generation time of the last update received by MSP $n$ from MU $m$ is $s(t)$.
-Then, the instantaneous AoI of MU $m$’s learning model measured at MSP $n$, $\Delta_{mn}(t)$, is
+Then, the instantaneous AoI of MU $m$’s local model measured at MSP $n$, $\Delta_{mn}(t)$, is
 $\Delta_{mn}(t)=t-s(t),$ (5)
-which shows the elapsed time since the generation of the latest updated model.
-A smaller instantaneous AoI means that the latest model received is fresher at that time.
+which shows the elapsed time since the generation of the latest local updates. A smaller instantaneous AoI means that the latest local updates received are fresher at that time.
 To obtain a comprehensive understanding of the overall freshness of model updates over a certain period, we use average AoI to evaluate the freshness of the models.
-With the aid of Figure 5, the time-averaged AoI of the learning model from MU $m$ to MSP $n$ can be expressed as
-[…]
+With the aid of Figure 5, the time-averaged AoI of the local model from MU $m$ to MSP $n$ can be expressed as
+$\overline{\Delta}_{mn}$ $=\mathop{\lim}\limits_{Z\to\infty}\frac{1}{Z}\int_{0}^{Z}{{\Delta_{mn}}(t)}\,dt=\mathop{\lim}\limits_{Z\to\infty}\sum_{r=1}^{Z}\frac{{Q_{mn}}^{(r)}}{Z}$ (6)
 $=\frac{1}{2}\tau_{n}+T_{mn}^{c}+T_{mn}^{t},$
-where ${Q_{mn}}^{(r)}$ is the $r$-th trapezium under the curve. A lower average AoI indicates that the updated models are generally fresher over a long period.
-Since MU $m$ constantly collects data throughout the FL process, the amount of training data collected for training in each round is $|X_{mn}|=x_{m}\tau_{n}$, where $x_{m}$ denotes the data (number of floats) collected per unit time.
+where ${Q_{mn}}^{(r)}$ is the $r$-th trapezium under the curve. A lower average AoI indicates that the local updates are generally fresher over a long period.
+Since MU $m$ continuously collects data during the FL process, the amount of training data used in each round is $|X_{mn}|=x_{m}\tau_{n}$, where $x_{m}$ denotes the data (number of floats) collected per unit time.
 The cumulative time $T_{mn}^{c}$ for local training is determined by the amount of training data, computational resource $f_{mn}$ allocated to MSP $n$, and its local accuracy threshold $\theta_{m}$, which is equivalent to
@@
 A smaller value of $\theta$ indicates higher accuracy, but leads to a higher MU cost, i.e., the number of local iterations, which is upper bounded by $\log(1/\theta_{m})$ [34].
-Furthermore, the upload time $T_{mn}^{t}$ is directly related to the number of model parameters $b_{mn}$ and the resource $B_{mn}$ (i.e., uplink bandwidth) used to communicate between MU $m$ and MSP $n$, defined as
+Furthermore, the upload time $T_{mn}^{t}$ depends on the size of local updates $b_{mn}$ (in bits) and the resource $B_{mn}$ (in Hz) allocated for communication between MU $m$ and MSP $n$ (i.e., uplink bandwidth), defined as
 $T_{mn}^{t}=\frac{b_{mn}}{B_{mn}\log_{2}(1+\varsigma_{mn})},$ (8)
@@
 $g_{mn}$ and $p^{t}_{m}$ are the corresponding channel gain and the transmission power of MU $m$, respectively.
-$\sigma^{2}$ is additive white Gaussian noise (AWGN).
+$\sigma^{2}$ is the power of additive white Gaussian noise (AWGN).
 Finally, by substituting (7) and (8) into (6), $\overline{\Delta}_{mn}$ can be expressed as
@@
 Problem 1.
-$\max\,\,\Phi_{m}(\bm{f}_{m},\bm{B}_{m})=\sum_{n=1}^{N}(p_{mn}V_{mn}-C_{mn}),$ (11)
-$s.t.$ $C1:\sum_{n=1}^{N}f_{mn}<f_{m}^{max},\,\,\sum_{n=1}^{N}B_{mn}<B_{m}^{max},$
-$C2:\log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}\log_{2}(1+\varsigma_{mn})}\leq\tau_{n},$
+$\max\,\,\Phi_{m}(\bm{f}_{m},\bm{B}_{m})=\sum_{n=1}^{N}(p_{mn}V_{mn}-C_{mn})$ (11)
+$\mathrm{s.t.}$ $\ C1:\sum_{n=1}^{N}f_{mn}<f_{m}^{max},\,\,\sum_{n=1}^{N}B_{mn}<B_{m}^{max},$
+$\ C2:\log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}\log_{2}(1+\varsigma_{mn})}\leq\frac{1}{2}\tau_{n},$
 $\ C3:\frac{S_{m}}{f^{max}_{m}-\sum_{n=1}^{N}f_{mn}}<T^{req},$
-where $S_{m}$ indicates the minimum computational resource required to enjoy other basic services.
-Constraints C1 and C3 indicate that the MU’s total computational and communication resources are limited and cannot be fully used to contribute learning models, given the MU’s demand for basic services.
-Constraint C2 ensures that the total time for local training and uploading cannot exceed the time constraint $\tau_{n}$ for each round.
+where $p_{mn}$ is the reward offered by MSP $n$ to MU $m$ for their contribution.
+$S_{m}$ indicates the minimum computational resource required to enjoy other basic services.
+Constraints C1 and C3 indicate that the MU’s total computational and communication resources are limited and cannot be fully used to contribute local models, given the MU’s demand for basic services.
+Constraint C2 ensures that $V_{mn}$ is non-negative, guaranteeing that the value derived from the task meets the MSP’s minimum requirements.
+Specifically, the total time for local training and uploading in each round cannot exceed the time constraint $\frac{1}{2}\tau_{n}$.
 The goal of each MU is to choose the optimal allocation of computational $\bm{f}_{m}$ and communication $\bm{B}_{m}$ resources to maximize its utility.
-Note that $V_{mn}$ must be greater than 00.
-This is because if $V_{mn}\leq 0$ and MU $m$ is involved in contributing learning models, the utility is calculated as $p_{mn}V_{mn}-C_{mn}\leq 0$, resulting in MU $m$ choosing not to participate in the trading.
 IV-B Utility of MSP
-Each MSP has a gain function $\psi$ associated with the IoM to measure the benefits of all the learning models from MUs to the service.
+Each MSP has a gain function $\psi$ associated with the IoM to quantify the cumulative benefit of all local models contributed by MUs.
 In this paper, we adopt $\psi_{n}=\mu_{n}\ln(1+\sum_{m=1}^{M}V_{mn})$ as a monotonically increasing, differentiable, strictly concave IoM function, which is a simplification of the widely adopted function [36].
@@
 $\max\,\,\Psi_{n}(\bm{p}_{n},\bm{p}_{\textrm{-}n})=\psi_{n}-\sum_{m=1}^{M}p_{mn}V_{mn},$ (12)
-$s.t.\,\,p_{mn}>0,n\in\mathcal{N},m\in\mathcal{M},$
+$\mathrm{s.t.}\ p_{mn}>0,n\in\mathcal{N},m\in\mathcal{M},$
 where $\mu_{n}$ refers to the profit conversion coefficient [37] from IoM, which is adjusted according to the AR services offered by the different MSPs.
@@
 This thereby leads to reward competition among MSPs.
-Therefore, the optimization among MSPs can be considered as a non-cooperative game, termed a multi-MSP rewarding game $\Omega$ as follows.
+Therefore, the optimization among MSPs can be considered a non-cooperative game, termed a multi-MSP rewarding game $\Omega$, as follows.
 Definition 1.
@@
 $\Phi_{m}((\bm{f}_{m}^{*},\bm{B}_{m}^{*}),\bm{p}_{n}^{*})\geq\Phi_{m}((\bm{f}_{m},\bm{B}_{m}),\bm{p}_{n}^{*}),\forall m\in\mathcal{M},$ (13)
-$\Psi_{n}\left(\bm{f}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{B}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{p}_{n}^{*},\bm{p}_{-n}^{*}\right)\geq$
-$\Psi_{n}\left(\bm{f}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{B}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{p}_{n},\bm{p}_{-n}^{*}\right),\forall n\in\mathcal{N},$
-where $\bm{p}_{-n}^{*}$ denotes the optimal rewards offered by other MSPs.
+$\Psi_{n}\left(\bm{f}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{B}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{p}_{n}^{*},\bm{p}_{-n}^{*}\right)$
+$\geq\Psi_{n}\left(\bm{f}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{B}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{p}_{n},\bm{p}_{-n}^{*}\right),\forall n\in\mathcal{N},$
+where $\bm{p}_{-n}^{*}$ denotes the optimal reward vector for all MSPs except $n$.
 In summary, the MSPs’ optimization problems are formulated as the following EPEC problems:
-$\max_{p_{n}}\,\,\Psi_{n}=\psi_{n}-\sum_{m=1}^{M}\bigg{[}p_{mn}I_{mn}\bigg{(}\frac{1}{2}\tau_{n}$ (14)
-$-\frac{x_{m}\tau_{n}\log(1/\theta_{m})}{f_{mn}^{*}}-\frac{b_{mn}}{B_{mn}^{*}\log_{2}(1+\varsigma_{mn})}\bigg{)}\bigg{]},$
-[…]
+$\max_{p_{n}}\,\,\Psi_{n}=\psi_{n}-\sum_{m=1}^{M}\bigg[p_{mn}I_{mn}\bigg(\frac{1}{2}\tau_{n}-\frac{x_{m}\tau_{n}\log(1/\theta_{m})}{f_{mn}^{*}}$ (14)
+$\qquad\qquad\qquad\qquad\qquad-\frac{b_{mn}}{B_{mn}^{*}\log_{2}(1+\varsigma_{mn})}\bigg)\bigg],$
+$\mathrm{s.t.}\ \left\{\begin{aligned} &p_{mn}>0,\ n\in\mathcal{N},\ m\in\mathcal{M},\\ &(\bm{f}_{m}^{*},\bm{B}_{m}^{*})=\arg\max\Phi_{m}(\bm{f}_{m},\bm{B}_{m}),\\ &\qquad\qquad\quad\quad\mathrm{s.t.}\ C1,C2,C3.\end{aligned}\right$.
 To investigate the above EPEC, we address the lower level (Problem 1) and the upper level (Problem 2) by using the backward induction methods in the following section.
@@
 Below, we analyze and derive the unique optimal allocation.
-Definition 3.
-A maximization problem has a unique global optimum if the following conditions are satisfied [37]: (1) The objective function is strictly concave; (2) All constraints (equality and inequality) are affine; and (3) The strategy space is compact and convex.
 Theorem 1.
@@
 where $h^{B}_{mn}=\frac{2b_{mn}p_{mn}I_{mn}}{B_{mn}^{3}\log_{2}(1+\varsigma_{mn})}$.
-We find that $\bm{H}^{f}_{m}$ and $\bm{H}^{B}_{m}$ are both negative semi-definite.
-Hence, $\bm{H}_{m}$ is negative semi-definite, implying $\Phi_{m}(\bm{f}_{m},\bm{B}_{m})$ is strictly concave.
-Since the constraints of Problem 1 are affine and the strategy space is closed and bounded, Definition 3 ensures the existence of a unique global optimum $(\bm{f}_{m}^{*},\bm{B}_{m}^{*})$.
+Since $\bm{H}^{f}_{m}$ and $\bm{H}^{B}_{m}$ are diagonal matrices with strictly negative diagonal elements, they are negative definite.
+Consequently, the block-diagonal Hessian $\bm{H}_{m}$ is negative definite, implying $\Phi_{m}(\bm{f}_{m},\bm{B}_{m})$ is strictly concave and continuous.
+Moreover, the objective function $\Phi_{m}$ tends to -$\infty$ as any variable approaches the boundary of the feasible set (e.g., $f_{mn}\to 0^{+}$ and $B_{mn}\to 0^{+}$, or $f_{mn}\to\infty$ and $B_{mn}\to\infty$).
+Hence, the maximum is necessarily attained in the interior.
+Combined with strict concavity over the convex feasible region defined by constraints C1-C3, this ensures that Problem 1 admits a unique global optimum $(\bm{f}_{m}^{*},\bm{B}_{m}^{*})$ (see Section IV in [38]).
 ∎
-Consequently, MU $m$ has a unique best-response resource allocation for each MSP by solving Problem 1. From the Karush–Kuhn–Tucker (KKT) conditions, we establish the following proposition.
+Consequently, MU $m$ has a unique best-response resource allocation for each MSP by solving Problem 1. From the Karush–Kuhn–Tucker (KKT) conditions, we derive the following proposition.
 Proposition 1.
 Given the digital currency price offered by the MSPs and the Lagrange multipliers induced by the constraints, the optimal computational resource $f_{mn}$ allocated for MSP $n$ by MU $m$ satisfies:
-$f_{mn}^{*}=\begin{cases}\sqrt{\frac{p_{mn}I_{mn}x_{m}\tau_{n}}{c_{m}^{f}}},&\text{ $p_{mn}\geq\frac{F^{2}}{I_{mn}}$},\\ F\sqrt{\frac{x_{m}\tau_{n}}{c_{m}^{f}}},&\text{\,\,$0<p_{mn}<\frac{F^{2}}{I_{mn}}$},\\ 0,&\,\,\text{otherwise},\end{cases}$ (18)
+$f_{mn}^{*}=\begin{cases}\sqrt{\frac{p_{mn}I_{mn}x_{m}\tau_{n}}{c_{m}^{f}}},&p_{mn}\geq\frac{F_{mn}^{2}}{I_{mn}},\\ F_{mn}\sqrt{\frac{x_{m}\tau_{n}}{c_{m}^{f}}},&0<p_{mn}<\frac{F_{mn}^{2}}{I_{mn}},\\ 0,&\text{otherwise},\end{cases}$ (18)
 and the optimal communication bandwidth $B_{mn}$ is:
-$B_{mn}^{*}=\begin{cases}\sqrt{\frac{p_{mn}I_{mn}b_{mn}}{c_{m}^{B}\log_{2}(1+\varsigma_{mn})}},&\text{$p_{mn}\geq\frac{F^{2}}{I_{mn}}$},\\ F\sqrt{\frac{b_{mn}}{c_{m}^{B}\log_{2}(1+\varsigma_{mn})}},&\text{\,\,$0<p_{mn}<\frac{F^{2}}{I_{mn}}$},\\ 0,&\text{otherwise},\end{cases}$ (19)
-where $F=\log(1/\theta_{m})\sqrt{\frac{x_{m}c_{m}^{f}}{\tau_{n}}}+\frac{\sqrt{b_{mn}c_{m}^{B}}}{\tau_{n}\sqrt{\log_{2}(1+\varsigma_{mn})}}$.
+$B_{mn}^{*}=\begin{cases}\sqrt{\frac{p_{mn}I_{mn}b_{mn}}{c_{m}^{B}\log_{2}(1+\varsigma_{mn})}},&p_{mn}\geq\frac{F_{mn}^{2}}{I_{mn}},\\ F_{mn}\sqrt{\frac{b_{mn}}{c_{m}^{B}\log_{2}(1+\varsigma_{mn})}},&0<p_{mn}<\frac{F_{mn}^{2}}{I_{mn}},\\ 0,&\text{otherwise},\end{cases}$ (19)
+where $F_{mn}=2\log(1/\theta_{m})\sqrt{\frac{x_{m}c_{m}^{f}}{\tau_{n}}}+\frac{2\sqrt{b_{mn}c_{m}^{B}}}{\tau_{n}\sqrt{\log_{2}(1+\varsigma_{mn})}}$.
 For the detailed proof of Proposition 1, please refer to Appendix A.
@@
 Theorem 2.
-There exists a unique Nash equilibrium in the multi-MSP rewarding game $\Omega$, ensuring a unique optimal set of reward decisions $\{\bm{p}_{n}^{*}\}_{n\in\mathcal{N}}$.
+There exists a unique Nash equilibrium in the multi-MSP rewarding game $\Omega$, ensuring a unique optimal reward decision profile $\{\bm{p}_{n}^{*}\}_{n\in\mathcal{N}}$.
 Proof.
@@
 $\forall m\in\mathcal{M},\ m\neq m^{\prime}.$
-The proof of $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}<0$ and $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}\partial p_{m^{\prime}n}}<0$ can be found in Appendix B.
-We randomly choose a vector $\bm{h}\in\mathbb{R}^{M\times 1}$ with elements not all 00.
-For $\bm{\Lambda}_{n}$, it is easy to see that $\bm{h}^{T}\bm{\Lambda}_{n}\bm{h}=\sum_{m}\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}(h^{m})^{2}<0$.
-Furthermore, $\bm{h}^{T}\bm{H}_{n}\bm{h}=\frac{-\mu_{n}}{(1+\sum_{m}V_{mn})^{2}}\left(\sum_{m}h^{m}V_{mn}^{\prime}\right)^{2}<0$.
-Thus, we have $\bm{h}^{T}(\bm{\Lambda}_{n}+\bm{H_{n}})\bm{h}=\bm{h}^{T}\bm{\Lambda}_{n}\bm{h}+\bm{h}^{T}\bm{H}_{n}\bm{h}<0$, indicating that the utility function $\Psi_{n}$ is strictly concave.
-So far, the uniqueness of the Nash equilibrium [37] solution to the multi-MSP rewarding game $\Omega$ has been proved.
+We randomly choose a vector $\bm{h}\in\mathbb{R}^{M\times 1}$ with elements not all 0.
+Then, we have $\bm{h}^{T}(\bm{\Lambda}_{n}+\bm{H_{n}})\bm{h}=\sum_{m=1}^{M}(h^{m})^{2}(\frac{\mu_{n}V_{mn}^{\prime\prime}}{1+\sum_{m=1}^{M}V_{mn}}-2V_{mn}^{\prime}-p_{mn}V_{mn}^{\prime\prime})-\frac{\mu_{n}}{(1+\sum_{m=1}^{M}V_{mn})^{2}}(\sum_{m=1}^{M}h^{m}V_{mn}^{\prime})^{2}<0$, indicating that the utility function $\Psi_{n}$ is strictly concave.
+According to  [39], there exists a unique Nash equilibrium in the multi-MSP rewarding game $\Omega$.
+The proof of $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}<0$, $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}\partial p_{m^{\prime}n}}<0$, and $\bm{h}^{T}(\bm{\Lambda}_{n}+\bm{H_{n}})\bm{h}<0$ can be found in Appendix B.
 ∎
-1 Input: Action threshold $[p_{min},p_{max}]$ and number of episodes $\mathcal{T}$;
+1Input: Action threshold $p_{min},p_{max}]$ and number of episodes $\mathcal{T}$;
 2 Initialization: local observation $s_{n}$, actor network $\bm{\alpha}_{n}$, critic network $\bm{\beta}_{n}$, and episode buffer $\mathcal{D}_{n}$ for each MSP agent $n\in\mathcal{N}$;
-3 # Trading process using MDDR:
+3
+41ex# Trading process using MDDR:
 5 for episode = $1,2,\ldots,\mathcal{T}$ do
@@
 17 end for
-17Return: Reward decisions $\bm{p}$ of MSPs and resource allocation $[\bm{f},\bm{B}]$ of MUs;
-18 # FL training guided by trading results within time $T$:
+18Return: Reward decisions $\bm{p}$ of MSPs and resource allocation $\bm{f},\bm{B}]$ of MUs;
+19
+201ex# FL training guided by trading results within time $T$:
 21 foreach MSP agent $n\in\mathcal{N}$ do
 22 foreach MU $m\in\mathcal{M}$ do
-21             Based on the trading results, utilize $f_{mn}$ and $B_{mn}$ to perform the local FL task for MSP $n$, and then upload the model to MSP $n$.
+23      Based on the trading results, utilize $f_{mn}$ and $B_{mn}$ to perform the local FL task for MSP $n$, and then upload the local updates to MSP $n$.
 24   end foreach
-23      Aggregate received local models and issue an updated global model;
+25  Aggregate received local updates and issue an updated global model;
 26   Supply the updated model to the AR engine;
@@
 Utilizing Theorem 1 and Theorem 2, we prove the existence and uniqueness of the hierarchical equilibrium in the proposed EPEC model.
-Concretely, each MU has a unique best-response resource allocation strategy in reaction to any given MSP rewards, while each MSP determines a unique optimal pricing to maximize its own utility given the rewards of other MSPs.
+Concretely, each MU has a unique best-response resource allocation strategy in reaction to any given MSP rewards.
+At the same time, each MSP determines unique optimal pricing to maximize its own utility, given the rewards of other MSPs.
 Although the backward induction approach theoretically ensures an equilibrium solution to (14), solving it in real-world vehicular metaverse environments can be challenging due to dynamic network conditions and limited system knowledge (e.g., private cost parameters).
@@
 (ii) MSPs must have prior knowledge of private information about MUs (e.g., cost factors) to determine their rewards, which raises privacy concerns for MUs.
-To address these challenges, we further formulate the multi-MSP rewarding game as a multi-agent Markov decision process (MAMDP) [35], which adapts to dynamic channel conditions.
+To address these challenges, we further formulate the multi-MSP rewarding game as a multi-agent Markov decision process (MAMDP) [36], adapting to dynamic channel conditions.
 In this framework, each MSP is modeled as an individual agent that makes intelligent reward decisions.
@@
 Thus, we set the local observation of MSP $n$ at the $k$-th stage game defined as $s_{n}(k)=[V_{1n}(k),V_{2n}(k),\ldots,V_{mn}(k)]$, and the state of the environment is $\mathcal{S}(k)=[s_{1}(k),\ldots,s_{n}(k)]$.
-At the $k$-th stage game, MSP $n$ observes a state $s_{n}(k)$ and determines an action $\bm{p}_{n}(k)$ within $[p_{min},p_{max}]$.
+At the $k$-th stage game, MSP $n$ observes a state $s_{n}(k)$ and determines an action $\bm{p}_{n}(k)$ within $p_{min},p_{max}]$.
 When an action $\bm{p}_{n}(k)$ is applied to state $s_{n}(k)$, the agent $n$ receives a utility $\mathcal{U}_{n}(k)$ from the environment.
@@
 Finally, the episode buffer is cleared in preparation for the next episode (Line 15).
-The game will terminate when the maximum number of training episodes is reached and then the trading results will be obtained.
+The game terminates once the maximum number of training episodes is reached, after which the trading results are finalized.
 FL training guided by trading results: After obtaining the trading results (MSPs’ reward decisions $\bm{p}$ and resource allocation $(\bm{f},\bm{B})$), MUs and MSPs enter the FL phase.
@@
 During real-time interactions, the MSP’s AR service immersion and quality can be enhanced by utilizing the updated model.
-VI-C The Complexity and Convergence of MDDR
-The complexity of MDDR is mainly influenced by the structure of the actor-critic network.
-According to [39], the total computational complexity of the fully connected layer can be expressed as the number of multiplications $\mathcal{O}(\sum_{u=1}^{U}\varrho_{u}\cdot\varrho_{u-1})$, where $\varrho_{u}$ is the number of neurons in the $u$-th hidden layer.
-In this work, each agent has its actor-critic network, structured as $[5,256,10]$ and $[5,256,1]$, respectively.
-Therefore, its complexity can be expressed as $\mathcal{O}(5\times 256+256\times 10+5\times 256+256\times 1)$ without burden on the MSP.
-After several attempts, the MDDR can achieve convergence [40] (as per Fig.
-7 and Table II) with a suitable set of hyper-parameters.
-TABLE II: The Utilities of three reward approaches
+TABLE III: The Utilities of three reward approaches
 MSP 1 MSP 2 MSP 3 Total
@@
 VII-A Simulation Settings
-[…]
+We conduct experiments on the MNIST dataset with ResNet-18 and the GTSRB dataset with Faster R-CNN.
+The GTSRB was recorded during daytime driving on different types of roads in Germany for vehicle AR scenario testing [41].
+We set the number of MUs as $M=5$ and the number of MSPs as $N=3$.
+The configurations of the maximum computational and communication resource for each MU are randomly generated from the range [3,5][3,5]GHz and [1,4][1,4]MHz [42].
+For local training of MUs, we adopt the SGD optimizer with $\text{momentum}=0.9$ and $\text{learning rate}=0.001$.
+For model aggregation, MSPs adopt the Federated Averaging (FedAvg) algorithm, applying equal weights to all participants’ updates.
+The trading guidance time $T$ for MNIST and GTSRB are $30$s and $1,200$s, respectively.
+All baseline methods are configured with the same settings for fair comparison.
 VII-B Benchmark Schemes
@@
 Therefore, under the uniform rewards of MSPs, we develop and extend four benchmark schemes for multi-MSP and multi-MU scenarios while meeting FL deadlines and resource constraints of devices.
-  • x_based: Inspired by the incentive mechanism of the Stackelberg game for federated learning [31], the MU allocates computational and communication resources based on the amount of data involved in local training.
+  • x_based: Inspired by the incentive mechanism of the Stackelberg game for federated learning [32], the MU allocates its computational and communication resources based on the amount of data involved in local training.
   • w_based: The MU adjusts the resource allocation according to the inference loss [43], i.e., the potential value of local data in our work.
@@
   • fixed: Taking into account the differences in performance of the various devices that exist in reality, we randomly allocate a fixed set of computational and communication resources to reflect the heterogeneity of devices.
-TABLE III: Time comparison of five schemes for achieving the same performance.
+VII-C Performance Evaluation of MDDR
+First, we consider a relatively static communication environment (fixed channel conditions) to show the effectiveness of MDDR compared to the optimal solution and the general multi-agent depth approximation policy optimization (MAPPO) approach.
+We trained our proposed MDDR and MAPPO until convergence, and the results are shown in Fig.
+7 and Table III.
+We find that both MDDR and MAPPO achieve good results, but MAPPO slightly outperforms MDDR.
+This is attributed to MAPPO’s centrally controlled yet distributed execution architecture [44] that can utilize more information for decision making.
+However, this approach is not suitable for non-cooperation due to the need to share parameters among MSPs.
+For the fully distributed MDDR, it is more suitable for non-cooperative scenarios.
+The reason is that in MDDR, each agent has a self-contained actor-critic network, and information is not required to be shared.
+This suggests that the approach can achieve considerable utility while preserving privacy and saving communication costs significantly.
+Then, we deploy the MDDR in dynamic communication environments where channel conditions change over time.
+The trading behaviors of each MU and MSP are illustrated in Figs.
+8(a), 8(b), 8(c) and 8(d), reflecting that the MDDR can converge rapidly and show good stability in dynamic network environments.
+VII-D Benchmark Comparison
+TABLE IV: Time comparison of five schemes for achieving the same performance
 Schemes MNIST GTSRB
-Training Time to reach 90% Accuracy (s) Training Time to reach 50% mAP (s) Training Time to reach 70% mAP@5050{50}50 (s)
+Training Time to reach 90% Accuracy (s) Training Time to reach 50% mAP (s) Training Time to reach 70% mAP@50{50} (s)
 MSP 1 MSP 2 MSP 3 MSP 1 MSP 2 MSP 3 MSP 1 MSP 2 MSP 3
@@
 fixed 30+ 30+ 30+ 1200+ 1200+ 1200+ 1200+ 1200+ 1200+
-  • Note: $30$+ or $1200$+ indicates that the scheme required more than $30$ seconds or $1200$ seconds, respectively, to achieve the target accuracy, exceeding the guideline time for the trading results.
-VII-C Performance Evaluation of MDDR
-First, we consider a relatively static communication environment (fixed channel conditions) to show the effectiveness of MDDR compared to the optimal solution and the general multi-agent depth approximation policy optimization (MAPPO) approach.
-We trained our proposed MDDR and MAPPO until convergence, and the results are shown in Fig.
-7 and Table II.
-We find that both MDDR and MAPPO achieve good results, but MAPPO slightly outperforms MDDR.
-This is because MAPPO is a centrally controlled distributed execution [44] that can utilize more information for decision making.
-However, this approach is not suitable for non-cooperation due to the need to share parameters among MSPs.
-For the fully distributed MDDR, it is more suitable for non-cooperative scenarios.
-The reason is that in MDDR, each agent has a self-contained actor-critic network, and information is not required to be shared.
-This suggests that the approach can achieve considerable utility while preserving privacy and saving communication costs significantly.
-Then, we deploy the MDDR in dynamic communication environments where channel conditions change over time.
-The behaviors of each MU and MSP in the trading are shown in Figs.
-8(a), 8(b), 8(c) and 8(d), reflecting that the MDDR can converge rapidly and show good stability in dynamic network environments.
-VII-D Benchmark Comparison
+  • Note: 3030+ or 12001200+ indicates that the scheme required more than $30$ seconds or $1200$ seconds, respectively, to achieve the target accuracy, exceeding the guideline time for the trading results.
 Figs.
 9(a), 9(b), and 9(c) show the performance improvement in classification accuracy over time, where the immersion-aware outperforms the other schemes.
-This is because immersion-aware utilizes IoM to incentivize MUs to allocate computational and communication resources for providing learning models to MSPs.
-The IoM metric not only evaluates the freshness and accuracy of the model but also considers the amount and potential value of raw data used for training.
-As a result, resource-limited MUs are motivated to contribute more resources, providing valuable learning models to MSPs at an accelerated pace.
-In contrast, baseline schemes like x_based, w_based and w_x_based struggle to offer adequate incentives for MUs to contribute their available resources due to uniform rewards.
-For example, if a given MU has already reached the maximum resource it can contribute when the reward is $p$, then even if $p$ is increased further, the MSP’s gains will be impaired rather than increased.
-Consequently, the uniform reward $p$ will not be increased, leading to a situation where other MUs with available resources will not contribute more resources for providing higher-value learning models.
+This is because immersion-aware utilizes IoM to incentivize MUs to allocate computational and communication resources for providing local models to MSPs.
+The IoM metric not only evaluates the freshness and accuracy of the local model but also considers both the amount and potential value of raw data used for training.
+As a result, resource-constrained MUs are motivated to contribute more resources, providing valuable local models to MSPs more rapidly.
+In contrast, baseline schemes like x_based, w_based, and w_x_based struggle to offer adequate incentives for MUs to contribute their available resources due to uniform rewards.
+For example, if an MU has already reached its maximum allocatable resources at a given reward $p$, increasing $p$ further will not yield additional contributions.
+In this situation, raising the reward only increases the MSP’s cost without eliciting additional contributions, so its utility actually declines.
+Consequently, the uniform reward $p$ remains unchanged, preventing other MUs with more resources from contributing higher-value local models.
 In addition, Figs.
@@
 For example, at the $10$th second, the accuracy of MSP 3 is about $88\%$, which is $14\%$ and $3\%$ higher than those of MSP 1 and MSP 2, respectively.
-This is because the potential value $\omega$ is higher for MSP 3, leading to more rewards being given by MSP 3.
-As a result, the MUs will allocate more computational and communication resources to provide learning models for MSP 3 than for MSP 1 and MSP 2.
+This is because MSP 3 is associated with a higher potential value $\omega$, resulting in more substantial rewards being offered.
+As a result, the MUs will allocate more computational and communication resources to provide local models for MSP 3 than for MSP 1 and MSP 2.
 For the three different MSPs, some fluctuations are normal for model training (e.g., at time $10$s).
-It can be seen that the effect of fixed is the worst due to synchronous FL, which is limited by the MU with the least resource.
+Notably, the fixed scheme performs the worst due to the limitations of synchronous FL, which is bottlenecked by the least capable MU.
 Meanwhile, it is clear from Figs.
@@
 With the higher IoM, the performance is improved faster, resulting in a better immersive experience.
-Yet, the IoM of w_based in Fig.
-9(e) is higher than x_based and w_x_based but the performance improvement rate is not as fast as them.
-The reason is that the IoM here represents the sum of all IoM from MUs, whereas the rate of performance improvement is hampered by the worst MU.
-Additionally,from Figs.
+Although the IoM of w_based in Fig.
+9(e) is higher than x_based and w_x_based, its performance improvement rate is comparatively slower.
+This is because the IoM represents the aggregate value across all MUs, while the performance improvement rate is hampered by the worst MU.
+Furthermore, from Figs.
 9(b) and 9(c), it can be observed that both w_x_based and w_based perform better when compared to Fig.
@@
 For object detection on GTSRB, Figs.
-10(a), 10(b) and 10(c) show the mean average precision (mAP) and Table III gives the mAP@5050{50}50 (The value of mAP when the intersection over union (IOU) threshold is larger than $0.5$) results.
-We observe that compared with MNIST, the performance enhancement for object detection is slower due to the complexity of the detection model.
+10(a), 10(b) and 10(c) show the mean average precision (mAP) and Table IV gives the mAP@50{50} (The value of mAP when the intersection over union (IOU) threshold is larger than $0.5$) results.
+Compared to MNIST, the performance enhancement for object detection is slower due to the complexity of the detection task.
 Specifically, it takes at least $513.16$ seconds to achieve a $50\%$ mAP using the immersion-aware method.
 Our research introduces the flexibility to adjust decision guidance time, making it adaptable to various task types.
-Meanwhile, Figs.
-10(d)10(e)10(f) and Table III demonstrate that MSP 1 with greater IoM takes less time to achieve the same mAP, which indicates that IoM works well to capture immersion.
+Moreover, Figs.
+10(d)10(e)10(f) and Table IV demonstrate that MSP 1 with greater IoM takes less time to achieve the same mAP, which indicates that IoM works well to capture immersion.
 For instance, MSP 1 has the highest IoM, and the time that it takes to reach $50\%$ mAP is $513.16$s, which is $67.93$s and $110.73$s less than those of MSP 2 and MSP 3, respectively.
@@
 VIII Conclusion
-To facilitate the development of vehicular metaverse services, we propose an immersion-aware model trading framework that incorporates FL to incentivize MUs to contribute locally trained data for AR services such as object detection and classification.
+To facilitate the development of vehicular metaverse services, we propose an immersion-aware model trading framework, incorporating FL to incentivize MUs to contribute local models for AR services such as object detection and classification.
 Specifically, we construct an EPEC problem with MSPs as leaders and MUs as followers to achieve an equilibrium among their interests.
-A new metric called “IoM” is designed to comprehensively evaluate the enhancement brought by the learning models of MUs for AR services.
-In addition, considering the competitive relationship among MSPs and the dynamic network environment, we develop a fully distributed MDDR approach to obtain the reward decisions of MSPs.
+A new metric called “IoM” is designed to comprehensively evaluate the enhancement brought by the local models of MUs for AR services.
+Furthermore, considering the competitive relationship among MSPs and the dynamic network environment, we develop a fully distributed MDDR approach to obtain the reward decisions of MSPs.
 Extensive simulations on AR-related vehicle and MNIST datasets demonstrate that the proposed framework enables more efficient and immersive AR services in the vehicular metaverse.
@@
 First, we transform Problem 1 into an unconstrained optimization problem using the Lagrangian dual method [45, 46]. The Lagrangian function associated with MU $m$ is formulated as
-$L_{m}\left(\bm{f}_{m},\bm{B}_{m},\bm{p}_{m},\lambda_{m},\beta_{m},\delta_{mn},\gamma_{m}\right)=\sum_{n}p_{mn}V_{mn}$ (24)
-$-\sum_{n}\left[c_{m}^{f}f_{mn}\log(1/\theta_{m})+c_{m}^{B}B_{mn}\right]$
+$\ L_{m}\left(\bm{f}_{m},\bm{B}_{m},\bm{p}_{m},\lambda_{m},\beta_{m},\delta_{mn},\gamma_{m}\right)$ (24)
+$=$ $\ \sum_{n}p_{mn}V_{mn}-\sum_{n}\left[c_{m}^{f}f_{mn}\log(1/\theta_{m})+c_{m}^{B}B_{mn}\right]$
 $-\lambda_{m}(\sum_{n}f_{mn}-f_{m}^{max})-\beta_{m}(\sum_{n}B_{mn}-B_{m}^{max})$
-$-\sum_{n}\delta_{mn}[\log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}\log_{2}(1+\varsigma_{mn})}-\tau_{n}]$
+$-\sum_{n}\delta_{mn}[\log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}\log_{2}(1+\varsigma_{mn})}-\frac{1}{2}\tau_{n}]$
 $-\gamma_{m}[S_{m}-T^{req}(f_{m}^{max}-\sum_{n}f_{mn})],$
-where $\lambda_{m}\geq 0,\beta_{m}\geq 0$, $\delta_{mn}\geq 0$ and $\gamma_{m}\geq 0$ are the Lagrangian dual variables corresponding to constraints C1, C2 and C3, respectively.
+where ($\lambda_{m}\geq 0$ and $\beta_{m}\geq 0$), ($\delta_{mn}\geq 0$), and ($\gamma_{m}\geq 0$) are the Lagrangian dual variables corresponding to constraints C1, C2, and C3, respectively.
 Then, the first-order derivative of $L_{m}$ with respect to $f_{mn}$ and $B_{mn}$ can be derived as
@@
 $\frac{\partial L_{m}}{\partial B_{mn}}=$ $\ p_{mn}I_{mn}\frac{b_{mn}}{B_{mn}^{2}\log_{2}(1+\varsigma_{mn})}-c_{m}^{B}-\beta_{m}$ (26)
-$+\delta_{mn}\frac{b_{mn}}{B_{mn}^{2}\log_{2}(1+\varsigma_{mn})}.$
-From (25) and the constraints in (11), the Karush-Kuhn-Tucker (KKT) [47] conditions are given by
+$+\delta_{mn}\frac{b_{mn}}{B_{mn}^{2}\log_{2}(1+\varsigma_{mn})},$
+respectively. From (25) and the constraints in (11), the Karush-Kuhn-Tucker (KKT) [47] conditions are given by
 $\frac{\partial L_{m}}{\partial f_{mn}}=0,\ \lambda_{m}(\sum_{n}f_{mn}-f_{m}^{max})=0,$ (27)
-$\delta_{mn}[\log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}\log_{2}(1+\varsigma_{mn})}-\tau_{n}]=0,$
+$\delta_{mn}[\log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}\log_{2}(1+\varsigma_{mn})}-\frac{1}{2}\tau_{n}]=0,$
 $\gamma_{m}[S_{m}-T^{req}(f_{m}^{max}-\sum_{n}f_{mn})]=0,$
@@
 $\frac{\partial L_{m}}{\partial B_{mn}}=0,\ \beta_{m}(\sum_{n}B_{mn}-B_{m}^{max})=0,$ (29)
-$\delta_{mn}[\log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}\log_{2}(1+\varsigma_{mn})}-\tau_{n}]=0,$
+$\delta_{mn}[\log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}\log_{2}(1+\varsigma_{mn})}-\frac{1}{2}\tau_{n}]=0,$
 $\beta_{m},\delta_{mn}\geq 0,\ C1,\ C2.$
@@
 $(p_{mn}I_{mn}+\delta_{mn})\frac{b_{mn}}{B_{mn}^{2}\log_{2}(1+\varsigma_{mn})}-c_{m}^{B}=0.$ (30)
-Then, based on the value of $\delta_{mn}$, the optimal allocation of computation and communication resources exists in the following two cases:
+Then, based on the value of $\delta_{mn}$, the optimal allocation of computational and communication resources exists in the following two cases:
   • ($\delta_{mn}=0$): According to (28) and (30), we have
     $f_{mn}=\sqrt{\frac{p_{mn}I_{mn}x_{m}\tau_{n}}{c_{m}^{f}}},\ B_{mn}=\sqrt{\frac{p_{mn}I_{mn}b_{mn}}{c_{m}^{B}\log_{2}(1+\varsigma_{mn})}}.$ (31)
-  • ($\delta_{mn}>0$): Substituting $f_{mn}$ and $B_{mn}$ into $\delta_{mn}\left[\log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}\log_{2}(1+\varsigma_{mn})}-\tau_{n}\right]=0$, we can get
+  • ($\delta_{mn}>0$): Substituting $f_{mn}$ and $B_{mn}$ into $\delta_{mn}\left[\log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}\log_{2}(1+\varsigma_{mn})}-\frac{1}{2}\tau_{n}\right]=0$, we can get
     $\delta_{mn}=F^{2}-p_{mn}I_{mn},$ (32)
-    $F=\log(1/\theta_{m})\sqrt{\frac{x_{m}c_{m}^{f}}{\tau_{n}}}+\frac{\sqrt{b_{mn}c_{m}^{B}}}{\tau_{n}\sqrt{\log_{2}(1+\varsigma_{mn})}}.$
+    $F=2\log(1/\theta_{m})\sqrt{\frac{x_{m}c_{m}^{f}}{\tau_{n}}}+\frac{2\sqrt{b_{mn}c_{m}^{B}}}{\tau_{n}\sqrt{\log_{2}(1+\varsigma_{mn})}}.$
     Next, bring $\delta_{mn}$ into $f_{mn}$ and $B_{mn}$, we have
     $f_{mn}=F\sqrt{\frac{x_{m}\tau_{n}}{c_{m}^{f}}},\ B_{mn}=F\sqrt{\frac{b_{mn}}{c_{m}^{B}\log_{2}(1+\varsigma_{mn})}}.$ (33)
-Appendix B Proof of $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}<0$ and $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}\partial p_{m^{\prime}n}}<0$
+Appendix B Proof of Theorem 2
 Proof.
@@
 Consequently, we can obtain
-$\frac{\partial f_{mn}}{\partial p_{mn}}=\frac{I_{mn}f_{mn}}{2(p_{mn}I_{mn}+\delta_{mn})}>0.$ (34)
+$\frac{\partial f_{mn}}{\partial p_{mn}}=\frac{I_{mn}f_{mn}}{2(p_{mn}I_{mn}+\delta_{mn})}>0,$ (34)
 the second-order derivative of $\frac{\partial L_{m}}{\partial f_{mn}}=0$ with respect to $p_{mn}$ is expressed as
-$\frac{\partial^{2}f_{mn}}{\partial p_{mn}^{2}}=\frac{I_{mn}\frac{\partial f_{mn}}{\partial p_{mn}}(p_{mn}I_{mn}+\delta_{mn})-I_{mn}^{2}f_{mn}}{2(p_{mn}I_{mn}+\delta_{mn})^{2}},$ (35)
+$\frac{\partial^{2}f_{mn}}{\partial p_{mn}^{2}}=\frac{I_{mn}\frac{\partial f_{mn}}{\partial p_{mn}}(p_{mn}I_{mn}+\delta_{mn})-I_{mn}^{2}f_{mn}}{2(p_{mn}I_{mn}+\delta_{mn})^{2}}.$ (35)
 By substituting (34) into (35), we have
@@
 Similarly, we obtain the second-order derivative of $\frac{\partial L_{m}}{\partial B_{mn}}=0$ with respect to $p_{mn}$ as
-$\frac{\partial^{2}B_{mn}}{\partial p_{mn}^{2}}=\frac{I_{mn}\frac{\partial B_{mn}}{\partial p_{mn}}(p_{mn}I_{mn}+\delta_{mn})-I_{mn}^{2}B_{mn}}{2(p_{mn}I_{mn}+\delta_{mn})^{2}},$ (38)
+$\frac{\partial^{2}B_{mn}}{\partial p_{mn}^{2}}=\frac{I_{mn}\frac{\partial B_{mn}}{\partial p_{mn}}(p_{mn}I_{mn}+\delta_{mn})-I_{mn}^{2}B_{mn}}{2(p_{mn}I_{mn}+\delta_{mn})^{2}}.$ (38)
 By substituting (37) into (38), we have
 $\frac{\partial^{2}B_{mn}}{\partial p_{mn}^{2}}=\frac{-I_{mn}^{2}B_{mn}}{4(p_{mn}I_{mn}+\delta_{mn})^{2}}<0.$ (39)
-Based on (34), (36), (37), and (39), we can easily infer that $V_{mn}^{\prime}>0$ and $V_{mn}^{\prime\prime}<0$, as shown in (40).
-Then, we can obtain $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}\partial p_{m^{\prime}n}}<0$.
-For $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}$, We substitute the specific expressions for $V_{mn}^{\prime}$ and $V_{mn}^{\prime\prime}$ into (22) and obtain the following result as shown in (41).
-Next, we randomly choose a vector $\bm{h}\in\mathbb{R}^{M\times 1}$ with elements not all 00.
-For $\bm{\Lambda}_{n}$, it is easy to see that $\bm{h}^{T}\bm{\Lambda}_{n}\bm{h}=\sum_{m}\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}(h^{m})^{2}<0$.
-Furthermore, $\bm{h}^{T}\bm{H}_{n}\bm{h}=\frac{-\mu_{n}}{(1+\sum_{m}V_{mn})^{2}}\left(\sum_{m}h^{m}V_{mn}^{\prime}\right)^{2}<0$.
-Thus, we have $\bm{h}^{T}(\bm{\Lambda}_{n}+\bm{H_{n}})\bm{h}=\bm{h}^{T}\bm{\Lambda}_{n}\bm{h}+\bm{h}^{T}\bm{H}_{n}\bm{h}<0$, indicating that the utility function $\Psi_{n}$ is strictly concave.
-So far, the uniqueness of the Nash equilibrium [37] solution to the multi-MSP rewarding game $\Omega$ has been proved.
 $V_{mn}^{\prime}=\frac{\partial V_{mn}}{\partial p_{mn}}=I_{mn}$ $\left(\frac{x_{m}\tau_{n}\log(1/\theta_{m})}{f_{mn}^{2}}\cdot\frac{\partial f_{mn}}{\partial p_{mn}}+\frac{b_{mn}}{B_{mn}^{2}\log_{2}(1+\varsigma_{mn})}\cdot\frac{\partial B_{mn}}{\partial p_{mn}}\right)>0,$ (40)
@@
 $\left.-\frac{2b_{mn}}{B_{mn}^{3}\log_{2}(1+\varsigma_{mn})}\cdot(\frac{\partial B_{mn}}{\partial p_{mn}})^{2}+\frac{b_{mn}}{B_{mn}^{2}\log_{2}(1+\varsigma_{mn})}\cdot\frac{\partial^{2}B_{mn}}{\partial p_{mn}^{2}}\right]<0.$
-$\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}=-\left[\frac{4a_{1}I_{mn}^{2}\delta_{mn}+a_{1}I_{mn}^{3}p_{mn}}{4f_{mn}(\delta_{mn}+I_{mn}p_{mn})^{2}}+\frac{4b_{mn}I_{mn}^{2}\delta_{mn}+b_{mn}I_{mn}^{3}p_{mn}}{4a_{2}B_{mn}(\delta_{mn}+I_{mn}p_{mn})^{2}}+I_{mn}^{2}\frac{\left(\frac{a_{1}a_{2}I_{mn}B_{mn}+b_{mn}I_{mn}f_{mn}}{2a_{2}f_{mn}B_{mn}(\delta_{mn}+I_{mn}p_{mn})}\right)^{2}}{(1+\sum_{m}V_{mn})^{2}}\right.$ (41)
-$\left.+\frac{3I_{mn}^{3}(a_{1}a_{2}B_{mn}+b_{mn}f_{mn})}{4a_{2}f_{mn}B_{mn}(\delta_{mn}+I_{mn}p_{mn})^{2}(1+\sum_{m}V_{mn})}\right]<0,\text{where}\,a_{1}=x_{m}\tau_{n}\log(1/\theta_{m}),a_{2}=\log_{2}(1+\varsigma_{mn}).$
+$\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}=-\Bigg[\frac{aI_{mn}^{3}p_{mn}+4aI_{mn}^{2}\delta_{mn}}{4f_{mn}(p_{mn}I_{mn}+\delta_{mn})^{2}}+\frac{bI_{mn}^{3}p_{mn}+4bI_{mn}^{2}\delta_{mn}}{4B_{mn}(p_{mn}I_{mn}+\delta_{mn})^{2}}+\frac{\mu_{n}I_{mn}^{2}\left(\frac{aI_{mn}}{2f_{mn}(p_{mn}I_{mn}+\delta_{mn})}+\frac{bI_{mn}}{2B_{mn}(p_{mn}I_{mn}+\delta_{mn})}\right)^{2}}{\left(1+\sum_{m}V_{mn}\right)^{2}}$ (41)
+$\qquad\qquad\qquad+\frac{\mu_{n}I_{mn}\left(\frac{3aI_{mn}^{2}}{4f_{mn}(p_{mn}I_{mn}+\delta_{mn})^{2}}+\frac{3bI_{mn}^{2}}{4B_{mn}(p_{mn}I_{mn}+\delta_{mn})^{2}}\right)}{\left(1+\sum_{m}V_{mn}\right)}\Bigg]<0,$
+$\text{where}\,a=x_{m}\tau_{n}\log(1/\theta_{m}),\ b=\frac{b_{mn}}{\log_{2}(1+\varsigma_{mn})}.$
+Based on (34), (36), (37), and (39), we can easily infer that $V_{mn}^{\prime}>0$ and $V_{mn}^{\prime\prime}<0$, as shown in (40).
+Then, we can obtain $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}\partial p_{m^{\prime}n}}<0$.
+For $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}$, we substitute the specific expressions for $V_{mn}^{\prime}$ and $V_{mn}^{\prime\prime}$ into (22) and obtain the following result as shown in (LABEL:result).
+Next, by randomly selecting a vector $\bm{h}\in\mathbb{R}^{M\times 1}$ with elements not all 0, we obtain
+$\ \bm{h}^{\top}(\bm{\Lambda}_{n}+\bm{H}_{n})\bm{h}$
+$=$ $\underbrace{\sum_{m=1}^{M}\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}(h^{m})^{2}}_{\text{diagonal part}}+\underbrace{\sum_{m\neq m^{\prime}}h^{m}\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}\partial p_{{m^{\prime}}n}}h^{m^{\prime}}}_{\text{off-diagonal part}}.$ (42)
+Expanding the expressions and letting $S=\sum_{m=1}V_{mn}$, we have
+$\ \quad\bm{h}^{\top}(\bm{\Lambda}_{n}+\bm{H}_{n})\bm{h}$
+$=\sum_{m\neq{m^{\prime}}}h^{m}\left(-\mu_{n}\frac{V_{mn}^{\prime}V_{{m^{\prime}}n}^{\prime}}{(1+S)^{2}}\right)h^{m^{\prime}}$
+$+\sum_{m=1}^{M}\left[\mu_{n}\frac{V_{mn}^{\prime\prime}(1+S)-(V_{mn}^{\prime})^{2}}{(1+S)^{2}}-2V_{mn}^{\prime}-p_{mn}V_{mn}^{\prime\prime}\right](h^{m})^{2}.$ (43)
+Specially, the off-diagonal part is transformed using the identity
+$\ \sum_{m\neq m^{\prime}}h^{m}V_{mn}^{\prime}V_{{m^{\prime}}n}^{\prime}h^{m^{\prime}}$
+$=$ $\left(\sum_{m=1}^{M}h^{m}V_{mn}^{\prime}\right)^{2}-\sum_{m=1}^{M}(h^{m})^{2}(V_{mn}^{\prime})^{2},$ (44)
+thus,
+$\sum_{m\neq{m^{\prime}}}h^{m}\left(-\mu_{n}\frac{V_{mn}^{\prime}V_{{m^{\prime}}n}^{\prime}}{(1+S)^{2}}\right)h^{m^{\prime}}$
+$=$ $-\frac{\mu_{n}}{(1+S)^{2}}\left[\left(\sum_{m=1}^{M}h^{m}V_{mn}^{\prime}\right)^{2}-\sum_{m=1}^{M}(h^{m})^{2}(V_{mn}^{\prime})^{2}\right].$ (45)
+Then, combining both parts, we have
+$\ \bm{h}^{\top}(\bm{\Lambda}_{n}+\bm{H}_{n})\bm{h}$
+$=$ $\sum_{m=1}^{M}\left[\mu_{n}\frac{V_{mn}^{\prime\prime}(1+S)-(V_{mn}^{\prime})^{2}}{(1+S)^{2}}-2V_{mn}^{\prime}-p_{mn}V_{mn}^{\prime\prime}\right](h^{m})^{2}$
+$\quad-\frac{\mu_{n}}{(1+S)^{2}}\left[\left(\sum_{m=1}^{M}h^{m}V_{mn}^{\prime}\right)^{2}-\sum_{m=1}^{M}(h^{m})^{2}(V_{mn}^{\prime})^{2}\right].$ (46)
+Moreover, the expression can be simplified to
+$\ \bm{h}^{\top}(\bm{\Lambda}_{n}+\bm{H}_{n})\bm{h}$
+$=$ $-\frac{\mu_{n}}{(1+S)^{2}}\left(\sum_{m=1}^{M}h^{m}V_{mn}^{\prime}\right)^{2}$
+$+\sum_{m=1}^{M}(h^{m})^{2}\left(\mu_{n}\frac{V_{mn}^{\prime\prime}}{1+S}-2V_{mn}^{\prime}-p_{mn}V_{mn}^{\prime\prime}\right).$ (47)
+Finally, we analyze the sign of this expression:
+  1. 1.
+    Since $\mu_{n}>0$ and $(1+S)^{2}>0$, the first term is non-positive.
+  2. 2.
+    Plugging (40) into the second term yields $(\mu_{n}\frac{V_{mn}^{\prime\prime}}{1+S}-2V_{mn}^{\prime}-p_{mn}V_{mn}^{\prime\prime})<0$, as shown in (LABEL:second-term).
+    $\mu_{n}\frac{V_{mn}^{\prime\prime}}{1+S}-2V_{mn}^{\prime}-p_{mn}V_{mn}^{\prime\prime}=-\Bigg[\frac{aI_{mn}^{3}p_{mn}+4aI_{mn}^{2}\delta_{mn}}{4f_{mn}(p_{mn}I_{mn}+\delta_{mn})^{2}}+\frac{\mu_{n}I_{mn}\left(\frac{3aI_{mn}^{2}}{4f_{mn}(p_{mn}I_{mn}+\delta_{mn})^{2}}+\frac{3bI_{mn}^{2}}{4B_{mn}(p_{mn}I_{mn}+\delta_{mn})^{2}}\right)}{\left(1+S\right)}\Bigg]<0,$ (48)
+    $\text{where}\,a=x_{m}\tau_{n}\log(1/\theta_{m}),\ b=\frac{b_{mn}}{\log_{2}(1+\varsigma_{mn})}.$
+Based on the above, we can have $\bm{h}^{T}(\bm{\Lambda}_{n}+\bm{H_{n}})\bm{h}<0$, indicating that the utility function $\Psi_{n}$ is strictly concave.
+According to [39], there exists a unique Nash equilibrium in the multi-MSP rewarding game $\Omega$.
+∎